A Complete Bibliography of Publications in Numerische Mathematik (2010–2019)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

30 November 2023 Version 1.65

Title word cross-reference

(−1, 1) [DMN12]. 1 [FR19, RS11]. 2 [HR19, HOR17a, TWW15]. 2/3 [BT15]. 3 [BKW10, CCZ13, EMR19, LCH12]. 5 [BKW10]. α [ABC15]. B [BCM17, MMMKV16]. C^{0} [BNRS17, CCQ13, SH18, DLT12]. C^{r} [Mat13]. δ [YS13]. dG(1) [CLA19]. H [BPZ12, FKS18, FKS19, FLOP10, GMS19, dVBRS11]. H(curl; Ω) [HLZ12]. H(div) [dVBMR16]. H(curl) [dVBMR16]. H(curl, div) [DLT13]. H(div) [AT19]. h/2 [FLOP10]. H^{1} [BY14, LC14]. H^{2}H^{2} [Bör10]. H^{2} [LMcS17]. hp [BDGQ12, HM13, dVCMR18, dVMM19]. k [dVBRS11]. L [Pru14]. L^{2} [Har18, LC14, Mak18, NW18]. L^{∞} [Hac13, LC14, CC10b]. L^{p} [Li15]. L_{2} [BY14, Wil19, DS11]. L_{p} [BLM11a]. M [JLL15, XL17]. hp [CNSV17]. Lp [CM18]. H [DHS17, FMP15]. n = 5 [HJZ14]. ∇ · B = 0 [HMX17]. p [EGLTP17, dVBRS11]. P_{1} [LL10]. QR [CH19]. r [Dic12]. S^{2} [BD12b]. T [CC13, CFL11]. θ [GAB13]. H(div)^{∗} [CPS16]. W [LV13]. W^{1,p} [Mir12].

-AFEM [CNSV17]. -BEM [HM13, GMS19]. -coercivity [CC13]. -conforming [AT19, dVBMR16]. -D [CCZ13, FR19, HR19]. -designs [CFL11]. -discretization [Har18, NW18]. -elements [LMcS17]. -elliptic [DLT13, HLZ12]. -error [CM18]. -matrices [FKS19, Bör10, FKS18]. -matrix [DHS17, FMP15, JLL15, XL17]. -method [GAB13]. -methods [LV13]. -nonconforming [CCQ13]. -norm [EGLTP17]. -projection [BY14]. -refinement [dVBRS11]. -series [BCM17, MMMKV16]. -singularities [YS13]. -stability [BY14]. -stable [Pru14].

1 [SB13a].

2 [SB13b].

A-posteriori [RW12]. Absorbing [ABK13]. Abstract [SDH^{+}14, Mas16, FGP10b]. Accelerated [SX15b]. Acceleration [Sid11a, SCHH13]. Accuracy [HR14, BGR18, MR17]. Accurate [Koe19, XXL12b, XXL12a, BZ17, CCLM15, CGSW16, DWWW17, DK11, DH17, Lin17, NP15, XL17]. acoustic [KK15c, MP18b, dVMRR17]. action [BL13a]. active [GGGS19]. actual [NN11]. adaptations [DMG19]. adapted [EÖ17, Mir12]. Adaptive [AORS17, CHTV18, DM16, EPS17, FGH^{+}16, FHPS19, Gan13, GG12, GMSS11, Git14, GY17, KSD11, Mir16, SV16, BC17, CGMM11, CG11, CPR13, CGMM14, CSW16, CPB17, CXH10, CCZ13, DS11, Dem16, Dem17, EMR19, FGHP17, FLOP10, Gal15, GK19a, HSX12, KdS18, KS11c, OB17, Pla15, Rei12, RS18, SZ13, SH18]. adaptivity [KK15a, Rad19, dVMM19]. Additive [LMR16, Bre13, DK16b, GS16]. ADI [BBKS18, HJS15, MOR17]. ADI-type [BBKS18]. Adiabatic [JM18]. adjoint [AN19, BBB17, CEM14, GGMO16]. adjoints [BMWB14]. advection [AH14, AH17, BGGH16, DS18c]. advection-diffusion [AH14]. AFEM [CNSV17]. affine [KR16, MMMKV16]. aggregation [PCCC18]. ALE [BKN13]. algebra [BCM17]. algebraic [AR16, BBK17, BM13, BBKS18, BMP10, Men11, Not16b, PSV18, XXL12a, XL17]. Algorithm [SCHH13, VVVF10, BCN19, BCPS10, BBKS18, CGSW16, CH19, Dic12, EMB10, HHYY14, HLS11, Ise11, JMM12, KLL19, MMMS10, OYV17, Rie19, SG11, ST16, Tu11]. algorithms [BLF19, BCGS11, CHTV18, DMG19, HH11, Kir11, NP15, Nie10, PT10, SX13, SX15a, SX15b, XL17]. aliasing [BT15]. all-at-once [Tak15]. Allen [BM11, CCS17]. Almost [WMS10, FHPS19, OTMS13, TOMS13]. alternating [HY15]. alternative [Pin18]. Ampère [BCMO16, Nei10, NZ19]. analog [BY19]. Analysis [ABER10, BD12a, BGJ18, BIM11, Ber16, CLLS19, DLT13, FKNP11, GGM14, GV15, GGK^{+}19, GGRG15, HPS16, KP17, LOSV16, PS11, QRB11, RS11, RE14, Run14, SS18, SCHH13, TWZ17, XZ10, Zha16a, ABM18, AMORB16, AL17, ABC15, AG10, BDQ10, BL13a, BB12, BRK17, BJLZ15, BBD16, BDLM18, BL13b, BC15a, BDQ11, BKN13, BM17, Bre13, BDS13, CDNP16, CPS16, CGG19, CC19, CDL19, CP14, CS12, DES11, DS18c, DWWW17, DM14, DHM19, DGS15, EH18, EH13, GMS19, GW19, GMP^{+}14, GKN^{+}18, GMT11, GSV19, HR12, HL16, HSD18, HS19a, HLZ12, HP17, HPS18, HKT14, HL18, KS11a, KPR14, LC14, LS16, LN18, LCWW17, LJS18, MOS11, Mak18, NGKN10, NvPZ10, Nou19, PS18, PGW18, RMS17, ST16, SB13a, SX13, TSGU13, TWW15, TT15, VGG14, Wen10]. analysis [Wil19, WS19, ZS17, ZDQ17, dHQS15, dVBRS11]. Analytic [CG17]. and/of [TWW15]. angle [HKK12]. Anisotropic [Mon19, BGN12, CCP13, CCZ13, Kop17, LN18, LHQ14, Mir16, SY17]. anti [AR19]. anti-Gauss [AR19]. any [ZZ15]. application [Adc10, BK12, BC15a, CEM14, Che12, CC13, CLM15, GV15, HSLZ19, HW18, Hor17b, HSZ13, NTT16]. applications [ABM18, AHR14, BL13a, BCPS10, BDQ11, HW13, JZ13, KKN14, Nie10, NTW18, NOS16, YS13, ZT12]. applied [FGP10a, FGP10b]. Applying [GGS15]. approach [BO19, BG17, CGMM11, CCGP17, CH18b, CM10, DS18c, FY11, GWCH10, GP11, GGGS19, Mak18, NPVW19, Not16b, OPS15, BEG14]. approximability [FMP15]. Approximate [BH11, BAK19, BCCH19, BKW10, Rus10]. Approximating [BV14, Pin18, Wan12]. Approximation [BMWB14, BS17, Bör10, BC16, BM17, BS19, CHW12, CMP13, vyKS14, ABB19, AP11, ABB15, BLM11a, BBL11, BY15, BO17, BGN12, BM11, Bar13b, BHK10, BKV15, BG12, BBD16, BDQ11, BLP19, BHL^{+}19, CD11, CMR18, DMG19, DCC18, DU15, DHS17, DMMS11, Egg19, Eng14, EZ15, EH12, GT14, GK19a, GSY17, GGRG15, HSW17, HR10, KKR12, KOZ16, KK15c, KS11b, LSW11, LMMR19, LLY15, NTT16, NTW18, Nou19, Nov14, OMS13, PC13, RHM13, SY17, SG… 12, Wac17, XCW10, ZT12]. approximations [AN19, BCS14a, Bal12, BFG13, BPS10, BBK17, BV12, BR17, BMO15, BEJ14, CCP13, CEM14, EK18, ER18, Fro18, GGMO16, GNS15, GEEF17, HR14, HR12, HJT14, HL18, KK15a, KR16, LM11, LHQ14, LM13, NS14, TWZ17]. arbitrary [Mir12, Mon19]. argument [Fuk10]. arising [BLM11b, BG16, DD16, Eng14, GMP^{+}14, PG17]. Arnold [CGG19]. Arnoldi [BG12, ST16]. Artificial [BMGN18]. assimilation [BO18]. associated [BCD15, GK19b]. assumption [CGG16]. assumptions [CC19]. Asymptotic [FS14, HL19, OYV17, SH17, WZH13, ADBN16, BDF12, CGMM14, HJNS17]. asymptotically [Pla15, SV12]. Asymptotics [LT10, BLM11a]. Asynchronous [MSV17]. atomistic [LOSV16, MOS11, OLOV18]. atomistic/continuum [LOSV16, MOS11, OLOV18]. attractors [Pie18, ZT12]. Augmented [CC10a, BB12, Jin17, LM15]. Automatic [Pla15, BMWB14]. auxiliary [ZLL^{+}12]. average [DH17, LMR16]. Averaging [Dal10, CLM17]. axisymmetric [BGN19].

B [RMS17]. B-splines [RMS17]. Babuska [Mak18]. backscattering [HHR11]. Backward [ST16, AL15, DLPV18, GMT11, LLW10]. Balanced [Sin12, BL17, CDKLM18]. Balancing [Zam13]. Banach [FN12, JZ13, JZ14, JY16, KdS18, Ste15, WZ10, ZWJ19, dHQS15]. band [Eng14, GG12]. barrier [MOGO17]. barycentric [BDHK12, BRZ13, CH18b, Mas14, MdC17]. based [Bal12, BBK17, BKW10, CGG19, CHTV18, CS17, CH12, DS16, DSX16, EH18, FLOP10, FHN^{+}14, GS17, GGM14, GH14a, GH14b, GP11, HS19a, HT18, KT12, LA11, LFS16, MOS11, Mat13, NTW18, Nou19, OLOV18, RMS17, WX12, Wei11]. bases [BV12, MV15]. basis [CQR16, HSW17, PT14, RHM13, SV13]. BDDC [Sou13, Tu11]. BDF [BMWB14]. BE [GMSS11]. beam [DGBL^{+}15]. Bear [CLLS19]. Beavers [CGHW11]. Bellman [BMZ10, SS16]. Beltrami [BHL^{+}19]. BEM [HM13, Bal12, BBSV13, BGIS17, BN16, FFKP15, FHPS19, GMS19, GHS12, OB17, Wei11]. BEM-based [Wei11]. BEM-FEM [GHS12]. bending [Bar13b, Zha16a, dVM11]. Benjamin [BMGN18, DHKR16]. Bernoulli [Kre17]. Bernstein [Kir11, KT12, Kir17, PS16]. Bessel [Hor17b]. best [BY15, BO17, HL18]. between [CMR18, Fer13b, HU17, Jun15]. Bézout [NNT15]. BGK [ADBN16]. bidiagonal [Bar13a]. bidomain [Zam13]. Bielak [GHS12]. bifurcations [Chá11, HJNS17]. biharmonic [BAK19, CG14, PS13]. binary [YBTB14]. Bingham [CRS16]. biology [GT14]. bioluminescence [GWCH10]. Biot [AKYZ18]. Birkhoff [FGP10a, FGP10b]. bisection [CNX12]. Bivariate [EKAB16, Che12, NNT15]. blended [LOSV16]. blending [OLOV18]. Block [DLPV18, AR19, BLM11a, BS13, Hac16, Har15, LZ15, OYV17]. block-Jacobi [OYV17]. block-structured [Hac16]. blocks [Koe19]. blow [Cho18, HW18, ZS17]. blow-up [Cho18, HW18, ZS17]. Bona [BMGN18]. both [HHS15, HM16]. Bouligand [CN19]. bound [FLMP12, GMT11]. bound-constrained [GMT11]. Boundary [GMÖ^{+}18, HJHA17, Ste14, VGG14, Adc10, ABK13, BY19, BHT17, BIM11, BMGN18, BHL^{+}19, CWHLT15, CGS14, DU15, EÖ17, Era15, EOS17, FM18, FGH^{+}16, FGHP17, FL15, FH17, Gan13, GKL19, HK19a, HQ12, HT18, HK17, JL13b, JSW18, KOZ16, Kre17, LA11, LJS18, Mas16, OS14, OPS15, QRSZ19, SV13, Say13, SH17, SW11b, ST18, ZZ15, vyKS14]. boundary-driven [FH17]. bounded [Che12, KR16, MOR17, Wen10]. bounds [AV11, BBB17, BCCH19, BKUV17, CDM^{+}18, CG14, DS16, FGL14, Har18, HHS15, LN15, PS16, SB13a, SB13b, XCW10]. Boussinesq [OS19]. box [PT10]. branches [GL11]. Bregman [SH19]. Brent [WG13]. Brinkman [AMORB16, GGM14, GKSS13, LDQ11]. Bubble [BS10, GN14]. Buchholz [LT10]. build [Maz11a]. bulk [BHLZ16]. bulk-surface [BHLZ16]. Burgers [LP15].

c [Dic12]. Cahn [BM11, CCS17, CHW17, CP14, DWWW17, ER15, GL11, GWW14, LCWW17]. calculation [Hor17b]. calculations [Eng14]. calculus [Chr11, EHRS12]. Camassa [ADM19]. canonical [EH12]. capillary [KP17]. capturing [HM14]. cardiac [Zam13]. Carlo [BSZ11, LRS20, BC15a, BD12b, GGK^{+}19, GKSS13, GKN^{+}15, HANT16, HMOU16, LRS19, PGW18, TSGU13]. case [ABER10]. Cauchy [BR11]. cbc [Dic12]. CCC [BR17]. cell [CCP13, Era15, GGGS19, YWY14]. cell-centered [CCP13, Era15]. cells [HJHA17]. centered [CCP13, Era15, ZZ15]. Central [CKL14]. Central-upwind [CKL14]. certain [Smy10]. CG [ZBJ18]. chain [MOS11]. chains [MS19]. changing [CC13, DCC18]. channel [BHT17]. characteristics [SWS16]. characterization [Kim19]. Chebyshev [Mas14, Maz11b, XCW10]. Chebyshevian [Maz11a]. chemistry [BCS14a]. circle [HJT14]. Circulant [GKN^{+}18]. clarifier [BKTT10]. clarifier-thickener [BKTT10]. class [BMN11, BB13, BBC^{+}18, CWHLT15, DD16, NvPZ10, Not19, RS11, Ven11]. classes [iTSMM10]. classical [AN18, BS13, CG17, LMS19, SX15a]. classification [JO10]. Clenshaw [HS15]. closed [BGN12, BEG14, KLL19]. clouds [DM14]. clusters [Gal15, LZ15]. co [DL11]. co-dimension [DL11]. coarse [EMR19, SDH^{+}14]. coefficient [HQ12, LS18]. coefficients [AY18, BD11, BSZ11, CC13, DCC18, GGK^{+}19, GKN^{+}15, GKN^{+}18, mHZZ17, Ise11, Li15, LMR16, SS18, SS16, TSGU13, TWZ17]. coercivity [CC13]. Collocation [ST18, AR16, LSW12, Lui17, MNT13, Nak17]. combined [SS15]. common [NNT15]. commutator [HK19b]. commutator-free [HK19b]. compactly [SV13]. compactness [HHX19]. companion [VVVF10]. comparison [BD12a, BHK10, BS19, LM15]. complementarity [GAB13]. Complete [HK19a]. completion [BGJ18, GK19a]. complex [BGHL14, Seg13, Zha16b]. complexes [CHH18]. complexity [Bel11]. component [Nou19]. Componentwise [NP15]. composites [PC13]. composition [CMMV14]. compressed [VD11]. compressible [AH17, CR12, FKM16, GMN19, Kar13, SS15, Wil19]. compressor [Str15]. computable [AV11]. Computation [MP15, Cho18, CCSS16, Fuk10, Fuk13, GL11, GST19, GP11, Ise11, JO10, Man13, Rie19, SG11, SB13b, WZH13]. Computational [CFL11, FHPS19]. computations [DK11, NP15]. Computing [ACWL14, NNT15, Seg13, ZLL^{+}12, BL13b, EMB10, GG12, JLL15, KLS19, MV15]. condensation [Loz19]. condition [CGHW11, DP16, HKK12, HK19b, JL13a, KdS18, KOZ16, Nak12, PZ18]. conditional [Fuk13]. conditions [AHR14, ABK13, BMGN18, DU15, GKL19, HK19a, HW11, LA11, LJS18, vyKS14]. conductivity [HKQ18, PC13]. cone [GS13, KdS18]. cone-constrained [GS13]. cones [Mir16]. confinement [GY12]. conformal [GT19, KPR14]. Conforming [Wac17, DO17, GN14, KK15b, SZ13, AT19, dVBMR16]. conjecture [HJZ14]. conjugate [EGLTP17, PT10]. connection [Jun15]. consecutive [Dri12]. conservation [AGS10, BBC^{+}18, BG16, BKTT10, GM14, HM14, MR17, Tow18]. conservative [EZ15, JP13]. Consistency [KSD10, Kla15]. consistent [ADF18, CEM14]. constant [AP11, BDHK12, FLZ19]. constants [OMS13, RHM13]. Constrained [Str15, AGS10, ABC15, CQR16, CHL13, GS13, GMT11, HS19a, NPVW19, PG17, SG11]. constraint [Owe16]. constraints [EMB10, FLMP12, HHYY14, NW18, OW11, PT10, RW12, dHQS15]. Constructing [BV12, HHS15, ZBJ18]. Construction [AT19, DGSY17, BD11, DL11]. contact [BK12, BGIS17, CMR18, DH17, HR10, KSD10, KSD11, Kla15, Rad19, WHC14]. contact-stabilized [KSD11, Kla15]. Continuation [BEK11, Bel11]. Continuous [HW18, LL12, BRK17, CC13, DLT13, Li15, YBTB14]. continuum [Cal17, LOSV16, MOS11, OLOV18]. control [AORS17, ABM18, BM11, CQR16, DD16, GY17, HSLZ19, HFL12, JSW18, KK15a, KSD11, LV13, NV12, NPVW19, OPS15, OW11, RW12, SG11, SA10, Tak15]. controllability [BHL11]. controlling [DS11, Dem16]. controls [CHW12]. convection [ABH13, BBFP18, BBK17, BS12, BC12, BHLZ19, CJ14, FKNP11, FH17, HJT14, KKR12, LHQ14, TWW15]. convection-diffusion [BBFP18, BBK17, BS12, BC12, CJ14, FKNP11, HJT14, KKR12]. convection-reaction-diffusion [ABH13]. Convergence [AG10, BBD16, BC15a, BMZ10, BDS13, CNSV17, CPB17, CLA19, CXH10, CDL19, CS12, DFL16, DS11, Dem17, DWWW17, DHKR16, EH18, FKM16, FLOP10, GW19, GS17, GH14b, GLW19, HS19a, HQ12, Har15, HLZ12, HJS15, HKT14, HSX12, HL18, KdS18, KS11a, KLLG17, KPR14, LS18, LZ15, Nak17, NTT16, PGW18, Pie18, PCCC18, SCHH13, Tow18, Ais15, AR16, BL13a, BT15, BHK10, BG12, BR17, BC17, Cal17, CvN13, DE16, ER18, FM18, FGH^{+}16, FGHP17, GGS15, Gan13, GLS17, GY17, HL19, HKK12, HY15, HH10, KL18b, NN11, NZ19, ÖS13, OTMS13, OYV17, SW19, Sid11a, Sin12, SH19, SH18, SX13, SX15a, SX15b, TOMS13, Wan18, ZRK16, dVCMR18]. Convergent [CR12, AKST14, ABAC16, CLM15, FLS10, GWW14, Kar13, KL17, KLL19, NKL21, Rie19, Sch16]. Convex [CLMS18, DKS13, HHR11, EÖ17, FN12, FLMP12, GNS15, GWW14, HHYY14, JZ14, JY16, LMSS17, Mir16, Wac17]. convex-nonconvex [LMSS17]. convexity [EMB10]. convolution [BBSV13, BLM11b, Hac11, LFS16]. coordinates [DS10]. Cordes [SS16]. Coriolis [CDKLM18]. corner [dVCMR18]. corners [FL15]. corrected [HS15, JSW18, Not16a]. Correcting [ABB19]. Correction [FKS19, LRS20, NKL21, OZ23, BBK17, Fer13a]. corrections [CCP13]. costs [FHPS19]. Coulomb [GMSS11]. counts [GMS19]. coupled [BDG^{+}18, BHLZ16, CDL19, DO17, LVY14, OS14, PS17b, SzS12]. Coupling [OB17, ADF18, AN18, AG10, BDQ10, BLS15, Era15, EOS17, GHS12, GMSS11, HJHA17, KL17, NKL21]. couplings [FFKP15]. Covariance [DHS17]. Crank [KK15a, Wan18]. critical [PSV10]. cross [BKV15, BC16, KK11, Mur17]. cross-diffusion [Mur17]. Crouzeix [HM16, LMR16]. crystals [GG12]. cubature [Che12]. cubic [ABB15]. cuboid [CCQ13]. cuboidal [AT19]. curl [KK15b, Sun16]. Curtis [HS15]. curvature [BGN19, KLL19, ZLL^{+}12]. Curve [PS17b]. curves [BGN12, EKAB16]. curvilinear [ABB15]. customized [HHYY14]. Cut [BHLZ16, YBTB14]. CutFEM [BHLZ19]. cylindrical [AP10]. d [TWW15, CCZ13, EMR19, FR19, HR19, HOR17a, LCH12, RS11]. DAEs [SB13a, SB13b]. Damped [EK18]. damping [CGK18, PSV10]. Darcy [BDQ10, BB12, BDG^{+}18, BGG^{+}16, BOS18, CGHW11, CGSW16, CS17, CHW17, DO17, GVY14, LDQ11, LVY14]. Data [BO18, AdST12, BGJ18, BMZ10, CD11, GSY17, HR12, HSW17, HW13, LC14, WE16, ZS14]. Davidson [Ais15]. DDFV [GKL19]. de-aliasing [BT15]. DE-Sinc [OTMS13, TOMS13]. Decay [KS11c, Pet12]. decomposition [AL17, ABAC16, Ber16, CGHW11, CH12, EH18, FN12, FLS10, HSZ13, NGKN10, PT12, SW11b, YWY14, Zha18, ZDQ17]. decoupled [CHW17, GGGS11]. defined [NS14]. degenerate [ABH13, CJ14, DE16, EH18, EKAB16, KKR12, LS18, Rei12, SX15a, SX15b]. degree [CHTV18, Dri12, HLS11]. delamination [OB17]. delay [HV12, WMS10]. denoising [AHR14]. densities [Lin17]. density [SSS12]. dependent [Ber16, HL16, SG… 12, SS16]. derivative [AdST12, Nov14]. Derivatives [BCPS10, ADF18, Dal10]. derived [BC12]. descent [dHQS15]. description [AGK^{+}14]. Design [BDF12, ADF18, Maz11a]. designs [CFL11, GP11]. determinant [SAI17]. deterministic [Rus10, TWZ17]. DeTurck [Fri15]. developmental [GT14]. DG [BS19, DGS15, Kar13]. diagonal [Har15]. diagonalization [GW19, Str15]. diagonally [DK11]. difference [AL15, BY19, DHKR16, Fro18, GK19b, HR14, HV12, KKR12, LVY14, MR17, Pru14, SzS12]. different [Dri12, HHNS16]. differentiable [DGSY17]. differential [ABB15, AR12, BLF19, BM13, Bör10, BV14, CMMV14, CH18a, CS12, DM14, Git14, GKL15, GLW19, HHB16, HHX19, HV12, Mas16, Rus10, SV12, SA10, Sti10, WZ10, WMS10]. differential-algebraic [BM13]. Differentiation [dGKL19, BMWB14, DS16, DS18a]. diffuse [DS18c]. diffusion [AH14, AH17, AV11, ABH13, BO19, BBFP18, BBK17, BGGH16, BS12, BC12, BCCH19, BEJ14, CLLS19, CCP13, CJ14, CM15, DS18c, DK16a, FKNP11, FGL14, FH17, GV19, GGGS11, HPS16, HJT14, KKR12, KS11b, Kop17, KLLG17, LS18, LHQ14, Mur17, Mus15, Pie18, PS17b, SG… 12]. diffusion-convection-reaction [LHQ14]. diffusive [TWW15]. digital [HMOU16]. dilation [CCR14]. dimension [DL11, ZS14]. dimensional [ABB19, ABER10, ABK13, BGG^{+}16, CCQ13, CH19, FGP10a, GL11, HS18, Pet12, Wan12]. dimensionality [Pet12]. dimensionally [BDQ11]. dimensionally-heterogeneous [BDQ11]. dimensions [BGGH16, DGS15, KS18a]. Dirac [AORS17]. Direct [Man13]. direction [HY15]. directional [BM17, Dal10]. Dirichlet [BY19, FL15, GMS19, JSW18, LN18, OPS15, vyKS14]. discontinous [BMZ10]. discontinuities [SKW19, Tow18]. discontinuity [DL11]. Discontinuous [AH14, AP10, HM13, HK17, LVY14, MR17, SS16, WHC14, YS13, BB12, BCS14b, BC15b, BCGS11, BKTT10, BS10, CP18, CBHW18, CM15, DM16, DK16b, FKNP11, HM14, HW18, HKXZ16, JL13b, KP17, LS18, LMMR19, LP15, Loz19, MP18b, MNT13, Mus15, RSV13, SSS12, SA10, SH18, Wil19, ZS14, Zha16a]. Discrete [DGBL^{+}15, GEEF17, JLZ18, KS18b, LV17, ABM18, BFG13, BIM11, BKN13, CLLS19, DFL16, DHM19, EZ15, GLS17, HR12, HHX19, HFK19, HSZ13, KK11, KL17, LA11, LM13, MOGO17, Mir16, NKL21, Pie18, PZ18, SW19, ZLL^{+}12, Zha16b, ZS17, dGKL19]. Discrete-time [GEEF17]. Discretising [LS17]. Discretization [BCMO16, BGN19, BGG^{+}16, CC15, DGS15, DK16b, FGP10a, Har18, HR19, HS18, KdS18, Kla15, KL18a, KPR14, LMR16, Man15a, NV12, NW18, RE14, RSV13, SV16, dVM11]. discretizations [BRK17, CCS17, FS17, HKXZ16, KL18b, LR17, Sch16]. discretized [BHL11, CPS16, FS19, HH11, PPS18]. Discretizing [Chá11]. dislocation [CCM12]. dispersion [JM18]. dispersion-managed [JM18]. dispersive [AY18]. displacement [CLLS19, CDL19, Fer13a, LM15]. displacement-correction [Fer13a]. dissipative [ADBN16, BEJ14, CMP13, Egg19]. dissipativity [GAB13]. dissolution [KPR14]. distorted [WXY12]. distributional [FY11]. div [BPZ12]. divided [MR17]. domain [AL17, ABAC16, Ber16, BG17, BGHL14, CGHW11, CH12, EH18, FLS10, HPS16, NS14, PT12, QRSZ19, Say13, SW11b, YWY14]. domains [AP10, FM18, FKM16, FL15, HPS16, LN18, Wen10]. dominant [DK11]. Dörfler [KS11c]. double [VVVF10, iTSMM10, TWW15]. double-diffusive [TWW15]. doublets [BLM18]. doubling [NP15, XL17]. Douglas [HHYY14, HY15]. DP [DGS15]. DPG [BDGQ12]. drift [BCCH19, LS18]. driven [FH17, GLW19, KLLG17]. dual [ABAC16]. duplication [Fuk13]. dyadic [BCD15]. dynamic [BGR18, KP17]. dynamical [Chá11, FS17, GAB13, KSD10]. dynamics [BGN16, CCM12, DGBL^{+}15, FS17, NDL14, TWW15, Wor19, ZS12].

early [MOGO17]. early-exercise [MOGO17]. Edge [BBK17]. Edge-based [BBK17]. effect [BBL11]. Effective [CM13, CCM12]. effects [KP17, MdC17]. Efficient [BLF19, HHB16, Mir14, RSV13, SSS16, CGSW16, EH12, Hac11, KVW15, Wan12]. Effortless [SM16]. eigendata [BSCZ12]. eigenpair [JLL15]. eigenproblem [LMMR19]. eigenproblems [CEM14, SDH^{+}14, SX15a, SX15b]. eigenvalue [Ais15, BSCZ12, BV17, BEK11, CGMM11, CG11, CGMM14, CG14, DBV11, DP16, GS13, Gal15, GGMO16, GGK^{+}19, HL19, JMM12, Kim19, LLW10, LZ15, LLY15, Men11, Nak12, Sun16, SX13]. eigenvalues [BBB17, BAK19, BEK11, CDM^{+}18, FPD12, Fro18, GHP17, HHS15, IN18, Koe19, MP15, SX15a, SX15b]. eigenvectors [CDM^{+}18]. eikonal [DES11]. elastic [BGN12, CMR18]. elasticity [BGIS17, BDGQ12, DL15, FFKP15, GWX19, GN14, HKXZ16, LMMR19, LMcS17, Smy10]. elastoplastic [CLA19]. elastoplasticity [CSW16]. elastostatics [HR10]. electric [BN16, HHR11]. electrical [GS17, HJHA17, HKQ18, HM18]. electromagnetic [CCZ13]. electronic [KY12, SY17]. electrorheological [BBD16]. element [AV12, AH14, AORS17, AKST14, AP11, ADM19, ABB15, BBL11, BPS10, BY14, BBK17, Bar13b, BHK10, BGJ18, BDG^{+}18, BBD16, BPZ12, BHLZ16, BO18, BHL^{+}19, CGPS17, CGH13, CGMM14, CPS16, CSW16, CPB17, CWHLT15, CXH10, CW12, CH18a, CHH18, CCSS16, DS11, Dem16, Dem17, DCC18, DWWW17, DKS13, DU15, DO17, DSX16, DLT13, EÖ17, ER18, ER15, Era15, EOS17, FM18, FGH^{+}16, FGHP17, Gan13, GT14, GGM14, GG12, GVY14, GKSS13, GY17, GKN^{+}15, GGRG15, GLW19, HKK12, HT18, HR10, HLZ12, HSX12, HHS15, HMX17, HH11, JP13, KOZ16, KK15a, KS18b, Kir11, KT12, KK15c, KLL19, LM11, LMS19, LV17, LA11, LMY18, LLY15, LCWW17, LMR16, LL10, LHQ14, Mak18, MSW13, Mat13, Nei10, NV12, OS14, OPS15, OR10, PS13, PC13, QRB11]. element [SS15, SG… 12, SZ13, SS16, SW11b, Ste14, Ste10, Wac17, WX12, WXY12, ZBB14, Zha16b, ZRK16, dVBMR16, dVMRR17, dVCMR18, BSZ11]. elementary [EHRS12]. elements [AT19, BRK17, CCQ13, CGS14, CK16, CGK18, DLT12, EH13, GMÖ^{+}18, GN14, HHNS16, HK17, HM16, KS18a, KK15b, KLLG17, KS11c, LL12, LMcS17, Mir12, Mon19, dVMM19]. ELLAM [CDL19]. elliptic [AV12, ABM18, BSZ11, Bör10, BS10, CDNP16, CCQ13, DK16a, DGS15, DK16b, DLT13, Fuk10, Fuk13, GV15, GGK^{+}19, Git14, GY17, GKN^{+}15, GKN^{+}18, HQ12, HPS16, mHZZ17, HSLZ19, HS19b, HLZ12, HHS15, JL13b, KS18a, LMR16, NW18, NTT16, OPS15, PT14, Smy10, Sti10, TSGU13, XZ10, ZZ15]. ellipticity [OS14]. embedded [DFL16, HHB16]. embedding [GKN^{+}18]. enclosure [KKN14]. Energy [JSW18, Kop17, Say13, CHW17, Dem16, FLMP12, LMY18, OPS15, ZS17, ZVKG11]. Energy-norm [Kop17]. enhancement [MR17]. enrichment [Sch11b]. entries [BV12]. Entropy [BEJ14, HM14, FH17, Wil19, ZS12]. entropy-dissipative [BEJ14]. Entropy-stable [BEJ14]. epitaxy [CW12]. equation [AY18, AKST14, ABH13, ABK13, ADM19, AP10, AN18, BFG13, BLS15, BO19, BD12a, BCST19, BY19, BDG^{+}18, BN16, BMGN18, BDS13, BDF12, BEJ14, BO18, Cal17, CG14, CP14, Cho18, CCSS16, CMP13, CR12, DES11, DS18c, DHKR16, ER18, ER15, FS14, FL15, GL11, GGS15, GMÖ^{+}18, GH14b, GWW14, HK19a, HT18, HS17, JM18, KK15a, KS11a, Kre17, LL12, LP15, MP18b, MNT13, Nei10, NZ19, OS14, OS19, PSV10, PT12, Pie18, QRSZ19, SW11b, VGG14, ZS14]. equations [AL15, AL17, AR16, AR12, BLF19, BZ17, BBK17, BM11, BGGH16, BB13, BBKS18, BDLM18, BC12, BMP10, BMZ10, Bör10, BGHL14, BHL11, BV14, BC17, CCP13, CG17, CP18, CHW12, CCFM17, CC19, CMMV14, CCLM15, CLM17, CQR16, CKL14, CDKLM18, CH12, CJ14, CS12, CLM15, DD16, DU15, DL15, DE16, DLT12, EH18, FGHP17, FH17, GY12, GMN19, GLS17, Git14, GK19b, GKL15, GLW19, HHNS16, HHB16, HR19, He13, HOR17a, HJS15, HP17, HPS18, HKT14, HJT14, HKXZ16, HJNS17, HS18, HV12, JLZ18, KKR12, Kar13, KOZ16, KS18b, KKN14, KK15c, KL17, KL18b, KNWW19, LA11, LCH12, LC14, Li15, LS16, LL10, MP18a, Man15a, Mas16, MOR17, Nak17, NvPZ10, PS13, Pet12, PT14, PCCC18, Pru14, Ren17, RS11, RE14, Rus10, SV13, Say13]. equations [Sch10, SH17, SSS16, SWS16, SS16, SV12, ST18, SA10, Sti10, TT15, Ver17, Ver10, WZ10, Wan12, Wil19, WMS10, XZ10, XXL12b, XXL12a, XL17, YS13, ZT12, ZS12, NKL21]. equidistant [BDHK12]. equilibration [CM13]. equilibria [GL11]. equivariant [MMMKV16]. ergodic [FS19, HY15]. Error [ADM19, ABC15, BCS14a, BM11, BRK17, BL13b, CCS17, CDNP16, CD11, CC15, DS16, DD16, FS19, GMN19, GK19b, GGGS11, GSV19, HS15, HP17, HPS18, LCWW17, LL10, OMS13, SZ13, XCW10, ZS14, AV12, AV11, AMN11, AMORB16, BLM11a, BB12, BDLM18, BKN13, BM17, BKUV17, CCM12, CGPS17, CM13, CGH13, CPS16, CGG19, Che12, CS17, CM18, CJ14, DS18c, DK16a, DWWW17, DSX16, FLOP10, GMS19, GGMO16, GM14, GMT11, HR12, HSV12, Har18, HZZ18, KKR12, KK15a, Kla15, Kop17, KVW15, LM11, LMS19, LLW10, LO14, LMY18, LN18, LM13, LJS18, MOS11, Mon19, NV12, NW18, NvPZ10, OW11, PS16, RW12, RHM13, ST16, SWS16, TT15, WX12, Wei11, YS13, ZS17, ZVKG11, dVMM19]. errors [ABB19, DS11, Dem16, DLPV18, FGH^{+}16, MdC17, PSV18]. estimate [HS15, KKR12, LL10, Mon19, Wei11]. estimates [AV12, AMN11, ADM19, BCS14a, CCM12, CCS17, CGPS17, CGH13, CD11, CC15, CHTV18, Che12, CM18, CJ14, DD16, DK16a, DWWW17, FS19, GMS19, GMN19, GGMO16, GNS15, GK19b, GGGS11, Hac16, HZZ18, KKN14, Kop17, KSZ13, LM11, LMY18, LM13, Mak18, MR17, NV12, NW18, OMS13, OW11, RW12, Say13, SWS16, Ste15, YS13, ZS14, ZS17, ZVKG11, dVBRS11]. Estimating [PSV18, Che16b]. estimation [DSX16, Hac13, HSV12, LMS19, LO14, Lin17, NPVW19, PGW18, RHM13, dVMM19]. estimator [CS17, Era15, KVW15]. estimators [BC15a, CM13, FLOP10, SAI17, WX12]. Euclidean [Man15b]. Euler [CvN13, GGGS11, He13, HP17, JT16, LS18, Pin18, Ren17, WZ10, Wil19]. Eulerian [NDL14]. evaluation [Bal12, SS11]. evanescent [AN18]. EVD [OYV17]. even [BCS14a, HLS11, OTMS13, TOMS13]. even-tempered [BCS14a]. evolution [CLA19, CLM17, Egg19, HP17, HPS18, JLZ18, KL18a, KL18b]. Evolving [ER15, GM14, KLLG17, KLL19, Man15a]. exact [DGBL^{+}15, Koe19]. exactly [HMX17, MMMKV16]. examples [GV19]. exercise [MOGO17]. Existence [BFG13, CFL11]. expanding [Wor19]. expansion [SH17]. expansions [LT10, MOGO17, SS11, WZH13]. Explicit [EH13, AL15, GMS19, He13, OMS13, SB13a, SB13b, ZS14]. Exponential [Pet12, dVCMR18, BM19, CCR14, DMN12, EK18, GV19, HS17, HK19b, KLS19, OS19, Pie18, iTSMM10, WMS10]. exponential-fitting [GV19]. exponential-type [HS17, OS19]. exponents [BL13b, BV14]. Extended [SCHH13, KS11a, MV15, dCM17, Maz11b]. extensible [Owe16]. Extension [AdST12]. exterior [ABK13, BBSV13, BLS15, KL17, NKL21]. Extrapolated [LM15]. extremal [LLY15].

factor [NN11]. factoring [BH11]. factorization [DK11, LGC14, RB11]. family [AHR14, BBC^{+}18, CDL19, OTMS13, Ste10, TOMS13]. Far [dCMT16]. Far-field [dCMT16]. Farin [Mat13]. Fast [Bal12, FKS18, Fuk10, GST19, Kir11, KT12, Kir17, LGC14, PG17, SS11, Fuk13, GP11, HT18, Ise11, MMMS10, Mir14, Run14, WZH13, FKS19]. Faster [DMMS11]. fat [BIM11]. FE [GMSS11]. Fejér [Not16a]. FEM [BS19, BC17, CLLS19, CG11, CPR13, CRS16, CGG19, EPS17, FMP15, FFKP15, Gal15, GHS12, GGGS11, JSW18, Kar13, Sun16, Wei11]. FEMs [CDNP16, GLS17, mHZZ17, Li15, SWS16, Wan18]. Fer [RI15]. FETI [DGS15, PS11]. Feynman [FS19]. Fickian [BBFP18]. fictitious [BG17, JL13b]. field [BN16, BMO15, CCS17, DL11, FKS18, FKS19, dCMT16]. fields [Ber16, DHS17]. Filippov [DL11]. film [CW12, FLMP12]. filtering [NGKN10]. finding [LGL17, WG13]. fine [PZ18]. Finite [AGS10, AP11, ABB15, BPS10, Bar13b, BSZ11, BDG^{+}18, BHL^{+}19, LLY15, PC13, SG… 12, SzS12, ZS17, AV12, AH14, AORS17, AKST14, ABH13, ADM19, AT19, BBL11, BY14, BBK17, BHK10, BRK17, BY19, BGJ18, BBD16, BC12, BPZ12, BDF12, BHLZ16, BO18, CCP13, CGH13, CGMM14, CPS16, CSW16, CPB17, CXH10, CW12, CH18a, CHH18, CCSS16, CK16, CGK18, DS11, Dem16, Dem17, DCC18, DWWW17, DKS13, DU15, DO17, DSX16, DLT13, DHKR16, ER18, ER15, Era15, EOS17, EH13, FGL14, FH17, Fro18, GT14, GGM14, GG12, GM14, GVY14, GKSS13, GY17, GKN^{+}15, GK19b, GGRG15, GLW19, GN14, HR14, HKK12, HZZ18, HR10, HLZ12, HW18, HHS15, HMX17, JP13, KKR12, KOZ16, KK15a, KS18a, KS18b, KK15b, Kir11, KT12, KK15c, KLLG17]. finite [KLL19, KS11c, LM11, LL12, LMS19, LV17, LA11, LCH12, LC14, LMY18, LVY14, LCWW17, LMR16, LL10, LHQ14, Mak18, Mir12, Mur17, Nei10, NV12, OS14, OPS15, OR10, Pet12, QRB11, SS15, SS18, SZ13, SS16, Ste10, Wac17, WX12, WXY12, XZ10, ZBB14, ZZ15, Zha16b, ZRK16, BHT17, BGHL14, FGP10a]. finite-difference [HR14]. Finite-dimensional [FGP10a]. finite-element [GGRG15]. Finsler [Mir14]. First [BHK10, NS14, BS12, Fuk10, MP18b, RS18]. First-order [BHK10, MP18b]. fitting [GV19, IN18, OTMS13, TOMS13]. fixed [EHRS12, HRS12]. Fliess [GEEF17]. Floater [dCM17]. flow [BGN12, BB12, BG16, BMO15, CRS16, CS17, DD16, DHM19, Fri15, GGM14, KP17, KLL19, KLS19, LDQ11, PS17b, SS15, YBTB14, ZLL^{+}12]. flows [AH17, BPS10, BGN19, BHT17, BCN19, BEG14, BGG^{+}16, CCS17, CHW17, GVY14, LVY14, RHM13]. fluid [AKYZ18, AG10, BJLZ15, BG17, DD16, Fer13a, GV15, LDQ11, NP15, SSS12]. fluid-poroelastic [AKYZ18]. fluid-rigid [SSS12]. fluid-structure [AG10, BG17, Fer13a, GV15]. fluidic [BGN16]. fluids [BBD16]. flux [BBK17, BHK10, BG16, BKTT10, DSX16, PSV18, Tow18, WXY12]. flux-correction [BBK17]. fluxes [CCSS16]. Force [OLOV18, MOS11]. Force-based [OLOV18, MOS11]. forces [CDKLM18]. forcing [RE14]. form [FGP10a, FGP10b, Har15]. format [EH12, GK19a]. formation [GT14]. formats [Nou19]. forms [ABB15, CH18a, Dem17, HHX19]. formula [Not16a, Pin18]. formulae [AL15, Not19, PS16]. formulas [DS18a, iTSMM10]. formulation [AMN11, BB12, DGBL^{+}15, EGH15, HJHA17, HFL12, MP18b]. formulations [BKN13]. Fortin [MSW13]. forward [RS11]. foundation [Smy10]. four [HK19b]. Fourier [Adc10, BZ17, BT15, CG17, GP11, KK11, NGKN10]. fourth [BMN11, BEJ14, CGH13, CCQ13, HH11, Zha18]. fourth-order [BEJ14, CGH13]. fractional [AN19, AdST12, BLF19, BO19, BLP19, CJ14, CM15, JLZ18, KY12, Mus15, Nov14]. fracture [DHM19]. fractured [BGG^{+}16, CS17, LDQ11]. framework [CDM^{+}18, HSV12, Man15b, Mas16, Olv12, YBTB14, Zha18]. free [BDGQ12, CG11, HK19b, Kre17, SAI17]. frequencies [CLM17]. Frequency [EÖ17, ABER10, BM17, CWHLT15, HL16, NGKN10, NDL14]. Frequency-adapted [EÖ17]. Fresnel [ACWL14]. friction [GMSS11]. frictional [BK12, BGIS17, CMR18, Rad19]. Froissart [BLM18]. front [BCS14b]. full [DGS15, KL18a, BEG14]. Fully [AV11, AL15, BFG13, BIM11, CLLS19, CDL19, GLS17, HR12, KL17, Nak17, Nei10, NKL21]. Function [iTSMM10, ABB19, FY11, HSW17, Man13, SG11, SV12]. functional [BV14, BMO15, EGLTP17, Mas16, Rus10, WZ10, YWY14, ZVKG11]. functionals [BBL11, BCMO16, Che12, DKS13, HW13, SS11]. functions [BG12, BLM18, BGR18, BAK19, BS17, BPZ12, CHW12, Che12, DV11, DBV11, DGSY17, Fro18, Fuk13, GH14a, GT19, GN14, HMOU16, Hor17b, KY12, KR16, LT10, MMMS10, Maz11b, Mir16, NNT15, NTT16, Not19, OTMS13, PS16, PT14, PR13, SV13, SY17, Seg13, TOMS13, Wac17, WG13]. Further [TSGU13].

Gaffney [HHX19]. gain [BD11]. Galerkin [AH14, AMN11, ADM19, AP10, AR12, BBSV13, BB12, BS12, BCS14b, BC15b, BCGS11, BS10, CP18, CBHW18, CC13, CC10a, CM15, DM16, DK16b, EÖ17, EPS17, EGH15, FKNP11, FLOP10, Fer13b, GLS17, HM13, HK17, HM14, HW18, HKXZ16, JL13b, KP17, KNWW19, LMMR19, LV17, LMY18, LVY14, LP15, Loz19, MR17, MP18b, Mus15, RSV13, RMS17, SSS12, SV13, Say13, SS16, Ste15, Sti10, SH18, TT15, WHC14, Wan18, Wil19, Wor19, YS13, ZS14, Zha16a]. gamma [AHR14]. gaps [GG12]. gas [NDL14, ZS12]. Gasca [HJZ14]. gauge [HS13]. Gauss [BCD15, AR19, GST19, JZ13, Kal15, Man15a, PS16, SY17, dS18b]. Gaussian [BCS14a, DMN12]. Gegenbauer [Dri12]. General [DMG19, ABH13, CCR14, DKS13, FKM16, GWX16, HW13, JY16, Loz19, Olv12, Sid11a, SX13, WS19]. generalising [Man15b]. Generalized [AR19, AGK^{+}14, BCN19, CH18a, HHX19, HM18, LFS16, MS19, ABC15, BJLZ15, BB13, Chá11, DBV11, DLT12, FN12, FY11, GV15, GS16, IG15, Nak12, PSV10, RMS17, SSS16, SDH^{+}14, ZBB14]. generalized- [ABC15]. generation [FKS18, FKS19]. generic [CCP13]. Geometric [BR11, GM14, CM10, Maz11a]. geometrically [DGBL^{+}15]. geometries [RHM13]. geometry [CHW17, Dic12]. Gilbert [AKST14]. Global [Ais15, CK16, CGK18, EMB10, FGL14]. globally [PZ18]. GMRES [GGS15, Olv10]. GMWB [HFL12]. Godunov [Tow18]. Golub [Bar13a]. good [Maz11a, OS19]. Gordon [BZ17, BD12a, CCLM15, FS14]. grade [BJLZ15]. graded [BN16, BCGS11, CNX12]. Gradient [BGG^{+}16, DL15, CPS16, EGLTP17, LMcS17, OW11, PT10, ZWJ19]. gradients [BGR18, CKL14]. Gram [BS13]. Grassmann [AV10]. Greedy [HSLZ19]. green [SZ13, BC16, FY11]. Gregory [FR19, JT16]. Gregory-like [FR19]. grid [Che16a, HT18, NTT16]. grids [BHK10, BCD15, CNX12, GWX16, LVY14, RW17, Zha16b]. Gross [AL17, HKT14]. Grothendieck [FLZ19]. ground [BFG13]. group [ABC15, DGBL^{+}15, HHNS16]. Guaranteed [CDM^{+}18, CG14, GGS15]. Gummel [BC12].

H [Bör10, Bör10]. H- [Bör10]. half [Fuk10]. Halley [Kal15]. Hamilton [BMZ10, Cal17, CLM15, GK19b, SS16]. Hamiltonian [BCM17, CCM12, CCFM17, CHL13, FGP10a, FGP10b, GKL15, HL16, KLS19]. hanging [SW11a]. Hardy [HHNS16]. harmonic [BPS10, CCZ13, Dem17, GH14a, HHNS16]. heat [BPS10, BDG^{+}18, BDF12, BO18, HT18, QRSZ19, SzS12]. Hele [LCWW17, GGGS19]. Helmholtz [BKV15, BM17, GMS19, GGS15, HK19a, HSZ13, SW11b, Wan18]. hemivariational [HSD18]. Herman [LS17]. Hermite [BCS14a, CH18b, CM10, GST19, MS19, SY17]. Hermitian [BV17, MV15, OYV17]. Hessian [Fro18]. heterogeneous [AH14, BDQ11, LMR16]. heuristic [Jin17]. hexahedra [AT19, WXY12]. hexahedral [EH13]. Hierarchical [ZVKG11, EPS17, LO14, Mir16, SM16]. High [CCFM17, ABB19, AHR14, ABER10, ADBN16, BLF19, BPZ12, BM17, BDS13, CWHLT15, DL11, HL16, mHZZ17, HZZ18, HKT14, JL13b, KNWW19, NDL14, Ren17, RMS17, Run14, WS19, ZS12]. high-dimensional [ABB19]. high-frequency [ABER10, BM17]. High-order [CCFM17, AHR14, BDS13, HZZ18, HKT14, KNWW19, Ren17, WS19]. Higher [Nou19, SH19, ZBB14, BKN13, BGHL14, HU17, HK19b, HL18, SKW19]. Higher-order [Nou19, HU17]. Highly [CLM17, XL17, CMMV14, CCLM15]. Highly-oscillatory [CLM17]. Hilbert [JT11, NTT16, Olv12, WZH13]. Hilbert-space-valued [NTT16]. Hilliard [CHW17, DWWW17, LCWW17, BM11, CCS17, CP14, ER15, GL11, GWW14]. Hölder [CvN13]. Holm [ADM19]. homogeneous [FHN^{+}14]. homogenization [AH17, AY18]. homotopy [CGMM11]. Hood [MSW13]. Hopf [Chá11]. horizontal [CKL14]. Hörmander [CGS14]. Hormann [dCM17]. hp [BGIS17, OB17]. hp-BEM [BGIS17, OB17]. hull [DKS13]. Hybrid [LRS19, BL13b, BGG^{+}16, BKUV17, LSW11, LOSV16, LRS20]. hybridizable [CM15, SH18]. hybridized [GWX19]. hydrostatic [GGRG15]. hyperbolic [ADBN16, BDF12, CMP13, HP17, HPS18, KK11, MR17, YS13, ZS14]. hyperelastic [CR12]. hypersingular [CH12, PT12]. hypotheses [GHS12]. ideal [BKW10, WS19]. identification [HQ12, JZ13]. Identifying [HKQ18]. IGA [BS19, FGHP17]. II [ABER10, BKW10, CG17, FGP10b, Not19, PS11, SX15b]. Il’in [GV19]. ill [CN19, HH10, Kal15]. ill-posed [CN19, HH10, Kal15]. image [CLMS18]. imaginary [AL17]. imaginary-time [AL17]. impedance [GS17, HHR11, HKQ18, HM18]. implementation [BKW10]. implementations [Rie19]. Implicit [VVVF10, AL15, AG10, CvN13, FGL14, GMN19, He13, HJHA17, HP17, HPS18, KL18a, SS18, WZ10]. implicit-explicit [AL15]. implicit-linear [CvN13]. implicit/explicit [He13]. improved [FR19, LJS18]. Improving [FS17]. incomplete [Fuk10, RB11]. incompressible [BCN19, CCS17, Fer13a]. Incremental [Fer13a]. indefinite [CDNP16, CXH10, DLT12, OMS13]. indefiniteness [HN16]. independent [GGS15]. index [HANT16, SB13a, SB13b]. index-one [SB13a, SB13b]. inequalities [AdST12, BK12, EH13, HSD18, NvPZ10, RW17, Ste14]. inequality [BL13a, HS19a, HHX19, NW18]. Inexact [JT11, CHTV18, FHPS19, JLL15, Jin12, SG11, SX13, ZBJ18]. Inf [HW11, RHM13]. Inf-sup [HW11, RHM13]. infinite [BM19, CH19, HHNS16, Man13]. infinite-dimensional [CH19]. infinitely [DGSY17]. inhomogeneous [LLY15]. initial [BMZ10, GW19, HR12, HW18, ZS14]. initial-value [GW19]. integral [AR16, BY19, BN16, BLP19, BC16, BDS13, CH12, FGHP17, FL15, Fuk10, HJHA17, Kre17, NvPZ10, PT12, QRSZ19, SV13, Say13, SH17, ST18, VGG14]. integral-equation [BDS13]. integrals [ACWL14, Bal12, BLF19, Olv10, Pin18]. integrands [DNP14]. integration [ABC15, BBL11, BM13, BD12b, HS19b, HMOU16, MMMS10, OMS13, Pla15, Rus10, iTSMM10]. integrator [BZ17, HS17]. integrators [EH18, HL15, HK19b, OS19, Ver17]. integro [GLW19]. integro-differential [GLW19]. interaction [AKYZ18, AG10, Fer13a, GV15]. interactions [BG17]. interconnection [HU17]. interface [CGHW11, DS18c, DLT13, HLZ12, LR17, PS11, Run14]. Interior [LMY18, LM13, BLS15, BNRS17, KP17, KS11b, KL17, NKL21, PG17, SG11]. interior-layer [KS11b]. interlaced [DGSY17]. Interlacing [Dri12]. interpolants [JT16, dCM17]. Interpolation [GH14a, IG15, BLM11a, BY15, BO17, BM17, BDHK12, BRZ13, CH18b, EKAB16, Mas14, MdC17, Mon19, NOS16, SM16]. interpretation [BR11]. intersection [dCMT16]. Interval [SB13a, SB13b]. Intrinsic [CCSS16, HAG17, Sor10]. invariant [BEK11, Rie19]. inverse [BSCZ12, BH11, DES11, GS13, HL19, Jin10, Jin17, KKN14, LO14, Man13, SH19, ZWJ19, dHQS15]. inverses [FMP15]. inversion [Kir17]. invert [ST16]. inviscid [BGHL14]. involving [AORS17, BHT17, BCMO16, YS13]. IRGNM [KdS18]. irreducible [JLL15, LGL17]. isentropic [Ren17]. Isogeometric [ZDQ17, EH13, GMP^{+}14, dVBRS11]. isometries [Bar13b]. isoparametric [Mon19]. isotropic [BGN12]. issues [BB13]. iterated [JZ14, Man13, SH19]. iteration [AV10, BL13a, CN19, CGS14, GMS19, JLL15, LGL17, dHQS15]. iterations [ABER10, HL19, Jin12]. Iterative [HFL12, AHR14, BV17, CHTV18, CH18b, GST19, He13, Man15b]. Iteratively [HW13, JZ13, Kal15]. Ivanov [KdS18].

Jacobi [BMZ10, Ais15, Cal17, CLM15, GK19b, Har15, Man13, OYV17, SS16]. Jacobi-Bellman [SS16]. Jacobi-type [Har15]. Jacobian [Fuk13]. joint [Str15]. Jordan [Koe19]. Joseph [CGHW11]. junction [CLM15, GK19b]. Justification [CGG16].

Kac [FS19]. Kahan [Bar13a]. Kármán [BNRS17]. karstic [CHW17]. Kato [KL18b]. KdV [HS17]. Keller [ZS17]. Kernel [FHN^{+}14, BM17, DS16, NTW18]. kernel-based [DS16, NTW18]. kernels [Che12, FY11]. kind [BK12, Fuk10, Mas14, Not16a]. Kinetic [BEG14, GV19]. Kirchhoff [LL12, SH18]. Klein [BD12a, BZ17, CCLM15, FS14]. Kluk [LS17]. Koiter [Zha16a]. Korteweg [PSV10]. Kronecker [BS17, DLPV18]. Kronrod [dS18b]. Krylov [KS11a, MV15]. Kummer [LT10]. Kutta [AMN11, BBSV13, BLM11b, GS16, HPS18, LFS16, Man15a, Ver10, WE16, ZS14].

labeling [YBTB14]. Lagrange [AKYZ18, BS12, BG17, EKAB16, Fer13b, HR10, Jun15, Ren17, SSS12]. Lagrange-projection [Ren17]. Lagrangian [Fer13b, Jin17, RS11]. Laguerre [GST19]. Lamé [BLS14]. Lanczos [Ais15, Bar13a, LZ15]. Landau [AKST14]. Landweber [CN19]. Langevin [FS17]. Laplace [BHL^{+}19, CDM^{+}18, vyKS14]. Laplacian [BLP19, DS10, GGS15]. Large [Pru14, AH17, Bar13b, BGR18, BBKS18, BS17, JLL15, LC14, Lin17, Ren17]. largest [GGS15]. lateral [PS17b]. Lattice [DNP14, GSY19, BD11, BFK^{+}16, BKUV17, Dic12, DGSY17]. Lavrentiev [BGJ18]. Lavrentiev-finite [BGJ18]. law [BG16, Tow18]. laws [AGS10, BBC^{+}18, BKTT10, GM14, HM14, MR17]. layer [CCZ13, KS11b]. layers [BHT17]. LDU [DK11]. least [BGR18, BC17, CPB17, DMMS11, IN18, JL13a, RS18]. least-squares [BGR18, BC17, IN18]. Lebesgue [BDHK12]. Legendre [BCD15, Ise11, Lui17]. lemniscates [HN16]. level [BK12, BSZ11, GT14, GKSS13, LCH12, Tu11, dHQS15]. Levenberg [BGR18, HH10, Jin10, JY16]. lexicographic [HHB16]. Lie [ABC15, BCM17, DGBL^{+}15]. Lifschitz [AKST14]. like [EGLTP17, FR19, GW19, GGGS19, Ren17, SX15a, SX15b]. limit [AN18, BD12a, BZ17, Cal17, CG17, FS14, GV19]. linear [AL15, AL17, AP10, BBL11, BGIS17, BM13, BDGQ12, CDNP16, CHW12, Che12, CvN13, DL15, DE16, GWX19, GAB13, GGGS11, HHYY14, HHB16, HKXZ16, JMM12, KK15a, KKN14, KK15c, KL18b, LL12, LMcS17, LM15, LHQ14, Mur17, Sid11a, Sin12, Smy10, XZ10, ZS14]. linearization [Chr11, HM18]. linearizations [DLPV18]. linearized [BMGN18, NDL14]. Linearly [KL18a, AL15, PCCC18]. Liouville [RI15]. Lipschitz [Li15]. Lissajous [EKAB16]. Lobatto [BCD15]. Local [BBB17, BDLM18, FM18, HH11, SX13, SX15a, SX15b, BV12, BG16, Dem16, DH17, HS19b, LP15, Sch11b]. localizing [PSV18]. locally [AGS10, JP13]. Locating [Men11]. location [DV11]. locking [BDGQ12]. locking-free [BDGQ12]. log [SAI17]. log-determinant [SAI17]. logarithmic [BCST19, BM11]. lognormal [GKN^{+}15, GKN^{+}18, HS19b]. lognormal-parametric [HS19b]. Long [HL16, TWW15, CGSW16, HOR17a, HS18, Wan12]. Long-term [HL16]. Long-time [TWW15, CGSW16, HOR17a, HS18]. Love [LL12]. Low [BB13, GKL15, BBKS18, CP18, CC19, Hac16, Nou19, SSS16]. low-order [CP18]. low-rank [BBKS18, Hac16, Nou19, SSS16]. lower [CG14, HHS15]. LU [LGC14]. Lur’e [MOR17]. Lyapunov [BB13, BL13b, BV14, KS11a, SSS16].

M [XXL12b, XXL12a]. M-matrix [XXL12b, XXL12a]. MAC [GMN19]. MacCamy [GHS12]. Maclaurin [JT16, Pin18]. macro [Mat13]. macro-element [Mat13]. made [Lin17]. Maeztu [HJZ14]. magnetohydrodynamics [WS19]. Mahony [BMGN18]. majorization [LMSS17]. majorization-minimization [LMSS17]. managed [JM18]. manifold [GSY17]. manifold-valued [GSY17]. manifolds [Har18, HRS12, HAG17, Man15b]. map [BPS10]. mapping [HPS16]. maps [BPS10, GT19, Har18, ZLL^{+}12]. marching [BCPS10, Mir14]. marginal [BCN19]. marking [KS11c]. Marquardt [BGR18, HH10, Jin10, JY16]. Maruyama [LS18]. mass [GGGS11, Kir17]. massively [DK16b]. matched [CCZ13]. matching [YWY14]. materials [LLY15, SzS12]. Mathematical [Smy10]. matrices [VVVF10, BS17, BM19, Bör10, DK11, DP16, FMP15, FKS18, GMP^{+}14, Hac16, Koe19, LGC14, Lin17, Man13, MV15, OYV17, RB11, RMS17, Sid11b, FKS19]. matrix [BG12, BMP10, CCR14, CCGP17, Che16b, DHS17, DLPV18, DHM19, FMP15, FPD12, FLZ19, GKL15, HLS11, HAG17, JLL15, Kir17, KLS19, Men11, OR10, SAI17, Str15, XXL12b, XXL12a, XL17, ZBJ18]. matrix-free [SAI17]. Max [GNS15, YBTB14]. max-flow [YBTB14]. Max-norm [GNS15]. maximal [JLZ18, KS18b, LV17, Li15]. Maximum [DK16a, HZZ18, Li15, LHQ14, DKS13, HKK12, Pru14]. Maximum-norm [DK16a, Li15]. Maximum-norms [HZZ18]. Maxwell [NKL21, BBSV13, DLT12, HJS15, KL17, LS16]. mean [DFL16, KLL19]. mean-variance [DFL16]. means [Rus10]. Meany [BL13a]. measure [SG… 12, dGKL19]. measure-valued [SG… 12]. measures [AORS17, LRS19, LRS20]. mechanical [ABC15]. media [BBFP18, BGG^{+}16, CLLS19, CS17, CDL19, GGM14, KP17, KPR14, LDQ11, SS15]. Median [Nie10]. meets [HANT16]. Meixner [JJT13]. membranes [BGN16]. meromorphic [DV11]. Mesh [DCC18, Rad19, Ber16, GMÖ^{+}18, JP13, MLLR14, PZ18]. mesh-dependent [Ber16]. meshes [ABH13, ABB15, BN16, BCGS11, BDF12, DH17, GGRG15, HZZ18, HM13, Kop17, LS16, LN18, Loz19, Mak18, Mir12, SW11a, Wei11, WS19, YWY14, ZZ15]. Meshfree [Fro18, Nak17]. meshless [KNWW19, Sch10]. metamaterial [LS16]. method [AH14, AH17, AKYZ18, ABH13, AR16, ABAC16, BZ17, BT15, BGN16, BSZ11, BG12, BGR18, Bel11, BIM11, Ber16, BBC^{+}18, BL13b, BGHL14, BDGQ12, BNRS17, BDS13, BS10, CEM14, CGPS17, CGMM14, CSW16, CP18, CWHLT15, CW12, CCZ13, CQR16, CH12, CM15, DS11, DWWW17, DO17, DH17, DK16b, DLT13, ER15, EOS17, EH12, EGLTP17, FM18, FKM16, FKNP11, FL15, FGL14, FLS10, FR19, GY12, GHP17, GW19, GGM14, Gau18, GV15, GKSS13, GY17, GKL19, GWX16, GAB13, GSV19, HL16, HKK12, HPS16, HT18, HR19, HY15, HOR17a, HR10, HKXZ16, HSX12, HL18, HH11, Jin10, JZ13, JY16, Jin17, JL13b, Kal15, KdS18, Kar13, KOZ16, KS18b, Kim19, KSD11, Kla15, KS11a, KNWW19, LS18, LCH12, LZ15, LMY18, LP15, LCWW17, LMR16]. method [Loz19, LL10, LJS18, Man15b, MLLR14, MOR17, MP18b, MNT13, Mur17, Nei10, NZ19, NDL14, OR10, PS17a, PS18, PT12, PCCC18, QRB11, RS18, Rus10, SSS12, Sau17, SV13, Sch11a, Sch11b, SV12, Sti10, SH18, Tak15, TT15, Wac17, WZ10, WXY12, WG13, YS13, YWY14, ZS14, ZBB14, Zha16a, ZBJ18, dVMRR17, dVCMR18, BHT17]. methods [AV12, Ais15, AMN11, AORS17, AHR14, AL17, ADM19, AP10, ABC15, BK12, BL13a, BD12a, BCST19, BHK10, BMWB14, BGJ18, BB13, BR11, BS12, BDG^{+}18, BV17, BCGS11, Bre13, BLS14, BOS18, BHLZ16, BO18, CGHW11, CG17, CGH13, CPB17, CBHW18, CMMV14, CXH10, CNX12, CC13, CCSS16, CJ14, CC10b, CS12, CHL13, DM16, Dem16, Dem17, DE16, DSX16, EÖ17, EMR19, EGH15, Era15, FGP10a, FGP10b, FGH^{+}16, FGHP17, Gan13, GG12, GGK^{+}19, GVY14, Git14, GWX19, GKN^{+}15, GS16, HR14, HS19a, Har15, HLZ12, HP17, HPS18, HKT14, HW13, HW18, HHS15, HMX17, HFL12, HV12, JP13, JT11, Jun15, KSD10, LV13, LMS19, LRS19, LRS20, LV17, LA11, LC14, LOSV16, LVY14, LFS16, LM15, MSV17, Man15b, MMMKV16, MR17, Mus15, Nak17, NTW18, OS14, OTMS13]. methods [OZ17, OZ23, PS11, Pru14, Rei12, RE14, RSV13, RMS17, Run14, Sch10, SZ13, SS16, SW11b, Ste14, SzS12, SX15a, SX15b, TOMS13, TSGU13, Ver10, WX12, WHC14, Wil19, Wor19, XZ10, Zam13, ZRK16, Zha18, ZWJ19, ZS17, dVBMR16]. metric [ZLL^{+}12]. metrics [BCPS10, Mir14]. metropolized [FS17]. MFS [Smy10]. MHD [BKW10, BL17, HMX17]. micromagnetics [FLMP12]. midpoint [JM18]. mimetic [LVY14, dVM11]. min [YBTB14]. min-cut [YBTB14]. Mindlin [BC15b, PS17a, dVM11]. Minimal [DS18a, CD11, SW19]. minimisers [DKS13]. minimization [EGLTP17, FLMP12, FLS10, HHYY14, LMSS17, YWY14]. minimum [ZS12]. minmax [BV17]. miscible [CLLS19, CDL19]. misfit [HW13]. Mixed [CK16, CGK18, LMMR19, AMORB16, AT19, BGIS17, BHK10, BRK17, BEG14, BKUV17, CDNP16, CGG19, CW12, DO17, DSX16, ER18, GGM14, GWX19, GN14, KK15c, KY12, LM11, LCWW17, Ste10, ST18, Sti10, Sun16, WXY12]. model [AKYZ18, BBD16, CDL19, DHM19, GT14, GGM14, GGGS11, GGGS19, KPR14, LDQ11, Rad19, Sin12, Zam13]. Modeling [BDQ11, BG16, HJHA17, QRSZ19]. modelling [BKTT10]. models [BBFP18, CCS17, CK16, CGK18, HMX17, MOS11, SKW19]. Modified [BS12, NGKN10, Ver17, dS18b, Adc10, ACWL14, SSS12, SWS16, WG13]. moisture [SzS12]. Molodensky [CGS14]. Monge [BCMO16, Nei10, NZ19]. Monotone [CCP13, BMZ10]. monotonicity [GS17]. monotonicity-based [GS17]. Monro [DMG19]. Monte [LRS20, BSZ11, BC15a, BD12b, GGK^{+}19, GKSS13, GKN^{+}15, HANT16, HMOU16, LRS19, PGW18, TSGU13]. Morley [HSX12, HM16, HH11, Nei10, PS13]. Mortar [GVY14, GSV19, WX12]. motion [DL11, GGGS19]. moving [DS18c]. Muckenhoupt [NOS16]. Multi [BSZ11, CMMV14, HANT16, LP11, BHK10, BCN19, BL17, GKSS13, LCH12, dHQS15]. Multi-index [HANT16]. Multi-level [BSZ11, GKSS13, LCH12, dHQS15]. multi-marginal [BCN19]. Multi-parameter [LP11]. multi-point [BHK10]. Multi-revolution [CMMV14]. multidimensional [LS18]. multifidelity [PGW18]. Multigrid [BFK^{+}16, BCGS11, BLS14, BOS18, GWX16, HKXZ16, NN11, Not16b, Tak15]. multilattices [OLOV18]. Multilevel [CQR16, Sch11a, BC15a, CNX12, DMG19, GWX19, TSGU13]. multilinear [NTW18]. multiple [AH14, ABER10, CLM17, Che12, Nak12]. multiplication [FLZ19]. multiplicative [Bre13, CvN13]. multiplicity [Men11]. multiplier [AKYZ18, BG17, HR10, Jun15]. multipliers [HY15]. multipoint [DSX16, WXY12]. Multirate [GS16]. Multiscale [LSW12, MP18a, Wen10, AH14, AH17, BZ17, EMR19, EGH15, GVY14, LDQ11, PS11]. multispectral [GWCH10]. Multistep [SCHH13, CHL13]. Multivalued [ZT12]. Multivariate [Adc10, CCR14, HMOU16, Sor10]. multiwave [BKW10]. Mumford [SKW19]. Muscl [CC10b].

Nash [CGS14]. natural [CRS16, CPB17, PS18]. Navier [DWWW17, BDQ10, DO17, FKM16, GMN19, GNS15, GKL19, HR19, He13, HOR17a, HJNS17, Kar13, LA11, LCH12, LC14, LL10, SWS16, Wan12, ZT12]. nearest [Men11]. nearly [GWX16]. nearness [GKL15]. necessary [HKK12]. negative [YS13]. negative-order [YS13]. Nested [BCD15, Sou13, BV12, BKV15, BC16, BM17]. nets [HMOU16]. network [PC13]. networks [EK18]. Neumann [CD11, Zam13]. neutral [Mas16, WZ10]. Newmark [KSD10, KSD11, Kla15]. Newton [EH12, HW13, JT11, Jin12, JZ13, Kal15, Kim19, LGL17, Man15b, SX13, SX15a, SX15b, ZBJ18]. Newton-Halley [Kal15]. Newton-like [SX15a, SX15b]. Newton-type [HW13, SX13]. Nicolson [KK15a, Wan18]. Nitsche [CH12, CMR18, GSV19, JL13b, Jun15, LR17, LJS18, MLLR14]. Nitsche-based [CH12]. Noda [JLL15, LGL17]. Nodal [CHH18, AMN11]. node [EKAB16]. nodes [BDHK12, MdC17]. noise [CvN13, GLW19, ZRK16]. noisy [HSW17]. Non [BBFP18, CEM14, HK19b, AP10, BV12, BG16, BCM17, BMZ10, BEG14, CGMM11, CDNP16, CLMS18, CCR14, CCGP17, Che12, Che16a, CP14, CN19, Dri12, DL15, DE16, EOS17, EKAB16, FS14, GGMO16, GSY19, HY15, LCH12, LS16, WE16, YWY14]. non-consecutive [Dri12]. Non-consistent [CEM14]. non-convex [CLMS18]. non-cylindrical [AP10]. non-degenerate [EKAB16]. non-ergodic [HY15]. Non-Fickian [BBFP18]. non-linear [DL15, DE16]. non-local [BV12, BG16]. non-matching [YWY14]. non-monotone [BMZ10]. non-orthogonal [YWY14]. non-periodic [GSY19]. non-permeable [CP14]. non-product [Che12]. non-relativistic [FS14]. non-rooted [BCM17]. Non-satisfiability [HK19b]. non-self-adjoint [GGMO16]. non-selfadjoint [CGMM11, CDNP16]. non-singular [LCH12]. non-smooth [CN19, WE16]. non-staggered [Che16a]. non-stationary [CCR14, CCGP17]. non-symmetric [EOS17]. non-uniform [LS16]. nonconforming [CPR13, CGH13, CDNP16, CRS16, CXH10, CCQ13, Era15, HHS15, LMcS17, LL10, Nei10, SS15, SZ13]. nonconstant [BPS10]. Nonconvex [LMSS17, CWHLT15, GLS17, PT10]. nondominated [Cal17]. Nonlinear [CBHW18, AV12, BFG13, BT15, BBK17, BGR18, BC12, BV17, BEK11, CG17, CCLM15, Cho18, CGK18, FN12, FFKP15, FKNP11, GLS17, GMSS11, HH10, HW18, HFK19, JM18, JMM12, Jin10, Kal15, KKN14, LA11, MR17, Mur17, Nak17, Nei10, SS11, SzS12, SX13, SX15a, SX15b, WZ10, Wan18, dHQS15]. nonlocal [GWW14]. nonmatching [DH17]. nonmonotone [AV12, PZ18]. Nonnegative [BSCZ12, Koe19, LN15, LGL17, ZBJ18]. nonobtuse [Dal10]. nonorthogonal [Str15]. nonoverlapping [ABAC16, DK16b]. nonperiodic [DNP14]. nonrelativistic [BD12a, BZ17]. nonsmooth [LMSS17, RE14]. nonstandard [Ste10]. Nonstationary [JZ14, HT18]. nonsymmetric [CXH10, FFKP15, LGC14, RB11]. norm [CPS16, DK16a, EGLTP17, FLZ19, GNS15, Kop17, Li15, Pru14, YS13]. normal [FGP10a, FGP10b]. norms [AdST12, HZZ18, Mir12, PS18]. note [BY15, CM18]. novel [GWCH10]. number [JL13a]. numbers [DP16, Nak12]. Numerical [AH17, BDQ10, BBSV13, BM13, BLP19, Cal17, CC19, DV11, DES11, DM14, DHM19, FLMP12, Fri15, HSD18, NDL14, Nov14, RI15, Rus10, AV12, AN18, BBL11, BLS15, BD12a, BCST19, BGN16, BKW10, BD12b, CD11, CCLM15, CHW17, CP14, CJ14, CC10a, CR12, DS16, DS18a, DE16, FKM16, GGGS19, HS19a, HMOU16, HJT14, KKN14, LP11, MP15, MMMS10, Mas16, PSV18, PSV10, RSV13, Ven11, WMS10, ZS17]. numerically [Olv12]. numerics [CLM17, LT10, TWW15]. Nyström [BDS13, FL15].

objects [Rie19]. observability [EZ15]. observers [HR12]. obstacle [BC15a, CM18, ZVKG11]. obstacles [BCS14b, CWHLT15, CC10a]. Oliker [NZ19]. once [Tak15]. One [BK12, HL18, OS14, SB13a, SB13b, ZS14]. One- [BK12]. one-dimension [ZS14]. one-equation [OS14]. only [Pin18]. Ono [DHKR16]. onto [BY14]. open [BGN12]. operator [BMN11, BAK19, BCMO16, BFK^{+}16, BHL^{+}19, Eng14, LO14, MSW13, Nov14, Sch10]. operators [AN19, ABM18, BLM11b, BBB17, BV12, Bör10, BC16, BR17, DM14, FY11, Git14, GEEF17, HFK19, HHS15, KKN14, KT12, MS19, Wor19, vyKS14]. Optimal [CPR13, CNX12, FGHP17, HMOU16, LR17, LC14, Rei12, SA10, AORS17, ADF18, BC17, CRS16, CSW16, CQR16, DFL16, DD16, FGH^{+}16, FHPS19, Gal15, GY17, GWX19, GWX16, HSLZ19, JSW18, LV13, NV12, NPVW19, NTT16, OW11, Pla15, RS18, RW12, SG11, SH18, TWZ17, dGKL19]. optimality [CNSV17, CXH10, DS11, Dem16, Dem17, GY17, HSX12, Jin12]. Optimally [Mir12]. Optimization [AY18, GV15, GP11, GMT11, LMSS17, NW18, PG17, Str15]. Optimized [BGGH16, BPZ12, MSV17]. options [MOGO17]. order [AHR14, AdST12, ADBN16, BLF19, BMN11, BHK10, BS12, BPZ12, BKN13, BGHL14, BDS13, BEJ14, BKTT10, CGH13, CDNP16, CP18, CCFM17, CCQ13, CGSW16, CS12, DWWW17, HU17, mHZZ17, HZZ18, HMOU16, HKT14, HK19b, HL18, HH11, Jin12, JL13b, KS18a, KY12, KNWW19, Mir12, MP18b, MNT13, NN11, NS14, Nou19, RS18, Ren17, Run14, SV12, SH19, SKW19, Wan12, WS19, YS13, ZS12, ZS14, ZBB14, ZZ15, ZRK16, Zha18]. ordinary [AR12, HHB16]. orthogonal [DBV11, ÖS13, PR13, YWY14, ZDQ17]. Osborn [Mak18]. oscillation [CG11]. oscillation-free [CG11]. oscillatory [AN18, CMMV14, CCLM15, CLM17, Olv10, WZH13]. OSLC [Tow18]. other [FHN^{+}14, Man15b]. outflow [GKL19]. over-penalized [BC15b]. overdamped [FS17]. overdetermined [Str15]. overlap [BFK^{+}16, MSV17]. Overlapping [EMR19, FLS10, MLLR14]. overlaps [SDH^{+}14]. Overrelaxation [LM15].

P1 [KOZ16]. P1/P1 [KOZ16]. Padé [BRZ13]. Padé-type [BRZ13]. pair [LGL17, Zha16b]. pairs [BEK11, Che16b]. Palindromic [HLS11, LLW10]. parabolic [AH17, AL15, ABH13, BRK17, BHL11, BS10, CC19, DE16, EH18, HFK19, KS18b, KKN14, KNWW19, LM11, LV17, Li15, LM13, MP18a, Nak17, NV12, Pru14, Rei12, RSV13, Sti10, ZDQ17]. parabolic-elliptic [Sti10]. paraboloids [dCMT16]. parallel [DK16b, HH11]. parallelepiped [KK15b]. parameter [JZ13, LP11, NPVW19]. parameterized [BEK11, DP16, TWZ17]. parameters [dGKL19]. Parametric [BGN12, EH13, HHB16, HS19b]. parametrized [RHM13]. Part [FGP10a, FGP10b, PS11, SB13a, SB13b]. partial [BSCZ12, Bör10, Git14, GLW19, HV12, Rus10, Sti10]. particle [PC13, PCCC18, Sch11a, Sch11b]. particle-partition [Sch11a, Sch11b]. partition [Sch11a, Sch11b]. partitions [BLM11a]. Pathwise [CvN13]. pattern [GT14]. PCG [FHPS19]. PDE [BO19, GKN^{+}18, NPVW19, PG17, Sin12]. PDE-constrained [NPVW19, PG17]. PDEs [FGP10a, FGP10b, GKN^{+}15, HS19b, KS18a, LSW12, Lui17, NTT16, OR10, PZ18, Rei12, SDH^{+}14, TSGU13, TWZ17, WE16, ZRK16, BSZ11, PS11]. penalization [vyKS14]. penalized [BC15b, LL10]. Penalty [DU15, KOZ16, BNRS17, HS19a, HR19, JZ14, KP17]. penalty-projection [HR19]. perfectly [CCZ13]. periodic [AY18, CMP13, GW19, GSY19, GWW14, HMOU16, HJNS17, MOS11]. periodic-like [GW19]. permeable [CP14]. Perron [LGL17]. Perturbation [DK11]. perturbed [AV11, DK16a, KS11b, Kop17, SH17]. Petrov [CP18, CBHW18, EGH15, HK17, Ste15]. Phase [BMO15, CCS17, CS17, CHW17, DHM19, HHNS16, Hor17b, KP17, SS15, SV12]. Phase-field [BMO15, CCS17]. photonic [Eng14, GG12]. Piecewise [KR16, NOS16, CHW12, LL12]. pipes [BEG14]. Pitaevskii [AL17, HKT14]. placement [NPVW19]. plane [GN14, IG15]. plate [BNRS17, LL12, dVM11]. plates [BC15b, PS17a, SH18]. POD [Sin12]. point [BHK10, BLS14, BOS18, DM14, FGH^{+}16, HW11, KSZ13, PG17, PPS18, SG11, Sou13, Tu11, ZWJ19]. points [DFL16, EKAB16, Mas14, XCW10]. Pointwise [NZ19, NW18, OW11]. Poisson [CCFM17, CCSS16, HW13, LP15, RS11]. polarization [AGK^{+}14]. polefinding [IN18]. poles [DV11]. Polyak [DMG19]. polygon [GLS17]. polygonal [LVY14, Wei11]. polyhedra [GNS15]. polyhedral [BN16, FM18, LN18, LVY14]. polynomial [BD11, Bel11, CHTV18, DGSY17, LLW10, MS19, NOS16, SV12, TWZ17]. polynomial-degree-robust [CHTV18]. Polynomials [HN16, CCR14, DLPV18, Dri12, Eng14, HLS11, JO10, JJT13, Kir11, KT12, LT10, Not19, ÖS13, ST18, dS18b]. poroelastic [AKYZ18]. porous [BBFP18, BGG^{+}16, CLLS19, CS17, CDL19, GGM14, KP17, KPR14, LDQ11, SS15]. posed [CN19, HH10, Kal15, Sch16]. posedness [BG16]. positive [AN19, BM13, MOR17, WS19]. positivity [HK19b, JLL15]. post [ER18]. post-processing [ER18]. posteriori [AV11, AMN11, AMORB16, BB12, CCM12, CDM^{+}18, CGPS17, CM13, CGH13, CPS16, CGG19, CHTV18, CS17, DK16a, DSX16, Era15, HSV12, KK15a, KKN14, Kop17, KVW15, LM11, LMS19, LO14, LM13, NvPZ10, RW12, RHM13, WX12, dVMM19]. postprocessing [CM13]. potential [ABK13, BM11, LO14]. potentials [GY12, SV16]. power [HL18, Pet12]. powers [AN19]. precipitation [KPR14]. Precise [Fuk13, AKST14]. Preconditioned [PT10, LM15]. preconditioner [DGS15]. preconditioners [LR17, PPS18]. Preconditioning [BMN11, BFK^{+}16, GGS15, LSW11, Sch11b]. prescribed [ZBJ18]. presence [dVCMR18]. Preserving [WZ10, BDF12, Egg19, FS14, GAB13, HMX17, HLS11, JLL15, KLS19]. prespecified [Men11]. pressure [LMS19]. pressure-robust [LMS19]. Pricing [MOGO17, HFL12]. primal [ABAC16, Loz19]. primal-dual [ABAC16]. primitive [BGHL14, HS18, Pet12]. principal [Nou19]. principle [DKS13, LHQ14, ZS12]. priori [AV12, AMORB16, BKN13, CM18, LN18, LJS18, MOS11, NV12, NW18, OW11]. probability [LRS19, LRS20]. problem [AORS17, AMORB16, BJLZ15, BY19, BDG^{+}18, BHLZ19, CPR13, CRS16, CD11, CC15, CHTV18, CCQ13, CH18a, CM18, CN19, CGS14, DES11, DBV11, DO17, DLT13, FL15, GMS19, GKL19, GGRG15, HSV12, HQ12, HFL12, JMM12, JL13a, KS11b, KVW15, Kre17, LN18, LLY15, MLLR14, Nak12, OB17, QRB11, Sou13, ST18, Sun16, Tak15, Tu11, WHC14, Zha16a, Zha16b, dVM11, dVMRR17, dCMT16, GKSS13]. problems [AV12, AH14, AH17, Adc10, AV11, Ais15, ADF18, ABER10, AG10, BSCZ12, BBSV13, BLS15, BMN11, BGIS17, BT15, BRK17, BKV15, BGR18, BGJ18, BS12, BV17, BEK11, BC15a, BDQ11, BLS14, BOS18, BDS13, BS10, BHLZ16, CGMM11, CG11, CGH13, CGMM14, CDNP16, CXH10, CCZ13, CQR16, CC13, CM15, DD16, DK16a, DCC18, DGS15, DK16b, EÖ17, Egg19, EMR19, EMB10, FN12, FFKP15, FKNP11, FGL14, GS13, GW19, GGMO16, GV15, GGK^{+}19, GMSS11, GY17, GKL15, HL19, HS19a, HPS16, mHZZ17, HSLZ19, HR10, HLZ12, HH10, HW18, HW11, HH11, Jin10, JZ13, Jin17, JL13b, JSW18, Kal15, Kim19, KKN14, KSD10, Kop17, KSZ13, LM11, LR17, LLW10, LMR16, LHQ14, LM13, Mas16, Mus15, NV12, NW18, NPVW19, NTW18, Not16b, OS14, OPS15, Olv12]. problems [OW11, PG17, PPS18, PT10, QRSZ19, Rad19, RI15, RSV13, RW12, Sch16, SH19, SX13, YBTB14, ZZ15, Zha18, ZWJ19, ZDQ17, ZVKG11, dHQS15]. procedure [LA11]. processes [Ven11]. processing [ER18]. product [Che12, DS10, HS19b, SS11]. program [HHB16]. programming [PG17]. projected [dHQS15]. projection [AG10, BY14, BV17, HR19, Ren17]. projections [GH14a, Ste15]. prolongation [NN11]. Prony [Sau17]. proofs [CFL11]. propagation [BLM11b, BCS14b, HR14]. propagator [LS17]. Proper [FN12, ZDQ17]. properties [ABB15, BHL11, Che12, FS17, FS19, GAB13, HFK19, IG15, KSZ13, OR10, Wan12]. property [BY15, BV17, DKS13, OTMS13, TOMS13]. Provably [Rie19, WS19]. Prussner [NZ19]. pseudo [HKT14]. pseudo-spectral [HKT14]. pseudodifferential [PT14]. pseudospectral [BZ17]. pseudostress [GGM14]. pseudostress-based [GGM14]. punctual [BDLM18].

QCD [BFK^{+}16]. QMC [GKN^{+}18, HS19b]. QR [VVVF10]. QR-algorithm [VVVF10]. quad [Sun16]. quad-curl [Sun16]. Quadratic [BR17, BMP10, Eng14, FLMP12, OYV17, PG17, XZ10]. quadratization [HLS11]. quadrature [AV12, AR19, BBSV13, BLM11b, BC16, DMN12, FR19, FHN^{+}14, HS15, KT12, LFS16, Not16a, Not19, OMS13, Owe16, PS16, Pla15, dS18b]. quadrature-based [KT12]. quadratures [GST19]. quadrilateral [HZZ18, Mon19, PS13, ZZ15, Zha16b]. quadrilaterals [WXY12]. Qualitative [HFK19]. Quantized [KS18a]. quantum [BCS14a, BEJ14]. Quasi [BD12b, Dem16, GKN^{+}15, Maz11b, SH18, AL15, BM19, CRS16, CXH10, DBV11, DS11, Dem17, GGK^{+}19, HM13, HMOU16, KL18b, NTT16, SM16, TWZ17]. quasi-interpolation [SM16]. quasi-linear [AL15, KL18b]. Quasi-Monte [BD12b, GKN^{+}15, GGK^{+}19, HMOU16]. Quasi-optimal [SH18, CRS16, NTT16, TWZ17]. Quasi-optimality [Dem16, CXH10, DS11, Dem17]. quasi-orthogonal [DBV11]. quasi-Toeplitz [BM19]. quasi-uniform [HM13]. quasiconformal [ZLL^{+}12]. quasilinear [CD11, CC19, DD16, HP17, HPS18]. quasiseparable [DP16]. quasistatic [WHC14]. quaternionic [JO10]. queue [NP15]. quotient [AV10].

Rachford [HHYY14, HY15]. Radau [PS16]. RADI [BBKS18]. Radial [HSW17, PT14]. radiation [HK19a]. radii [BPS10]. radius [LN15]. Radon [GH14a]. Random [Dic12, OZ17, AR12, BBC^{+}18, BC15a, DHS17, FKS18, FKS19, Git14, GKN^{+}15, HPS16, MNT13, NTT16, TSGU13, OZ23]. Randomized [Lin17, SAI17]. rank [BB13, BBKS18, BKUV17, EHRS12, GK19a, GKL15, Hac16, HRS12, HL18, Nou19, SSS16]. rank-1 [BKUV17]. rank-adaptive [GK19a]. rate [AR16, HY15, HL18, SH18]. rates [BC17, Cal17, Gan13, HQ12, KS11c, NZ19, SH19]. Rational [AN19, BG12, BLM18, BDHK12, BRZ13, CH18b, DBV11, GT19, GN14, IN18, MMMS10, PR13]. Raviart [HM16, LMR16]. raw [DM14]. Rayleigh [AV10, BJLZ15, GHP17]. RBF [LSW12]. reaction [AV11, ABH13, BGGH16, DK16a, FGL14, KS11b, Kop17, LHQ14, Pie18]. reaction-diffusion [AV11, DK16a, KS11b, Kop17, Pie18]. real [HU17, Kim19, MOR17]. realizable [ZBJ18]. realization [LP11, RSV13]. Reconstructing [HR12]. reconstruction [GS17]. reconstructions [PSV18]. recovery [DV11]. rectangles [CGH13]. rectangular [JP13, LS16]. Recurrence [PR13]. recurrences [MV15]. recursive [Hac16]. red [SZ13]. red-green [SZ13]. Reduced [RHM13, CQR16, Zha18]. reduction [Bar13a, SZ13, Sin12]. Refined [LMS19, PPS18]. refinement [SZ13, dVBRS11]. refinements [GMÖ^{+}18]. reflector [dCMT16]. Regge [Chr11]. regime [BD12a, BZ17, FS14, NDL14]. regions [AN18]. Regular [Zha18, Che16b, Dal10]. Regularity [CCGP17, LN18, CG17, CD11, CC19, DHS17, JLZ18, KS18b, KY12, LV17, Li15, Pru14, dGKL19]. Regularization [ZWJ19, GS17, HQ12, HKQ18, JT11, Jin12, JZ14, JY16, Jin17, LA11, LP11, OB17, SH19]. Regularized [BCST19, EH12, HW13, Jin10, JZ13, Kal15]. regularizing [HH10]. Reissner [BC15b, PS17a, dVM11]. related [BB13, BR11, BGHL14, FL15, Not19]. relation [BBK17]. relations [PR13]. relationship [Fer13b]. relative [FH17]. relativistic [FS14]. relaxation [ADF18, AL17, BGGH16, BKW10, GW19]. Relaxing [GHS12]. reliable [GST19, KVW15]. remarks [BO17]. reordering [OZ17, OZ23]. Reorthogonalization [Bar13a]. Reorthogonalized [BS13]. Representation [GT19, HAG17]. representations [EPS17, VD11]. represented [EH12]. Reproducing [FY11]. Reproduction [CCR14]. requirements [DCC18]. Residual [CGG19, DSX16, Wei11, CS17, KVW15, WX12]. Residual-based [CGG19, DSX16, CS17, WX12]. residual-type [KVW15]. residue [DV11]. resolution [BGHL14]. respect [BCPS10, dGKL19]. restarted [Ais15]. Restarting [BV17]. resultants [NNT15]. results [KSD10]. retarded [BV14, SV13, SV16]. reverse [BMWB14]. revolution [CMMV14]. Rham [CHH18]. Riccati [BBKS18, BMP10, XXL12a, XL17]. Ricci [Fri15]. Riemann [BKW10, Olv12]. Riemannian [Har18, ZBJ18]. rigid [SSS12]. Rigorous [GL11]. Ritz [GHP17]. Robbins [DMG19]. Robin [CGHW11, LJS18]. Robust [GGMO16, RB11, AV11, CDM^{+}18, CHTV18, Dic12, HKXZ16, LMS19, LMcS17, Ren17, SDH^{+}14, Tak15]. rod [CR12]. roles [RHM13]. rooted [BCM17]. rotating [BHT17]. rotational [HKT14]. roto [CP19]. roto-translational [CP19]. rough [DHS17, SS18]. rounding [MdC17]. row [BL13a]. row-action [BL13a]. rule [HS15, JM18, Jin17]. rules [ACWL14, AR19, BD11, BD12b, DMN12, Dic12, DNP14, DGSY17, GSY19, GMT11, HMOU16, Owe16, dS18b]. Runge [Man15a, AMN11, BBSV13, BLM11b, GS16, HPS18, LFS16, Ver10, WE16, ZS14]. Ruppert [DMG19].

Sacker [BV14]. saddle [BLS14, BOS18, HW11, KSZ13, PPS18, Sou13, Tu11]. saddle-point [PPS18, Sou13]. same [OTMS13, TOMS13]. Sampling [RW17, AdST12, BBC^{+}18, BKUV17, DFL16, HANT16, LRS19, LRS20]. Sard [Che12]. satisfiability [HK19b]. saturation [BY15, CGMM14, CGG16]. scalar [ABER10, MR17, NS14, Tow18]. scale [BBKS18]. scales [AH14, JT11]. Scattered [GSY17, HSW17]. scattering [ABER10, BDS13, CWHLT15, CCZ13, CC10a, EÖ17]. Scharfetter [BC12]. Scheidegger [CLLS19]. scheme [AKST14, AMORB16, AG10, BC12, BMZ10, BCS14b, BL17, Che16a, CLM15, CvN13, FH17, GMN19, GWW14, GK19b, GGGS11, He13, HJHA17, HH10, HS18, KP17, LS16, Ren17, SS15, SS18, Tow18, Ven11, Wan12, YWY14]. schemes [AGS10, ADBN16, BBK17, BC15b, BDF12, BKTT10, Cal17, CCR14, CCLM15, CHW17, CDL19, CKL14, CDKLM18, CC10a, CR12, DL15, DHKR16, EZ15, FS14, Fer13a, FLOP10, Fer13b, GM14, GV19, HZZ18, HM14, JLZ18, MS19, PSV10, RS11, WS19, ZS12, ZZ15]. Schmidt [BS13]. Schoenberg [BR17]. Schrödinger [ABK13, AL17, AP10, AN18, BFG13, BZ17, BCST19, CG17, CCLM15, GY12, GH14b, JM18, KK15a, LO14, Wan18]. Schwarz [AL17, BK12, BGGH16, Bre13, DK16b, EMR19, GV15, LMR16, MSV17]. SDEs [LS18, NS14]. SE [OTMS13, TOMS13]. Second [BKTT10, BK12, BJLZ15, BS12, CDNP16, CS12, DWWW17, KS18a, Mas14, MNT13, Not16a, SV12, Wan12]. second-grade [BJLZ15]. Second-order [BKTT10, CDNP16, CS12, KS18a, SV12]. Segel [ZS17]. segmentation [CLMS18]. selecting [DL11]. self [AN19, BBB17, CEM14, GGMO16]. self-adjoint [AN19, BBB17, CEM14]. selfadjoint [CGMM11, CDNP16]. Sell [BV14]. Semi [LA11, AdST12, AN18, AG10, BM19, CG17, CvN13, HR12, HJHA17, Pie18, RS11, SB13a, SB13b, dGKL19, Fer13b]. semi-classical [AN18, CG17]. Semi-discrete [LA11, HR12, Pie18, dGKL19]. semi-explicit [SB13a, SB13b]. semi-implicit [AG10, HJHA17]. semi-infinite [BM19]. semi-Lagrangian [RS11]. semi-linear [CvN13]. semi-norms [AdST12]. semiclassical [GH14b]. semidiscretizations [Say13, WE16]. semigroups [FS19]. semilinear [CHW12, FGP10a, FGP10b, KS11b, KNWW19, NV12, WE16]. Sensitivity [FPD12]. sensor [NPVW19]. separable [HHYY14]. separation [Che16b]. sequences [Dri12, Sid11a]. serial [OYV17]. series [Adc10, BCM17, MMMKV16]. several [BL13a, Sau17, SX13]. Shah [SKW19]. shallow [BL17, CKL14, CDKLM18, TT15]. Shanks [Sid11a, SCHH13]. Shannon [MOGO17]. shape [AGK^{+}14, ADF18, GS17]. Sharp [BLM11a]. Shaw [GGGS19, LCWW17]. Shaw-like [GGGS19]. shear [BC15b]. shell [Zha16a]. shift [VVVF10, GGS15, ST16]. shift-and-invert [ST16]. Shifted [Olv10, GGS15]. shock [HM14]. Short [MV15]. shortcomings [BT15]. shortening [PS17b]. sided [AV10, BBB17, JO10]. sign [CC13, DCC18]. sign-changing [DCC18]. signals [SKW19]. Signorini [KVW15]. signs [HHNS16]. similarity [VD11]. simple [FLOP10, Ise11]. simplices [BPZ12]. Simplicial [HS13, Kir11, KT12, Kir17, XZ10]. Simpson [Pla15]. Simultaneous [LM15, Sid11b]. Sinc [OMS13, OTMS13, TOMS13]. single [CS17]. single-phase [CS17]. singular [AR16, Bal12, FGHP17, HU17, HFL12, LCH12]. singularities [YS13, dVCMR18]. singularly [AV11, DK16a, KS11b, Kop17, SH17]. Sinkhorn [BCN19]. Sliding [DL11]. slip [DU15, KOZ16, LA11]. small [BD11, CC10a]. smallest [JLL15]. smooth [CH18a, CN19, DNP14, GY12, GL11, SV13, WE16]. Smoothing [NN11, SKW19]. smoothness [BKUV17, RMS17]. Sobolev [AP11, AdST12, FY11, GSY19, KY12, NOS16, Wen10]. Solution [BY19, BBSV13, Bör10, CC10a, HL16, HHB16, HFL12, Mas16, OB17, PG17, RI15, Rus10, SSS16]. solution-dependent [HL16]. solutions [BCCH19, Cho18, CMP13, EMB10, FGL14, GLW19, HJNS17, KKN14, KS11b, LCH12, Pet12, PZ18, SB13a, SB13b, WMS10, XXL12b, XXL12a]. solvable [CHW17]. solve [Bel11, SV12]. solver [BKW10, FHPS19]. solvers [GWX19]. Solving [GS13, Jin10, Kal15, LSW12, Olv12, PT12, WHC14]. Some [BO17, GV19, LN15, WX12, dVBRS11, BR11, CXH10, Rus10]. SOR [OZ17, OZ23]. SOR-type [OZ17, OZ23]. sorting [Cal17]. source [BDLM18, SG… 12]. space [BO19, BRK17, BHL11, FKNP11, HHNS16, HT18, HM14, KdS18, Lui17, Man15b, MP18b, NV12, NTT16, OPS15, Rei12, SG11, Ste15, Tow18, WZ10]. space-fractional [BO19]. space-time [HT18, MP18b]. space/time [BHL11]. space/time-discretized [BHL11]. spaces [BY14, BKUV17, EMR19, EGLTP17, FN12, FY11, FHN^{+}14, GSY19, JZ13, JZ14, JY16, KY12, Maz11a, Maz11b, NOS16, SW11a, SM16, SDH^{+}14, Wen10, ZWJ19, dHQS15]. spacetime [HS13]. Sparse [NTW18, ABB19, CHW12, DS10, HT18, NPVW19, NTT16, RW17]. sparse-grid [NTT16]. Sparsity [BPZ12, HANT16]. spatial [ZRK16]. spatially [YBTB14]. spatiotemporal [Kla15]. SPDEs [CvN13]. special [Seg13]. Spectral [ABM18, Eng14, HL15, ÖS13, RMS17, Wor19, BT15, GY12, HL19, HKT14, KSZ13, LN15, Lin17, Lui17, vyKS14]. spectrum [BV14, GMP^{+}14, Pet12, ZBJ18]. speed [MNT13]. sphere [BD12b, HSW17, LSW11, PT12, PT14, TT15, VGG14]. spheres [BPS10, FHN^{+}14, KNWW19, LSW12, Nie10]. spherical [CFL11, GP11, PT12, TT15]. Spline [BAK19, SW11a, Maz11a]. splines [PT12, RMS17, Sor10, TT15]. split [Mat13]. splitting [BMN11, CG17, CCFM17, FGP10a, FGP10b, Gau18, GH14b, GWW14, HHYY14, HJS15, HKT14]. square [LO14]. squares [BGR18, BC17, CPB17, DMMS11, IN18, JL13a, RS18]. Stabilised [KS11b]. Stability [BT15, DS18c, FFKP15, GLS17, KL18b, KSZ13, LSW11, SCHH13, BFG13, BY14, CC10b, EK18, GHP17, HOR17a, HJNS17, HS18, KK11, Li15, Mas14, RHM13, SWS16, WZ10, Wan18, Wil19, WMS10, dCM17, dHQS15]. stabilization [Ber16, PSV10]. Stabilized [BGIS17, BHLZ19, BS10, BO18, HR10, Jun15, KSD11, Kla15, LA11, LCH12, MLLR14]. Stable [BLS15, GK19a, HMX17, IN18, KL17, Sch11b, SW11b, Zha16b, BGN16, BEJ14, CHW17, HM14, HV12, MSW13, Pru14, Sch16, ZBB14, NKL21]. staggered [Che16a]. staggering [Che16a]. staircase [HR14]. state [CGHW11, NW18, OW11, RW12, SG11]. states [BFG13]. static [Loz19]. Stationary [AN18, CCR14, CCGP17, HSD18, He13, LCH12, LC14]. statistical [Wan12]. steady [BBD16, CGHW11]. steady-state [CGHW11]. steepest [dHQS15]. step [Pru14, SG11]. stepping [HW18, JLZ18, Mus15]. steps [Ren17]. Stieltjes [Not19, dS18b]. stiff [CS12, HJT14]. stiffness [GMP^{+}14]. stimulation [HJHA17]. Stochastic [AR12, BSZ11, CQR16, CS12, DMG19, EPS17, GGK^{+}19, GKSS13, GLW19, HL19, HR19, MNT13, TWZ17, WMS10, ZRK16]. Stokes [CHW17, AKYZ18, BDQ10, BJLZ15, BDLM18, BLS14, BC17, CGHW11, CPR13, CP18, CHTV18, CGSW16, CQR16, CH18a, DWWW17, DU15, DO17, FKM16, GMN19, GVY14, GNS15, GKL19, GGRG15, HSV12, HR19, He13, HOR17a, HKXZ16, HJNS17, Kar13, KOZ16, LMS19, LA11, LCH12, LC14, LVY14, LL10, MLLR14, Not16b, PPS18, QRB11, RE14, RHM13, SWS16, Tak15, Wan12, ZT12, Zha16b, vyKS14]. Stokes/Darcy [BDQ10, DO17]. Stopping [GMT11, Jin17]. Störmer [HL16]. strain [LMcS17]. Strassen [FLZ19]. streamers [RI15]. stresses [BDGQ12]. Strong [ZRK16, NS14]. Strongly [PT14, ABAC16, GMSS11, SzS12]. structural [KSZ13]. Structure [Egg19, AKYZ18, AG10, BS17, BG17, Eng14, Fer13a, FH17, GV15, HLS11, KLS19]. structure-preserving [HLS11]. Structured [DP16, LLW10, Hac16, KS18a, LGC14]. structures [AY18, CMR18]. studies [GWCH10]. Sturm [RI15]. subdivision [CCR14, CCGP17, CM10, MS19, Rie19]. subgradient [BCPS10]. subject [dHQS15]. submanifolds [LRS19, LRS20]. subspace [KS11a]. subspaces [GSY19]. Successive [LM15]. summation [Pin18]. sums [EHRS12, Pin18]. sup [HW11, RHM13]. Super [ER18, AR16]. super-algebraic [AR16]. Super-convergence [ER18]. supercell [CEM14]. Superconvergence [HM16, LS16, AMN11]. Superlinear [BG12]. supersmoothness [Sor10]. support [HHR11]. supported [SV13]. supports [HS19b]. sure [WMS10]. surface [BDS13, BHLZ16, BHL^{+}19, DS18c, ER15, KLLG17, KL18a, OR10]. surfaces [BN16, BHLZ19, DM16, DL11, GM14, KLL19, Man15a, SW19]. sweeping [Ven11]. Sylvester [XXL12b]. Symmetric [CHL13, GN14, Ais15, BDGQ12, BCGS11, CG11, CGG19, EOS17, FFKP15, Sid11b]. Symplectic [KK15c]. Symplectic-mixed [KK15c]. system [BCCH19, BL17, BGHL14, CGHW11, CGSW16, CHW17, CKL14, DWWW17, FKM16, Gau18, GGGS19, LCWW17, RS18, SSS12, SzS12, Wan18, ZS17]. systems [ADBN16, ABC15, BM13, Bel11, BBC^{+}18, BLS14, BOS18, Chá11, CH18a, CHL13, EK18, EZ15, GAB13, HL16, HM14, LM15, Man13, Mur17, Sin12, Smy10, SDH^{+}14, SA10]. Szego [PS16].

tangent [HAG17]. tangential [KdS18, NGKN10]. Taylor [MSW13, MS19]. TDNNS [PS17a, PS18]. technique [GW19]. techniques [AR12, Che16a, EZ15, HM18, MP18a, SH17]. temperature [CKL14]. tempered [BCS14a]. Temple [BBC^{+}18]. Temple-class [BBC^{+}18]. temporal [HS18, RSV13, SV13]. tensor [Bal12, CLLS19, DS10, EPS17, EH12, FN12, FLZ19, GK19a, Hac13, KS18a, LGL17, SS11]. tensor-structured [KS18a]. Tensorisation [Hac11]. tensors [AGK^{+}14, EH12, EHRS12, HU17, HRS12, HL18, LN15, Nou19]. term [BDLM18, HL16, RE14]. terms [JZ14, JY16]. tetrahedral [CCQ13]. tetrahedron [Mat13]. textile [SzS12]. Their [SCHH13, ABB15, BK12, DLPV18, Hac11, ZT12]. theoretical [BS19]. theory [ABM18, DK11, HS13, Nie10, Sin12]. thermistor [GLS17]. thickener [BKTT10]. thin [CW12, FLMP12]. thin-film [FLMP12]. third [CGSW16, ZS14]. third-order [CGSW16]. three [ABER10, Ber16, CCQ13, GL11, GT14, HS18, Pet12, Tu11]. three-dimensional [ABER10, CCQ13]. three-level [Tu11]. tide [CK16, CGK18]. Tight [BKUV17]. Tikhonov [HQ12, JZ14, KdS18]. Time [ADBN16, BKN13, HR19, Mus15, QRSZ19, AL17, ABC15, BZ17, BRK17, BS12, BCCH19, CCS17, CG17, CCZ13, CGSW16, DE16, EH18, EZ15, FKNP11, GH14b, GEEF17, HHNS16, HT18, HOR17a, HM14, HKT14, HW18, HJNS17, HS18, JLZ18, KL18b, LM13, Lui17, Man15a, MP18b, NV12, Pie18, Pru14, Rei12, Ren17, RE14, SV16, Say13, SG… 12, SS16, TWW15, Tow18, Wan12, WE16]. time-dependent [SG… 12, SS16]. Time-discrete [BKN13, EZ15]. Time-discretization [HR19]. time-discretized [BHL11]. Time-domain [QRSZ19]. time-splitting [CG17, GH14b, HKT14]. Time-stepping [Mus15, JLZ18]. timestep [KSD11]. Toeplitz [BM19]. tomography [GS17, GWCH10, HHR11, HKQ18, HM18]. topography [BL17]. topological [ADF18]. Total [CP19, FLS10, HKQ18, JL13a, PSV18]. totally [Koe19]. TR [SW11a]. TR-meshes [SW11a]. trace [EH13, SAI17]. tracking [CC15, Run14]. traffic [BG16, CLM15]. transfer [Wor19]. transform [KK11]. Transformation [SCHH13, Fuk10, Sid11a]. transformed [PCCC18]. Transforming [BMP10]. transforms [GP11, VD11, WZH13]. translational [CP19]. transmission [CC13, GMSS11, QRSZ19]. Transmutation [EZ15]. transport [BCN19, SzS12, dGKL19]. transports [HAG17]. trapezium [ACWL14]. tree [Nou19]. tree-based [Nou19]. trees [BCM17]. Trefftz [Kre17, MP18b]. triangles [ÖS13]. triangular [BHK10]. triangulations [Dal10]. Tricomi [LT10]. tridiagonal [FPD12]. tridiagonalization [Sid11b]. trigonometric [CS12, ST18]. trivariate [Mat13]. truncation [Hac16]. truncations [Hac13]. TT-format [GK19a]. TT-rank [HRS12]. turnpike [HSLZ19]. Two [AV10, LMcS17, OS19, ABK13, BK12, BBB17, BGGH16, CCS17, Che16b, CHW17, CMR18, DHM19, DGS15, HQ12, JO10, KP17, KS18a, MOS11, NNT15, SS15, Sid11b, Wan12, ZWJ19]. two-coefficient [HQ12]. two-dimensional [ABK13]. two-level [BK12]. two-phase [CCS17, CHW17, DHM19]. two-point [ZWJ19]. Two-sided [AV10, BBB17, JO10]. twofold [HW11]. type [BBKS18, BRZ13, DMG19, Har15, HS17, HW13, KL18b, KVW15, LO14, OS19, OZ17, OZ23, SX13].

UA [BZ17]. Ultraconvergence [mHZZ17]. unbiased [CMR18]. Unconditional [HOR17a, SWS16, Wan18]. Unconditionally [HV12]. unification [VD11]. unified [AMN11, CDM^{+}18, HSV12]. Uniform [BCCH19, BHL11, DE16, FGL14, PSV10, BEG14, EK18, HM13, LS16]. Uniform-in-time [BCCH19, DE16]. Uniformly [CCLM15, BZ17, JZ14, MSW13, Sch16, Wor19]. unilateral [BMP10]. Uniquely [CHW17]. Uniqueness [PZ18]. unit [BD12b]. unitary [MV15, VD11]. units [BKTT10]. unity [Sch11a, Sch11b]. unknown [ABB19]. unrestricted [GST19]. Unsteady [BEG14, LL10]. unstructured [BDF12, GWX16, GGRG15, Mak18]. Unsymmetric [Sch10, FPD12]. upper [HHS15]. upscaling [MP15]. upwind [CKL14, SS18]. UQ [NTW18]. used [LT10]. using [BRK17, BO18, GN14, HR12, Kir11, KT12, NP15, PSV18, PS18, Pla15, SKW19, ZLL^{+}12].

value [Adc10, GW19, HQ12, HW18, Kre17, Mas16, OS14, ST18, ZZ15]. valued [GSY17, NTT16, SG… 12]. values [HU17]. variable [BC15b, DD16, mHZZ17]. variables [Sau17]. variance [BC15a, DFL16]. variation [CP19, FLS10, HKQ18, KR16, PS11]. Variational [BGN19, EHRS12, SW19, BK12, DGBL^{+}15, EMB10, HL15, HSD18, NvPZ10, SH19, Ste14, Ver17]. vector [DL11, HAG17]. vectors [Hac11, HAG17]. velocities [HHNS16]. velocity [CC15, HOR17a]. velocity-vorticity [HOR17a]. Verlet [HL16]. Vertex [ZZ15]. Vertex-centered [ZZ15]. vertices [Dal10, SW11a]. very [Loz19]. via [ACWL14, FY11, GW19, IN18, Jin12, LMSS17, NNT15, RI15, SDH^{+}14, Tow18]. vibration [dVMRR17]. viewpoint [vyKS14]. virtual [CGPS17, dVBMR16, dVMRR17, dVCMR18, dVMM19]. viscosity [DD16]. viscous [HS18, Ver10]. Vlasov [CCFM17, RS11]. Volterra [WZ10]. Volume [BHT17, AGS10, ABH13, BC12, BDF12, CCP13, Era15, EOS17, FGL14, FH17, GM14, HZZ18, LCH12, LC14, LMR16, Mur17, QRB11, SS15, SS18, XZ10, ZZ15, ZS17, vyKS14, BGHL14]. volume-nonconforming [SS15]. Voronoi [FGL14]. vorticity [HOR17a]. Vries [PSV10].

walls [CP14]. Wang [HH11]. water [BL17, BEG14, CKL14, CDKLM18, TT15]. wave [AY18, BPS10, BLM11b, BLS15, BO19, Cho18, CR12, EK18, ER18, GMÖ^{+}18, HR14, HHNS16, KK15c, KY12, Man15a, MP18b, MNT13, Ver10]. waveform [AL17, BGGH16, GW19]. wavefunctions [SY17]. waveguides [HK19a]. wavelet [DS10, Git14, MOGO17, Rei12, RS18, SS11]. Wavenumber [GMS19, GGS15]. Wavenumber-explicit [GMS19]. wavenumber-independent [GGS15]. wavepacket [GH14b]. waves [BKW10, IG15]. weak [BMWB14, LMY18, ZRK16]. Weakly [BC15b, AR16, FGHP17, LGL17]. weakly-singular [FGHP17]. weight [Maz11b, Not19, PS16]. weighted [CQR16, HSZ13, KY12, NOS16]. weights [DMN12, Dic12, HS19b]. Well [BG16, CDKLM18, BL17, Sch16]. Well-balanced [CDKLM18, BL17]. well-posed [Sch16]. Well-posedness [BG16]. which [GGS15]. white [GLW19, ZRK16]. whose [GV19]. Wideband [BKV15]. Willmore [BMO15]. Winther [CGG19]. without [BLM18, BC15b, MSV17, PZ18]. Worsey [Mat13].

XFEM [LR17]. Xu [HH11].

Yee [LS16]. yields [FHPS19].

Z [Che16a]. Z-grid [Che16a]. Zakharov [Gau18]. zero [Koe19]. zeros [Dri12, JO10, JJT13, NNT15, Seg13, WG13]. Zimmermann [Str15].

References

Attouch:2016:SCP

[ABAC16] Hédy Attouch, Luis M. Briceño-Arias, and Patrick L. Combettes. A strongly convergent primal-dual method for nonoverlapping domain decomposition. Numerische Mathematik, 133(3):443– 470, July 2016. CODEN NUMMA7. ISSN 0029-599X (print),

            0945-3245 (electronic). URL http://link.springer.com/
            accesspage/article/10.1007/s00211-015-0751-4;   http:
            //link.springer.com/article/10.1007/s00211-015-0751-4.

Arnold:2015:FED

[ABB15] Douglas N. Arnold, Daniele Boffi, and Francesca Bonizzoni. Finite element differential forms on curvilinear cubic meshes and their approximation properties. Numerische Mathematik, 129 (1):1–20, January 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-014-0631-3.

Adcock:2019:CUE

[ABB19] Ben Adcock, Anyi Bao, and Simone Brugiapaglia. Correcting for unknown errors in sparse high-dimensional function approximation. Numerische Mathematik, 142(3):667–711, July 2019. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
019-01051-9.

Arnold:2015:EAG

[ABC15] Martin Arnold, Olivier Brüls, and Alberto Cardona. Error analysis of generalized-α Lie group time integration methods for constrained mechanical systems. Numerische Mathematik, 129(1): 149–179, January 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-014-0633-1.

Anand:2010:AMS

[ABER10] Akash Anand, Yassine Boubendir, Fatih Ecevit, and Fernando Reitich. Analysis of multiple scattering iterations for highfrequency scattering problems. II: The three-dimensional scalar case. Numerische Mathematik, 114(3):373–427, January 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=114&issue=3&spage=
373.

Angelini:2013:FVM

[ABH13] Ophélie Angelini, Konstantin Brenner, and Danielle Hilhorst. A finite volume method on general meshes for a degenerate parabolic convection-reaction-diffusion equation. Numerische Mathematik, 123(2):219–257, February 2013. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-012-0485-5.

Antoine:2013:ABC

[ABK13] Xavier Antoine, Christophe Besse, and Pauline Klein. Absorbing boundary conditions for the two-dimensional Schrödinger equation with an exterior potential. Numerische Mathematik, 125(2): 191–223, October 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-013-0542-8.

Allonsius:2018:SAD

[ABM18] Damien Allonsius, Franck Boyer, and Morgan Morancey. Spectral analysis of discrete elliptic operators and applications in control theory. Numerische Mathematik, 140(4):857–911, December 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL https://link.springer.com/article/10.
1007/s00211-018-0983-1.

Alazah:2014:CFI

[ACWL14] Mohammad Alazah, Simon N. Chandler-Wilde, and Scott La Porte. Computing Fresnel integrals via modified trapezium rules. Numerische Mathematik, 128(4):635–661, December 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-014-0627-z.

Aregba-Driollet:2016:TAH

[ADBN16] Denise Aregba-Driollet, Maya Briani, and Roberto Natalini. Time asymptotic high order schemes for dissipative BGK hyperbolic systems. Numerische Mathematik, 132(2):399–431, February 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-015-0720-y.

Adcock:2010:MMF

[Adc10] Ben Adcock. Multivariate modified Fourier series and application to boundary value problems. Numerische Mathematik, 115(4):511–552, June 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=115&issue=4&spage=511.

Amstutz:2018:CRO

[ADF18] Samuel Amstutz, Charles Dapogny, and Àlex Ferrer. A consistent relaxation of optimal design problems for coupling shape and topological derivatives. Numerische Mathematik, 140(1): 35–94, September 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-018-0964-4.

Antonopoulos:2019:EEG

[ADM19] D. C. Antonopoulos, V. A. Dougalis, and D. E. Mitsotakis. Error estimates for Galerkin finite element methods for the Camassa– Holm equation. Numerische Mathematik, 142(4):833–862, August 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL https://link.springer.com/article/10.
1007/s00211-019-01045-7.

Arcangeli:2012:ESI

[AdST12] Rémi Arcangéli, María Cruz López de Silanes, and Juan José Torrens. Extension of sampling inequalities to Sobolev seminorms of fractional order and derivative data. Numerische Mathematik, 121(3):587–608, July 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=121&issue=3&spage=587.

Astorino:2010:CAP

[AG10] M. Astorino and C. Grandmont. Convergence analysis of a projection semi-implicit coupling scheme for fluid-structure interaction problems. Numerische Mathematik, 116(4):721– 767, October 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=116&issue=4&spage=721.

Ammari:2014:GPT

[AGK^{+}14] Habib Ammari, Josselin Garnier, Hyeonbae Kang, Mikyoung Lim, and Sanghyeon Yu. Generalized polarization tensors for shape description. Numerische Mathematik, 126(2):199– 224, February 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-013-0561-5.

Andreianov:2010:FVS

[AGS10] Boris Andreianov, Paola Goatin, and Nicolas Seguin. Finite volume schemes for locally constrained conservation laws. Numerische Mathematik, 115(4):609–645, June 2010. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=115&issue=4&spage=

Abdulle:2014:DGF

[AH14] A. Abdulle and M. E. Huber. Discontinuous Galerkin finite element heterogeneous multiscale method for advection-diffusion problems with multiple scales. Numerische Mathematik, 126 (4):589–633, April 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-013-0578-9.

Abdulle:2017:NHM

[AH17] A. Abdulle and M. E. Huber. Numerical homogenization method for parabolic advection–diffusion multiscale problems with large compressible flows. Numerische Mathematik, 136(3):603–649, July 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Amat:2014:FHO

[AHR14] S. Amat, M. A. Hernández, and N. Romero. On a family of high-order iterative methods under gamma conditions with applications in denoising. Numerische Mathematik, 127(2):201–221, June 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-013-0589-6.

Aishima:2015:GCR

[Ais15] Kensuke Aishima. Global convergence of the restarted Lanczos and Jacobi–Davidson methods for symmetric eigenvalue problems. Numerische Mathematik, 131(3):405–423, November 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-015-0699-4.

Alouges:2014:CPF

[AKST14] François Alouges, Evaggelos Kritsikis, Jutta Steiner, and Jean-Christophe Toussaint. A convergent and precise finite element scheme for Landau–Lifschitz–Gilbert equation. Numerische Mathematik, 128(3):407–430, November 2014. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-014-0615-3.

Ambartsumyan:2018:LMM

[AKYZ18] Ilona Ambartsumyan, Eldar Khattatov, Ivan Yotov, and Paolo Zunino. A Lagrange multiplier method for a Stokes–Biot fluidporoelastic structure interaction model. Numerische Mathematik, 140(2):513–553, October 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/s00211-018-0967-1.

Akrivis:2015:FIL

[AL15] Georgios Akrivis and Christian Lubich. Fully implicit, linearly implicit and implicit-explicit backward difference formulae for quasi-linear parabolic equations. Numerische Mathematik, 131(4):713–735, December 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-015-0702-0.

Antoine:2017:ASW

[AL17] X. Antoine and E. Lorin. An analysis of Schwarz waveform relaxation domain decomposition methods for the imaginarytime linear Schrödinger and Gross–Pitaevskii equations. Numerische Mathematik, 137(4):923–958, December 2017. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
017-0897-3.

Akrivis:2011:GRK

[AMN11] Georgios Akrivis, Charalambos Makridakis, and Ricardo H. Nochetto. Galerkin and Runge–Kutta methods: unified formulation, a posteriori error estimates and nodal superconvergence. Numerische Mathematik, 118(3):429–456, July 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=118&issue=3&spage=
429.

Anaya:2016:PPE

[AMORB16] Verónica Anaya, David Mora, Ricardo Oyarzúa, and Ricardo Ruiz-Baier. A priori and a posteriori error analysis of a mixed scheme for the Brinkman problem. Numerische Mathematik, 133(4):781–817, August 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/link.springer.com/accesspage/article/10.1007/s00211-
015-0758-x;   http://link.springer.com/article/10.1007/
s00211-015-0758-x.

Arnold:2018:SSE

[AN18] Anton Arnold and Claudia Negulescu. Stationary Schrödinger equation in the semi-classical limit: numerical coupling of oscillatory and evanescent regions. Numerische Mathematik, 138(2):501–536, February 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https:/

/link.springer.com/article/10.1007/s00211-017-0913-7;
 https://link.springer.com/content/pdf/10.1007/s00211-
017-0913-7.pdf.

Aceto:2019:RAF

[AN19] Lidia Aceto and Paolo Novati. Rational approximations to fractional powers of self-adjoint positive operators. Numerische Mathematik, 143(1):1–16, September 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-019-01048-4.

Allendes:2017:AFE

[AORS17] Alejandro Allendes, Enrique Otárola, Richard Rankin, and Abner J. Salgado. Adaptive finite element methods for an optimal control problem involving Dirac measures. Numerische Mathematik, 137(1):159–197, September 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Antonopoulou:2010:DGM

[AP10] D. C. Antonopoulou and M. Plexousakis. Discontinuous Galerkin methods for the linear Schrödinger equation in non-cylindrical domains. Numerische Mathematik, 115(4): 585–608, June 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=115&issue=4&spage=585.

Antonietti:2011:FEA

[AP11] Paola F. Antonietti and Aldo Pratelli. Finite element approximation of the Sobolev constant. Numerische Mathematik, 117(1):37–64, January 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=117&issue=1&spage=37.

Augustin:2012:SGT

[AR12] F. Augustin and P. Rentrop. Stochastic Galerkin techniques for random ordinary differential equations. Numerische Mathematik, 122(3):399–419, November 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=122&issue=3&spage=399.

Arens:2016:CMW

[AR16] Tilo Arens and Thomas Rösch. A collocation method for weakly singular integral equations with super-algebraic convergence rate. Numerische Mathematik, 134(3):441–472, November 2016. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s00211-015-0783-9;   http:
//link.springer.com/article/10.1007/s00211-015-0783-9.

Alqahtani:2019:GBA

[AR19] Hessah Alqahtani and Lothar Reichel. Generalized block anti-Gauss quadrature rules. Numerische Mathematik, 143(3):605– 648, November 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-019-01069-z.

Arbogast:2019:CCM

[AT19] Todd Arbogast and Zhen Tao. Construction of H(div)- conforming mixed finite elements on cuboidal hexahedra. Numerische Mathematik, 142(1):1–32, May 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL

https://link.springer.com/article/10.1007/s00211-018-
0998-7.

Absil:2010:TSG

[AV10] P.-A. Absil and P. Van Dooren. Two-sided Grassmann– Rayleigh quotient iteration. Numerische Mathematik, 114 (4):549–571, February 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=114&issue=4&spage=549.

Ainsworth:2011:FCR

[AV11] Mark Ainsworth and Tomás Vejchodský. Fully computable robust a posteriori error bounds for singularly perturbed reaction-diffusion problems. Numerische Mathematik, 119(2): 219–243, October 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=119&issue=2&spage=219.

Abdulle:2012:PEE

[AV12] Assyr Abdulle and Gilles Vilmart. A priori error estimates for finite element methods with numerical quadrature for nonmonotone nonlinear elliptic problems. Numerische Mathematik, 121(3):397–431, July 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=121&issue=3&spage=397.

Allaire:2018:ODC

[AY18] G. Allaire and T. Yamada. Optimization of dispersive coefficients in the homogenization of the wave equation in periodic structures. Numerische Mathematik, 140(2):265–326, October 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00211-018-0972-4.

Ben-Artzi:2019:SFB

[BAK19] Matania Ben-Artzi and Guy Katriel. Spline functions, the biharmonic operator and approximate eigenvalues. Numerische Mathematik, 141(4):839–879, April 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/s00211-018-01018-2.

Ballani:2012:FES

[Bal12] Jonas Ballani. Fast evaluation of singular BEM integrals based on tensor approximations. Numerische Mathematik, 121 (3):433–460, July 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=121&issue=3&spage=433.

Barlow:2013:RGK

[Bar13a] Jesse L. Barlow. Reorthogonalization for the Golub–Kahan– Lanczos bidiagonal reduction. Numerische Mathematik, 124(2): 237–278, June 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-013-0518-8.

Bartels:2013:FEA

[Bar13b] Sören Bartels. Finite element approximation of large bending isometries. Numerische Mathematik, 124(3):415–440, July 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-013-0519-7.

Barrios:2012:PEA

[BB12] Tomás P. Barrios and Rommel Bustinza. An a posteriori error analysis of an augmented discontinuous Galerkin formulation for Darcy flow. Numerische Mathematik, 120(2): 231–269, February 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=120&issue=2&spage=231.

Benner:2013:LRM

[BB13] Peter Benner and Tobias Breiten. Low rank methods for a class of generalized Lyapunov equations and related issues. Numerische Mathematik, 124(3):441–470, July 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.com/article/10.1007/s00211-013-0521-0.

Barrenechea:2017:LTS

[BBB17] G. R. Barrenechea, L. Boulton, and N. Boussaïd. Local two-sided bounds for eigenvalues of self-adjoint operators. Numerische Mathematik, 135(4):953–986, April 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/link.springer.com/article/10.1007/s00211-016-0822-1;
 http://link.springer.com/content/pdf/10.1007/s00211-
016-0822-1.pdf.

Betancourt:2018:RSM

[BBC^{+}18] Fernando Betancourt, Raimund Bürger, Christophe Chalons, Stefan Diehl, and Sebastian Farås. A random sampling method for a family of Temple-class systems of conservation laws. Numerische Mathematik, 138(1):37–73, January 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
017-0900-z.

Berselli:2016:CAF

[BBD16] Luigi C. Berselli, Dominic Breit, and Lars Diening. Convergence analysis for a finite element approximation of a steady model for electrorheological fluids. Numerische Mathematik, 132 (4):657–689, April 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-015-0735-4.

Barbeiro:2018:NFC

[BBFP18] Sílvia Barbeiro, Somayeh Gh. Bardeji, José A. Ferreira, and Luís Pinto. Non-Fickian convection-diffusion models in porous media. Numerische Mathematik, 138(4):869–904, April 2018. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
017-0922-6.

Barrenechea:2017:EBN

[BBK17] Gabriel R. Barrenechea, Erik Burman, and Fotini Karakatsani. Edge-based nonlinear diffusion for finite element approximations of convection-diffusion equations and its relation to algebraic flux-correction schemes. Numerische Mathematik, 135(2):521–545, February 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/link.springer.com/article/10.1007/s00211-016-0808-z;
 http://link.springer.com/content/pdf/10.1007/s00211-
016-0808-z.pdf.

Benner:2018:RLR

[BBKS18] Peter Benner, Zvonimir Bujanović, Patrick Kürschner, and Jens Saak. RADI: a low-rank ADI-type algorithm for large scale algebraic Riccati equations. Numerische Mathematik, 138(2):301–330, February 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https:/

/link.springer.com/article/10.1007/s00211-017-0907-5;
 https://link.springer.com/content/pdf/10.1007/s00211-
017-0907-5.pdf.

Babuska:2011:ENI

[BBL11] Ivo Babuska, Uday Banerjee, and Hengguang Li. The effect of numerical integration on the finite element approximation of linear functionals. Numerische Mathematik, 117 (1):65–88, January 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=117&issue=1&spage=65.

Ballani:2013:NSE

[BBSV13] J. Ballani, L. Banjai, S. Sauter, and A. Veit. Numerical solution of exterior Maxwell problems by Galerkin BEM and Runge– Kutta convolution quadrature. Numerische Mathematik, 123(4): 643–670, April 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-012-0503-7.

Bessemoulin-Chatard:2012:FVS

[BC12] Marianne Bessemoulin-Chatard. A finite volume scheme for convection-diffusion equations with nonlinear diffusion derived from the Scharfetter–Gummel scheme. Numerische Mathematik, 121(4):637–670, August 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=121&issue=4&spage=637.

Bierig:2015:CAM

[BC15a] Claudio Bierig and Alexey Chernov. Convergence analysis of multilevel Monte Carlo variance estimators and application for random obstacle problems. Numerische Mathematik, 130(4): 579–613, August 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-014-0676-3.

Bosing:2015:WPD

[BC15b] Paulo Rafael Bösing and Carsten Carstensen. Weakly overpenalized discontinuous Galerkin schemes for Reissner–Mindlin plates without the shear variable. Numerische Mathematik, 130 (3):395–423, July 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-014-0672-7.

Borm:2016:AIO

[BC16] Steffen Börm and Sven Christophersen. Approximation of integral operators by Green quadrature and nested cross approximation. Numerische Mathematik, 133(3):409–442, July 2016. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s00211-015-0757-y;   http:
//link.springer.com/article/10.1007/s00211-015-0757-y.

Bringmann:2017:ALS

[BC17] P. Bringmann and C. Carstensen. An adaptive least-squares FEM for the Stokes equations with optimal convergence rates. Numerische Mathematik, 135(2):459–492, February 2017. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s00211-016-0806-1;   http://link.springer.com/article/
10.1007/s00211-016-0806-1.

Bessemoulin-Chatard:2019:UTB

[BCCH19] M. Bessemoulin-Chatard and C. Chainais-Hillairet. Uniform-in-time bounds for approximate solutions of the drift–diffusion system. Numerische Mathematik, 141(4):881–916, April 2019. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
018-01019-1.

Brix:2015:NDG

[BCD15] Kolja Brix, Claudio Canuto, and Wolfgang Dahmen. Nested dyadic grids associated with Legendre–Gauss–Lobatto grids. Numerische Mathematik, 131(2):205–239, October 2015. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-014-0691-4.

Brenner:2011:MAS

[BCGS11] S. C. Brenner, J. Cui, T. Gudi, and L.-Y. Sung. Multigrid algorithms for symmetric discontinuous Galerkin methods on graded meshes. Numerische Mathematik, 119(1): 21–47, September 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=119&issue=1&spage=21.

Bogfjellmo:2017:HSL

[BCM17] Geir Bogfjellmo, Charles Curry, and Dominique Manchon. Hamiltonian B-series and a Lie algebra of non-rooted trees. Numerische Mathematik, 135(1):97–112, January 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/link.springer.com/accesspage/article/10.1007/s00211-
016-0792-3;   http://link.springer.com/article/10.1007/
s00211-016-0792-3.

Benamou:2016:DFI

[BCMO16] Jean-David Benamou, Guillaume Carlier, Quentin Mérigot, and Édouard Oudet. Discretization of functionals involving the Monge–Ampère operator. Numerische Mathematik, 134 (3):611–636, November 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/link.springer.com/accesspage/article/10.1007/s00211-
015-0781-y;   http://link.springer.com/article/10.1007/
s00211-015-0781-y.

Benamou:2019:GIF

[BCN19] Jean-David Benamou, Guillaume Carlier, and Luca Nenna. Generalized incompressible flows, multi-marginal transport and Sinkhorn algorithm. Numerische Mathematik, 142(1):33–54, May 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL https://link.springer.com/article/10.
1007/s00211-018-0995-x.

Benmansour:2010:DRM

[BCPS10] F. Benmansour, G. Carlier, G. Peyré, and F. Santambrogio. Derivatives with respect to metrics and applications: subgradient marching algorithm. Numerische Mathematik, 116 (3):357–381, September 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=116&issue=3&spage=357.

Bachmayr:2014:EEH

[BCS14a] Markus Bachmayr, Huajie Chen, and Reinhold Schneider. Error estimates for Hermite and even-tempered Gaussian approximations in quantum chemistry. Numerische Mathematik, 128(1): 137–165, September 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-014-0605-5.

Bokanowski:2014:DGS

[BCS14b] Olivier Bokanowski, Yingda Cheng, and Chi-Wang Shu. A discontinuous Galerkin scheme for front propagation with obstacles. Numerische Mathematik, 126(1):1–31, January 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-013-0555-3.

Bao:2019:RNM

[BCST19] Weizhu Bao, Rémi Carles, Chunmei Su, and Qinglin Tang. Regularized numerical methods for the logarithmic Schrödinger equation. Numerische Mathematik, 143(2):461–487, October 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00211-019-01058-2.

Baldeaux:2011:CPL

[BD11] Jan Baldeaux and Josef Dick. A construction of polynomial lattice rules with small gain coefficients. Numerische Mathematik, 119(2):271–297, October 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=119&issue=2&spage=271.

Bao:2012:ACN

[BD12a] Weizhu Bao and Xuanchun Dong. Analysis and comparison of numerical methods for the Klein–Gordon equation in the nonrelativistic limit regime. Numerische Mathematik, 120(2):189–229, February 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=120&issue=2&spage=189.

Brauchart:2012:QMC

[BD12b] Johann S. Brauchart and Josef Dick. Quasi-Monte Carlo rules for numerical integration over the unit sphere s^{2}. Numerische Mathematik, 121(3):473–502, July 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=121&issue=3&spage=473.

Buet:2012:DAP

[BDF12] Christophe Buet, Bruno Després, and Emmanuel Franck. Design of asymptotic preserving finite volume schemes for the hyperbolic heat equation on unstructured meshes. Numerische Mathematik, 122(2):227–278, October 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=122&issue=2&spage=227.

Bernardi:2018:FEM

[BDG^{+}18] Christine Bernardi, Séréna Dib, Vivette Girault, Frédéric Hecht, François Murat, and Toni Sayah. Finite element methods for Darcy’s problem coupled with the heat equation. Numerische Mathematik, 139(2):315–348, June 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-017-0938-y.

Bramwell:2012:LFD

[BDGQ12] Jamie Bramwell, Leszek Demkowicz, Jay Gopalakrishnan, and Weifeng Qiu. A locking-free hp DPG method for linear elasticity with symmetric stresses. Numerische Mathematik, 122(4): 671–707, December 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-012-0476-6.

Bos:2012:LCB

[BDHK12] Len Bos, Stefano De Marchi, Kai Hormann, and Georges Klein. On the Lebesgue constant of barycentric rational interpolation at equidistant nodes. Numerische Mathematik, 121 (3):461–471, July 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=121&issue=3&spage=461.

Bertoluzza:2018:LEA

[BDLM18] Silvia Bertoluzza, Astrid Decoene, Loïc Lacouture, and Sébastien Martin. Local error analysis for the Stokes equations with a punctual source term. Numerische Mathematik, 140(3):677– 701, November 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-018-0976-0.

Badea:2010:NAN

[BDQ10] Lori Badea, Marco Discacciati, and Alfio Quarteroni. Numerical analysis of the Navier–Stokes/Darcy coupling. Numerische Mathematik, 115(2):195–227, April 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=115&issue=2&spage=195.

Blanco:2011:MDH

[BDQ11] Pablo J. Blanco, Marco Discacciati, and Alfio Quarteroni. Modeling dimensionally-heterogeneous problems: analysis, approximation and applications. Numerische Mathematik, 119(2): 299–335, October 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=119&issue=2&spage=299.

Bruno:2013:CAH

[BDS13] Oscar P. Bruno, Víctor Domínguez, and Francisco-Javier Sayas. Convergence analysis of a high-order Nyström integral-equation method for surface scattering problems. Numerische Mathematik, 124(4):603–645, August 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-013-0525-9.

Bourdarias:2014:UMF

[BEG14] Christian Bourdarias, Mehmet Ersoy, and Stéphane Gerbi. Unsteady mixed flows in non uniform closed water pipes: a Full Kinetic Approach. Numerische Mathematik, 128(2):217–263, October 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-014-0611-7.

Bukal:2014:ESE

[BEJ14] Mario Bukal, Etienne Emmrich, and Ansgar Jüngel. Entropystable and entropy-dissipative approximations of a fourth-order quantum diffusion equation. Numerische Mathematik, 127(2): 365–396, June 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-013-0588-7.

Beyn:2011:CEI

[BEK11] Wolf-Jürgen Beyn, Cedric Effenberger, and Daniel Kressner. Continuation of eigenvalues and invariant pairs for parameterized nonlinear eigenvalue problems. Numerische Mathematik, 119(3):489–516, November 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=119&issue=3&spage=489.

Beltran:2011:CMS

[Bel11] Carlos Beltrán. A continuation method to solve polynomial systems and its complexity. Numerische Mathematik, 117(1):89–113, January 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=117&issue=1&spage=89.

Bertoluzza:2016:AMD

[Ber16] Silvia Bertoluzza. Analysis of a mesh-dependent stabilization for the three fields domain decomposition method. Numerische Mathematik, 133(1):1–36, May 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-015-0742-5.

Bambusi:2013:ESG

[BFG13] Dario Bambusi, Erwan Faou, and Benoît Grébert. Existence and stability of ground states for fully discrete approximations of the nonlinear Schrödinger equation. Numerische Mathematik, 123 (3):461–492, March 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-012-0491-7.

Brannick:2016:MPO

[BFK^{+}16] James Brannick, Andreas Frommer, Karsten Kahl, Björn Leder, Matthias Rottmann, and Artur Strebel. Multigrid preconditioning for the overlap operator in lattice QCD. Numerische Mathematik, 132(3):463–490, March 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-015-0725-6.

Beckermann:2012:SCR

[BG12] Bernhard Beckermann and Stefan Güttel. Superlinear convergence of the rational Arnoldi method for the approximation of matrix functions. Numerische Mathematik, 121 (2):205–236, June 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=121&issue=2&spage=205.

Blandin:2016:WPC

[BG16] Sebastien Blandin and Paola Goatin. Well-posedness of a conservation law with non-local flux arising in traffic flow modeling. Numerische Mathematik, 132(2):217–241, February 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-015-0717-6.

Boffi:2017:FDA

[BG17] Daniele Boffi and Lucia Gastaldi. A fictitious domain approach with Lagrange multiplier for fluid-structure interactions. Numerische Mathematik, 135(3):711–732, March 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s00211-016-0814-1;   http://link.springer.com/article/
10.1007/s00211-016-0814-1.

Brenner:2016:GDH

[BGG^{+}16] Konstantin Brenner, Mayya Groza, Cindy Guichard, Gilles Lebeau, and Roland Masson. Gradient discretization of hybrid dimensional Darcy flows in fractured porous media. Numerische Mathematik, 134(3):569–609, November 2016. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s00211-015-0782-x;   http://link.springer.com/article/
10.1007/s00211-015-0782-x.

Bennequin:2016:OSW

[BGGH16] Daniel Bennequin, Martin J. Gander, Loic Gouarin, and Laurence Halpern. Optimized Schwarz waveform relaxation for advection reaction diffusion equations in two dimensions. Numerische Mathematik, 134(3):513–567, November 2016. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s00211-015-0784-8;   http:
//link.springer.com/article/10.1007/s00211-015-0784-8.

Bousquet:2014:HOF

[BGHL14] Arthur Bousquet, Gung-Min Gie, Youngjoon Hong, and Jacques Laminie. A higher order Finite Volume resolution method for a system related to the inviscid primitive equations in a complex domain. Numerische Mathematik, 128(3):431–461, November 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-014-0622-4.

Banz:2017:SMH

[BGIS17] Lothar Banz, Heiko Gimperlein, Abderrahman Issaoui, and Ernst P. Stephan. Stabilized mixed hp-BEM for frictional contact problems in linear elasticity. Numerische Mathematik, 135(1):217–263, January 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/link.springer.com/accesspage/article/10.1007/s00211-
016-0797-y;   http://link.springer.com/article/10.1007/
s00211-016-0797-y.

BenBelgacem:2018:ALF

[BGJ18] F. Ben Belgacem, V. Girault, and F. Jelassi. Analysis of Lavrentiev-finite element methods for data completion problems. Numerische Mathematik, 139(1):1–25, May 2018. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
017-0930-6.

Barrett:2012:PAI

[BGN12] John W. Barrett, Harald Garcke, and Robert Nürnberg. Parametric approximation of isotropic and anisotropic elastic flow for closed and open curves. Numerische Mathematik, 120 (3):489–542, March 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=120&issue=3&spage=489.

Barrett:2016:SNM

[BGN16] John W. Barrett, Harald Garcke, and Robert Nürnberg. A stable numerical method for the dynamics of fluidic membranes. Numerische Mathematik, 134(4):783–822, December 2016. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.com/article/10.1007/s00211-
015-0787-5;   http://link.springer.com/content/pdf/10.
1007/s00211-015-0787-5.pdf.

Barrett:2019:VDA

[BGN19] John W. Barrett, Harald Garcke, and Robert Nürnberg. Variational discretization of axisymmetric curvature flows. Numerische Mathematik, 141(3):791–837, March 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https:/

/link.springer.com/article/10.1007/s00211-018-1013-z;

 https://link.springer.com/content/pdf/10.1007/s00211-
018-1013-z.pdf.

Bellavia:2018:LMM

[BGR18] Stefania Bellavia, Serge Gratton, and Elisa Riccietti. A Levenberg–Marquardt method for large nonlinear least-squares problems with dynamic accuracy in functions and gradients. Numerische Mathematik, 140(3):791–825, November 2018. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
018-0977-z.

Byckling:2011:AFI

[BH11] Mikko Byckling and Marko Huhtanen. Approximate factoring of the inverse. Numerische Mathematik, 117(3): 507–528, March 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=117&issue=3&spage=507.

Bause:2010:FOC

[BHK10] Markus Bause, Joachim Hoffmann, and Peter Knabner. Firstorder convergence of multi-point flux approximation on triangular grids and comparison with mixed finite element methods. Numerische Mathematik, 116(1):1–29, July 2010. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=116&issue=1&spage=
1.

Boyer:2011:UCP

[BHL11] Franck Boyer, Florence Hubert, and Jérôme Le Rousseau. Uniform controllability properties for space/time-discretized parabolic equations. Numerische Mathematik, 118(4):601–661, August 2011. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://www.springerlink.com/
openurl.asp?genre=article&issn=0029-599X&volume=118&issue=
4&spage=601.

Burman:2019:FEA

[BHL^{+}19] Erik Burman, Peter Hansbo, Mats G. Larson, Karl Larsson, and André Massing. Finite element approximation of the Laplace– Beltrami operator on a surface with boundary. Numerische Mathematik, 141(1):141–172, January 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https:/

/link.springer.com/article/10.1007/s00211-018-0990-2;
 https://link.springer.com/content/pdf/10.1007/s00211-
018-0990-2.pdf.

Burman:2016:CFE

[BHLZ16] Erik Burman, Peter Hansbo, Mats G. Larson, and Sara Zahedi. Cut finite element methods for coupled bulk-surface problems. Numerische Mathematik, 133(2):203–231, June 2016. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-015-0744-3.

Burman:2019:SCC

[BHLZ19] Erik Burman, Peter Hansbo, Mats G. Larson, and Sara Zahedi. Stabilized CutFEM for the convection problem on surfaces. Numerische Mathematik, 141(1):103–139, January 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL

https://link.springer.com/article/10.1007/s00211-018-
0989-8;   https://link.springer.com/content/pdf/10.1007/
s00211-018-0989-8.pdf.

BenChaabane:2017:NFV

[BHT17] Soumaya Ben Chaabane, Makram Hamouda, and Mahdi Tekitek. New Finite Volume Method for rotating channel flows involving boundary layers. Numerische Mathematik, 136(3):651–678, July 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bertoluzza:2011:AFD

[BIM11] Silvia Bertoluzza, Mourad Ismail, and Bertrand Maury. Analysis of the fully discrete fat boundary method. Numerische Mathematik, 118(1):49–77, May 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=118&issue=1&spage=49.

Bazhlekova:2015:ARS

[BJLZ15] Emilia Bazhlekova, Bangti Jin, Raytcho Lazarov, and Zhi Zhou. An analysis of the Rayleigh–Stokes problem for a generalized second-grade fluid. Numerische Mathematik, 131(1):1–31, September 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/link.springer.com/article/10.1007/s00211-014-0685-2;
 http://link.springer.com/content/pdf/10.1007/s00211-
014-0685-2.pdf.

Badea:2012:OTL

[BK12] L. Badea and R. Krause. One- and two-level Schwarz methods for variational inequalities of the second kind and their application to frictional contact. Numerische Mathematik, 120 (4):573–599, April 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=120&issue=4&spage=573.

Bonito:2013:TDH

[BKN13] Andrea Bonito, Irene Kyza, and Ricardo H. Nochetto. Timediscrete higher order ALE formulations: a priori error analysis. Numerische Mathematik, 125(2):225–257, October 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-013-0539-3.

Burger:2010:SOS

[BKTT10] Raimund Bürger, Kenneth H. Karlsen, Héctor Torres, and John D. Towers. Second-order schemes for conservation laws with discontinuous flux modelling clarifier-thickener units. Numerische Mathematik, 116(4):579–617, October 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=116&issue=4&spage=579.

Byrenheid:2017:TEB

[BKUV17] Glenn Byrenheid, Lutz Kämmerer, Tino Ullrich, and Toni Volkmer. Tight error bounds for rank-1 lattice sampling in spaces of hybrid mixed smoothness. Numerische Mathematik, 136(4): 993–1034, August 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bebendorf:2015:WNC

[BKV15] M. Bebendorf, C. Kuske, and R. Venn. Wideband nested cross approximation for Helmholtz problems. Numerische Mathematik, 130(1):1–34, May 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-014-0656-7.

Bouchut:2010:MAR

[BKW10] François Bouchut, Christian Klingenberg, and Knut Waagan. A multiwave approximate Riemann solver for ideal MHD based on relaxation II: numerical implementation with 3 and 5 waves. Numerische Mathematik, 115(4):647–679, June 2010. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=115&issue=4&spage=
647.

Bai:2013:MIA

[BL13a] Zhong-Zhi Bai and Xin-Guo Liu. On the Meany inequality with applications to convergence analysis of several row-action iteration methods. Numerische Mathematik, 124(2):215–236, June 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-012-0512-6.

Beyn:2013:EAH

[BL13b] Wolf-Jürgen Beyn and Alexander Lust. Error analysis of a hybrid method for computing Lyapunov exponents. Numerische Mathematik, 123(2):189–217, February 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-012-0486-4.

Bouchut:2017:MWB

[BL17] François Bouchut and Xavier Lhébrard. A multi well-balanced scheme for the shallow water MHD system with topography. Numerische Mathematik, 136(4):875–905, August 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Banjai:2019:EHO

[BLF19] L. Banjai and M. López-Fernández. Efficient high order algorithms for fractional integrals and fractional differential equations. Numerische Mathematik, 141(2):289–317, February 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00211-018-1004-0;   https://link.springer.com/content/
pdf/10.1007/s00211-018-1004-0.pdf.

Babenko:2011:SAA

[BLM11a] Yuliya Babenko, Tatyana Leskevich, and Jean-Marie Mirebeau. Sharp asymptotics of the L_{p} approximation error for interpolation on block partitions. Numerische Mathematik, 117(3):397–423, March 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=117&issue=3&spage=397.

Banjai:2011:RKC

[BLM11b] Lehel Banjai, Christian Lubich, and Jens Markus Melenk. Runge–Kutta convolution quadrature for operators arising in wave propagation. Numerische Mathematik, 119(1):1– 20, September 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=119&issue=1&spage=1.

Beckermann:2018:RFF

[BLM18] Bernhard Beckermann, George Labahn, and Ana C. Matos. On rational functions without Froissart doublets. Numerische Mathematik, 138(3):615–633, March 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/s00211-017-0917-3.

Bonito:2019:NAI

[BLP19] Andrea Bonito, Wenyu Lei, and Joseph E. Pasciak. Numerical approximation of the integral fractional Laplacian. Numerische Mathematik, 142(2):235–278, June 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-019-01025-x.

Brenner:2014:MMS

[BLS14] Susanne C. Brenner, Hengguang Li, and Li-Yeng Sung. Multigrid methods for saddle point problems: Stokes and Lamé systems. Numerische Mathematik, 128(2):193–216, October 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-014-0607-3.

Banjai:2015:SNC

[BLS15] Lehel Banjai, Christian Lubich, and Francisco-Javier Sayas. Stable numerical coupling of exterior and interior problems for the wave equation. Numerische Mathematik, 129(4):611–646, April 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-014-0650-0.

Bartels:2011:ECA

[BM11] Sören Bartels and Rüdiger Müller. Error control for the approximation of Allen–Cahn and Cahn–Hilliard equations with a logarithmic potential. Numerische Mathematik, 119 (3):409–435, November 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=119&issue=3&spage=409.

Baum:2013:NIP

[BM13] A. K. Baum and V. Mehrmann. Numerical integration of positive linear differential-algebraic systems. Numerische Mathematik, 124(2):279–307, June 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-013-0514-z.

Borm:2017:AHF

[BM17] Steffen Börm and Jens M. Melenk. Approximation of the highfrequency Helmholtz kernel by nested directional interpolation: error analysis. Numerische Mathematik, 137(1):1–34, September 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bini:2019:ESI

[BM19] Dario A. Bini and Beatrice Meini. On the exponential of semiinfinite quasi-Toeplitz matrices. Numerische Mathematik, 141(2): 319–351, February 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-018-1006-y.

Besse:2018:ABC

[BMGN18] Christophe Besse, Benoît Mésognon-Gireau, and Pascal Noble. Artificial boundary conditions for the linearized Benjamin–Bona– Mahony equation. Numerische Mathematik, 139(2):281–314, June 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL https://link.springer.com/article/
10.1007/s00211-017-0943-1.

Bansch:2011:PCF

[BMN11] Eberhard Bänsch, Pedro Morin, and Ricardo H. Nochetto. Preconditioning a class of fourth order problems by operator splitting. Numerische Mathematik, 118(2):197–228, June 2011. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=118&issue=2&spage=
197.

Bretin:2015:PFA

[BMO15] Elie Bretin, Simon Masnou, and Édouard Oudet. Phase-field approximations of the Willmore functional and flow. Numerische Mathematik, 131(1):115–171, September 2015. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-014-0683-4.

Bini:2010:TAR

[BMP10] Dario A. Bini, Beatrice Meini, and Federico Poloni. Transforming algebraic Riccati equations into unilateral quadratic matrix equations. Numerische Mathematik, 116(4):553–578, October 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=116&issue=4&spage=
553.

Beigel:2014:AWA

[BMWB14] Dörte Beigel, Mario S. Mommer, Leonard Wirsching, and Hans Georg Bock. Approximation of weak adjoints by reverse automatic differentiation of BDF methods. Numerische Mathematik, 126(3):383–412, March 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-013-0570-4.

Bokanowski:2010:CNM

[BMZ10] Olivier Bokanowski, Nadia Megdich, and Hasnaa Zidani. Convergence of a non-monotone scheme for Hamilton–Jacobi–Bellman equations with discontinous initial data. Numerische Mathematik, 115(1):1–44, March 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=115&issue=1&spage=1.

Bespalov:2016:BGM

[BN16] A. Bespalov and S. Nicaise. The BEM with graded meshes for the electric field integral equation on polyhedral surfaces. Numerische Mathematik, 132(4):631–655, April 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.com/article/10.1007/s00211-015-0736-3.

Brenner:2017:IPM

[BNRS17] Susanne C. Brenner, Michael Neilan, Armin Reiser, and Li-Yeng Sung. A C^{0} interior penalty method for a von Kármán plate. Numerische Mathematik, 135(3):803–832, March 2017. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s00211-016-0817-y;   http:
//link.springer.com/article/10.1007/s00211-016-0817-y.

Bank:2017:SRI

[BO17] Randolph E. Bank and Jeffrey S. Ovall. Some remarks on interpolation and best approximation. Numerische Mathematik, 137 (2):289–302, October 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-017-0877-7.

Burman:2018:DAH

[BO18] Erik Burman and Lauri Oksanen. Data assimilation for the heat equation using stabilized finite element methods. Numerische Mathematik, 139(3):505–528, July 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https:/

/link.springer.com/article/10.1007/s00211-018-0949-3;
 https://link.springer.com/content/pdf/10.1007/s00211-
018-0949-3.pdf.

Banjai:2019:PAF

[BO19] Lehel Banjai and Enrique Otárola. A PDE approach to fractional diffusion: a space-fractional wave equation. Numerische Mathematik, 143(1):177–222, September 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-019-01055-5.

Borm:2010:ASO

[Bör10] Steffen Börm. Approximation of solution operators of elliptic partial differential equations by H H- and H^{2}H^{2}-matrices.

Numerische Mathematik, 115(2):165–193, April 2010. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=115&issue=2&spage=
165.

Brenner:2018:MMS

[BOS18] Susanne C. Brenner, Duk-Soon Oh, and Li-Yeng Sung. Multigrid methods for saddle point problems: Darcy systems. Numerische Mathematik, 138(2):437–471, February 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
017-0911-9.

Banas:2010:FEA

[BPS10] L’ubomír Bañas, Andreas Prohl, and Reiner Schätzle. Finite element approximations of harmonic map heat flows and wave maps into spheres of nonconstant radii. Numerische Mathematik, 115(3):395–432, May 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=115&issue=3&spage=395.

Beuchler:2012:SOH

[BPZ12] Sven Beuchler, Veronika Pillwein, and Sabine Zaglmayr. Sparsity optimized high order finite element functions for H(div) on simplices. Numerische Mathematik, 122(2):197– 225, October 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=122&issue=2&spage=197.

Berlinet:2011:GIS

[BR11] Alain F. Berlinet and Christophe Roland. Geometric interpretation of some Cauchy related methods. Numerische Mathematik, 119(3):437–464, November 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=119&issue=3&spage=437.

Bosner:2017:QCA

[BR17] Tina Bosner and Mladen Rogina. Quadratic convergence of approximations by CCC–Schoenberg operators. Numerische Mathematik, 135(4):1253–1287, April 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/link.springer.com/accesspage/article/10.1007/s00211-
016-0831-0;   http://link.springer.com/article/10.1007/
s00211-016-0831-0.

Brenner:2013:AAM

[Bre13] Susanne C. Brenner. An additive analysis of multiplicative Schwarz methods. Numerische Mathematik, 123(1):1–19, January 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-012-0479-3.

Bause:2017:EAD

[BRK17] Markus Bause, Florin A. Radu, and Uwe Köcher. Error analysis for discretizations of parabolic problems using continuous finite elements in time and mixed finite elements in space. Numerische Mathematik, 137(4):773–818, December 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL

https://link.springer.com/article/10.1007/s00211-017-
0894-6;   https://link.springer.com/content/pdf/10.1007/
s00211-017-0894-6.pdf.

Brezinski:2013:PTR

[BRZ13] Claude Brezinski and Michela Redivo-Zaglia. Padé-type rational and barycentric interpolation. Numerische Mathematik, 125(1): 89–113, September 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-013-0535-7.

Burman:2010:BSD

[BS10] Erik Burman and Benjamin Stamm. Bubble stabilized discontinuous Galerkin method for parabolic and elliptic problems. Numerische Mathematik, 116(2):213–241, August 2010. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=116&issue=2&spage=
213.

Bermejo:2012:MLG

[BS12] Rodolfo Bermejo and Laura Saavedra. Modified Lagrange– Galerkin methods of first and second order in time for convection-diffusion problems. Numerische Mathematik, 120 (4):601–638, April 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=120&issue=4&spage=601.

Barlow:2013:RBC

[BS13] Jesse L. Barlow and Alicja Smoktunowicz. Reorthogonalized block classical Gram–Schmidt. Numerische Mathematik, 123(3): 395–423, March 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-012-0496-2.

Benzi:2017:AFL

[BS17] Michele Benzi and Valeria Simoncini. Approximation of functions of large matrices with Kronecker structure. Numerische Mathematik, 135(1):1–26, January 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/link.springer.com/accesspage/article/10.1007/s00211-
016-0799-9;   http://link.springer.com/article/10.1007/
s00211-016-0799-9.

Bressan:2019:AFD

[BS19] Andrea Bressan and Espen Sande. Approximation in FEM, DG and IGA: a theoretical comparison. Numerische Mathematik, 143(4):923–942, December 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/s00211-019-01063-5.

Bai:2012:NIE

[BSCZ12] Zheng-Jian Bai, Stefano Serra-Capizzano, and Zhi Zhao. Nonnegative inverse eigenvalue problems with partial eigendata. Numerische Mathematik, 120(3):387–431, March 2012. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=120&issue=3&spage=
387.

Barth:2011:MLM

[BSZ11] Andrea Barth, Christoph Schwab, and Nathaniel Zollinger. Multi-level Monte Carlo Finite Element method for elliptic PDEs with stochastic coefficients. Numerische Mathematik, 119(1):123–161, September 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=119&issue=1&spage=123.

Bardos:2015:SSC

[BT15] Claude Bardos and Eitan Tadmor. Stability and spectral convergence of Fourier method for nonlinear problems: on the shortcomings of the 2/3 de-aliasing method. Numerische Mathematik, 129 (4):749–782, April 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-014-0652-y.

Bebendorf:2012:CNB

[BV12] M. Bebendorf and R. Venn. Constructing nested bases approximations from the entries of non-local operators. Numerische Mathematik, 121(4):609–635, August 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=121&issue=4&spage=609.

Breda:2014:ALE

[BV14] Dimitri Breda and Erik Van Vleck. Approximating Lyapunov exponents and Sacker–Sell spectrum for retarded functional differential equations. Numerische Mathematik, 126(2):225– 257, February 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-013-0565-1.

Betcke:2017:RIP

[BV17] Marta M. Betcke and Heinrich Voss. Restarting iterative projection methods for Hermitian nonlinear eigenvalue problems with minmax property. Numerische Mathematik, 135 (2):397–430, February 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/link.springer.com/article/10.1007/s00211-016-0804-3;
 http://link.springer.com/content/pdf/10.1007/s00211-
016-0804-3.pdf.

Bank:2014:HHS

[BY14] Randolph E. Bank and Harry Yserentant. On the H^{1}-stability of the L_{2}-projection onto finite element spaces. Numerische Mathematik, 126(2):361–381, February 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-013-0562-4.

Bank:2015:NIB

[BY15] Randolph E. Bank and Harry Yserentant. A note on interpolation, best approximation, and the saturation property. Numerische Mathematik, 131(1):199–203, September 2015. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-014-0687-0.

Beale:2019:SDP

[BY19] J. Thomas Beale and Wenjun Ying. Solution of the Dirichlet problem by a finite difference analog of the boundary integral equation. Numerische Mathematik, 141(3):605–626, March 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00211-018-1010-2.

Bao:2017:UAU

[BZ17] Weizhu Bao and Xiaofei Zhao. A uniformly accurate (UA) multiscale time integrator Fourier pseudospectral method for the Klein–Gordon–Schrödinger equations in the nonrelativistic limit regime. Numerische Mathematik, 135(3):833–873, March 2017. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s00211-016-0818-x;   http:
//link.springer.com/article/10.1007/s00211-016-0818-x.

Calder:2017:NSR

[Cal17] Jeff Calder. Numerical schemes and rates of convergence for the Hamilton–Jacobi equation continuum limit of nondominated sorting. Numerische Mathematik, 137(4):819–856, December 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00211-017-0895-5.

Carstensen:2018:NDP

[CBHW18] C. Carstensen, P. Bringmann, F. Hellwig, and P. Wriggers. Nonlinear discontinuous Petrov–Galerkin methods. Numerische Mathematik, 139(3):529–561, July 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https:/

/link.springer.com/article/10.1007/s00211-018-0947-5.

Claeys:2010:AGS

[CC10a] X. Claeys and F. Collino. Augmented Galerkin schemes for the numerical solution of scattering by small obstacles. Numerische Mathematik, 116(2):243–268, August 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=116&issue=2&spage=243.

Clain:2010:SMM

[CC10b] Stéphane Clain and Vivien Clauzon. L^{∞} stability of the Muscl methods. Numerische Mathematik, 116(1):31–64, July 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=116&issue=1&spage=
31.

Chesnel:2013:CCG

[CC13] Lucas Chesnel and Patrick Ciarlet, Jr. T -coercivity and continuous Galerkin methods: application to transmission problems with sign changing coefficients. Numerische Mathematik, 124(1):1–29, May 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-012-0510-8.

Casas:2015:EED

[CC15] Eduardo Casas and Konstantinos Chrysafinos. Error estimates for the discretization of the velocity tracking problem. Numerische Mathematik, 130(4):615–643, August 2015. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-014-0680-7.

Casas:2019:NAQ

[CC19] Eduardo Casas and Konstantinos Chrysafinos. Numerical analysis of quasilinear parabolic equations under low regularity assumptions. Numerische Mathematik, 143(4):749–780, December 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL https://link.springer.com/article/10.
1007/s00211-019-01071-5.

Casas:2017:HOH

[CCFM17] Fernando Casas, Nicolas Crouseilles, Erwan Faou, and Michel Mehrenberger. High-order Hamiltonian splitting for the Vlasov– Poisson equations. Numerische Mathematik, 135(3):769–801, March 2017. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s00211-016-0816-z;   http:
//link.springer.com/article/10.1007/s00211-016-0816-z.

Charina:2017:RNS

[CCGP17] M. Charina, C. Conti, N. Guglielmi, and V. Protasov. Regularity of non-stationary subdivision: a matrix approach. Numerische Mathematik, 135(3):639–678, March 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/link.springer.com/article/10.1007/s00211-016-0809-y;
 http://link.springer.com/content/pdf/10.1007/s00211-
016-0809-y.pdf.

Chartier:2015:UAN

[CCLM15] Philippe Chartier, Nicolas Crouseilles, Mohammed Lemou, and Florian Méhats. Uniformly accurate numerical schemes for highly oscillatory Klein–Gordon and nonlinear Schrödinger equations. Numerische Mathematik, 129(2):211–250, February 2015. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-014-0638-9.

Cacace:2012:PEE

[CCM12] S. Cacace, A. Chambolle, and R. Monneau. A posteriori error estimates for the effective Hamiltonian of dislocation dynamics. Numerische Mathematik, 121(2):281–335, June 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=121&issue=2&spage=
281.

Cances:2013:MCG

[CCP13] Clément Cancès, Mathieu Cathala, and Christophe Le Potier. Monotone corrections for generic cell-centered finite volume approximations of anisotropic diffusion equations. Numerische Mathematik, 125(3):387–417, November 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.com/article/10.1007/s00211-013-0545-5.

Chen:2013:NTC

[CCQ13] Hongru Chen, Shaochun Chen, and Zhonghua Qiao. C^{0}- nonconforming tetrahedral and cuboid elements for the threedimensional fourth order elliptic problem. Numerische Mathematik, 124(1):99–119, May 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-012-0508-2.

Charina:2014:REP

[CCR14] Maria Charina, Costanza Conti, and Lucia Romani. Reproduction of exponential polynomials by multivariate non-stationary subdivision schemes with a general dilation matrix. Numerische Mathematik, 127(2):223–254, June 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.com/article/10.1007/s00211-013-0587-8.

Cai:2017:EET

[CCS17] Yongyong Cai, Heejun Choi, and Jie Shen. Error estimates for time discretizations of Cahn–Hilliard and Allen–Cahn phase-field models for two-phase incompressible flows. Numerische Mathematik, 137(2):417–449, October 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/s00211-017-0875-9.

Ciarlet:2016:IFE

[CCSS16] P. G. Ciarlet, P. Ciarlet, Jr., S. A. Sauter, and C. Simian. Intrinsic finite element methods for the computation of fluxes for Poisson’s equation. Numerische Mathematik, 132(3):433–462, March 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-015-0730-9.

Chen:2013:AAP

[CCZ13] Zhiming Chen, Tao Cui, and Linbo Zhang. An adaptive anisotropic perfectly matched layer method for 3-D time harmonic electromagnetic scattering problems. Numerische Mathematik, 125(4):639–677, December 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.com/article/10.1007/s00211-013-0550-8.

Casas:2011:EEN

[CD11] Eduardo Casas and Vili Dhamo. Error estimates for the numerical approximation of a quasilinear Neumann problem under minimal regularity of the data. Numerische Mathematik, 117(1):115–145, January 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=117&issue=1&spage=115.

Chertock:2018:WBS

[CDKLM18] Alina Chertock, Michael Dudzinski, Alexander Kurganov, and Mária Lukácová-Medvid’ová. Well-balanced schemes for the shallow water equations with Coriolis forces. Numerische Mathematik, 138(4):939–973, April 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/s00211-017-0928-0.

Cheng:2019:CAF

[CDL19] Hanz Martin Cheng, Jérôme Droniou, and Kim-Ngan Le. Convergence analysis of a family of ELLAM schemes for a fully coupled model of miscible displacement in porous media. Numerische Mathematik, 141(2):353–397, February 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
018-1002-2.

Cances:2018:GRP

[CDM^{+}18] Eric Cancès, Geneviève Dusson, Yvon Maday, Benjamin Stamm, and Martin Vohralík. Guaranteed and robust a posteriori bounds for Laplace eigenvalues and eigenvectors: a unified framework. Numerische Mathematik, 140(4):1033–1079, December 2018. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
018-0984-0.

Carstensen:2016:EAN

[CDNP16] Carsten Carstensen, Asha K. Dond, Neela Nataraj, and Amiya K. Pani. Error analysis of nonconforming and mixed FEMs for second-order linear non-selfadjoint and indefinite elliptic problems. Numerische Mathematik, 133(3):557–597, July 2016. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s00211-015-0755-0;   http:
//link.springer.com/article/10.1007/s00211-015-0755-0.

Cances:2014:NCA

[CEM14] Eric Cancès, Virginie Ehrlacher, and Yvon Maday. Nonconsistent approximations of self-adjoint eigenproblems: application to the supercell method. Numerische Mathematik, 128(4):

663–706, December 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-014-0626-0.

Chen:2011:CEP

[CFL11] Xiaojun Chen, Andreas Frommer, and Bruno Lang. Computational existence proofs for spherical t-designs. Numerische Mathematik, 117(2):289–305, February 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=117&issue=2&spage=289.

Carstensen:2011:OFA

[CG11] Carsten Carstensen and Joscha Gedicke. An oscillation-free adaptive FEM for symmetric eigenvalue problems. Numerische Mathematik, 118(3):401–427, July 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=118&issue=3&spage=401.

Carstensen:2014:GLE

[CG14] Carsten Carstensen and Dietmar Gallistl. Guaranteed lower eigenvalue bounds for the biharmonic equation. Numerische Mathematik, 126(1):33–51, January 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.com/article/10.1007/s00211-013-0559-z.

Carles:2017:FTS

[CG17] Rémi Carles and Clément Gallo. On Fourier time-splitting methods for nonlinear Schrödinger equations in the semi-classical limit II. Analytic regularity. Numerische Mathematik, 136(1):315–342, May 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Carstensen:2016:JSA

[CGG16] C. Carstensen, D. Gallistl, and J. Gedicke. Justification of the saturation assumption. Numerische Mathematik, 134(1):1–25, September 2016. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s00211-015-0769-7;   http:
//link.springer.com/article/10.1007/s00211-015-0769-7.

Carstensen:2019:RBP

[CGG19] Carsten Carstensen, Dietmar Gallistl, and Joscha Gedicke. Residual-based a posteriori error analysis for symmetric mixed Arnold–Winther FEM. Numerische Mathematik, 142(2):205–234, June 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL https://link.springer.com/article/
10.1007/s00211-019-01029-7;   https://link.springer.com/
content/pdf/10.1007/s00211-019-01029-7.pdf.

Carstensen:2013:PEE

[CGH13] Carsten Carstensen, Dietmar Gallistl, and Jun Hu. A posteriori error estimates for nonconforming finite element methods for fourth-order problems on rectangles. Numerische Mathematik, 124(2):309–335, June 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-012-0513-5.

Cao:2011:RRD

[CGHW11] Yanzhao Cao, Max Gunzburger, Xiaoming He, and Xiaoming Wang. Robin–Robin domain decomposition methods for the steady-state Stokes–Darcy system with the Beavers– Joseph interface condition. Numerische Mathematik, 117 (4):601–629, April 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=117&issue=4&spage=601.

Cotter:2018:MFE

[CGK18] Colin J. Cotter, P. Jameson Graber, and Robert C. Kirby. Mixed finite elements for global tide models with nonlinear damping. Numerische Mathematik, 140(4):963–991, December 2018. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
018-0980-4.

Carstensen:2011:AHA

[CGMM11] C. Carstensen, J. Gedicke, V. Mehrmann, and A. Miedlar. An adaptive homotopy approach for non-selfadjoint eigenvalue problems. Numerische Mathematik, 119(3):557– 583, November 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=119&issue=3&spage=557.

Carstensen:2014:AFE

[CGMM14] C. Carstensen, J. Gedicke, V. Mehrmann, and A. Miedlar. An adaptive finite element method with asymptotic saturation for eigenvalue problems. Numerische Mathematik, 128(4):615– 634, December 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-014-0624-2.

Cangiani:2017:PEE

[CGPS17] Andrea Cangiani, Emmanuil H. Georgoulis, Tristan Pryer, and Oliver J. Sutton. A posteriori error estimates for the virtual element method. Numerische Mathematik, 137 (4):857–893, December 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https:/

/link.springer.com/article/10.1007/s00211-017-0891-9;
 https://link.springer.com/content/pdf/10.1007/s00211-
017-0891-9.pdf.

Costea:2014:NHI

[CGS14] Adrian Costea, Heiko Gimperlein, and Ernst P. Stephan. A Nash– Hörmander iteration and boundary elements for the Molodensky problem. Numerische Mathematik, 127(1):1–34, May 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-013-0579-8.

Chen:2016:ELT

[CGSW16] Wenbin Chen, Max Gunzburger, Dong Sun, and Xiaoming Wang. An efficient and long-time accurate third-order algorithm for the Stokes–Darcy system. Numerische Mathematik, 134(4):857–879, December 2016. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s00211-015-0789-3;   http:
//link.springer.com/article/10.1007/s00211-015-0789-3.

Chouly:2012:NBD

[CH12] Franz Chouly and Norbert Heuer. A Nitsche-based domain decomposition method for hypersingular integral equations. Numerische Mathematik, 121(4):705–729, August 2012. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=121&issue=4&spage=

Christiansen:2018:GFE

[CH18a] Snorre H. Christiansen and Kaibo Hu. Generalized finite element systems for smooth differential forms and Stokes’ problem. Numerische Mathematik, 140(2):327–371, October 2018. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
018-0970-6.

Cirillo:2018:IAB

[CH18b] Emiliano Cirillo and Kai Hormann. An iterative approach to barycentric rational Hermite interpolation. Numerische Mathematik, 140(4):939–962, December 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-018-0986-y.

Colbrook:2019:IDQ

[CH19] Matthew J. Colbrook and Anders C. Hansen. On the infinite-dimensional QR algorithm. Numerische Mathematik, 143(1):17–83, September 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-019-01047-5;
 https://link.springer.com/content/pdf/10.1007/s00211-
019-01047-5.pdf.

Chavez:2011:DDS

[Chá11] Joseph Páez Chávez. Discretizing dynamical systems with generalized Hopf bifurcations. Numerische Mathematik, 118 (2):229–246, June 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=118&issue=2&spage=229.

Chen:2012:PSK

[Che12] Chin-Yun Chen. On the properties of Sard kernels and multiple error estimates for bounded linear functionals of bivariate functions with application to non-product cubature. Numerische Mathematik, 122(4):603–643, December 2012. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-012-0471-y.

Chen:2016:STN

[Che16a] Qingshan Chen. On staggering techniques and the non-staggered Z-grid scheme. Numerische Mathematik, 132(1):1–21, January 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-015-0715-8.

Chen:2016:EST

[Che16b] Xiao Shan Chen. On estimating the separation of two regular matrix pairs. Numerische Mathematik, 134(2):223–247, October 2016. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s00211-015-0775-9;   http:
//link.springer.com/article/10.1007/s00211-015-0775-9.

Christiansen:2018:NFE

[CHH18] Snorre H. Christiansen, Jun Hu, and Kaibo Hu. Nodal finite element de Rham complexes. Numerische Mathematik, 139(2): 411–446, June 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-017-0939-x.

Console:2013:SMM

[CHL13] Paola Console, Ernst Hairer, and Christian Lubich. Symmetric multistep methods for constrained Hamiltonian systems. Numerische Mathematik, 124(3):517–539, July 2013. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-013-0522-z.

Cho:2018:CBS

[Cho18] Chien-Hong Cho. On the computation for blow-up solutions of the nonlinear wave equation. Numerische Mathematik, 138(3): 537–556, March 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-017-0919-1.

Christiansen:2011:LRC

[Chr11] Snorre H. Christiansen. On the linearization of Regge calculus. Numerische Mathematik, 119(4):613–640, December 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?

genre=article&issn=0029-599X&volume=119&issue=4&spage=
613.

Cermak:2018:AII

[CHTV18] Martin Cermák, Frédéric Hecht, Zuqi Tang, and Martin Vohralík. Adaptive inexact iterative algorithms based on polynomialdegree-robust a posteriori estimates for the Stokes problem. Numerische Mathematik, 138(4):1027–1065, April 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
017-0925-3.

Casas:2012:ASC

[CHW12] Eduardo Casas, Roland Herzog, and Gerd Wachsmuth. Approximation of sparse controls in semilinear equations by piecewise linear functions. Numerische Mathematik, 122(4):645–669, December 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-012-0475-7.

Chen:2017:USE

[CHW17] Wenbin Chen, Daozhi Han, and Xiaoming Wang. Uniquely solvable and energy stable decoupled numerical schemes for the Cahn–Hilliard–Stokes–Darcy system for two-phase flows in karstic geometry. Numerische Mathematik, 137(1):229–255, September 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Cifani:2014:NME

[CJ14] Simone Cifani and Espen R. Jakobsen. On numerical methods and error estimates for degenerate fractional convection-diffusion equations. Numerische Mathematik, 127(3):447–483, July 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-013-0590-0.

Cotter:2016:MFE

[CK16] Colin J. Cotter and Robert C. Kirby. Mixed finite elements for global tide models. Numerische Mathematik, 133(2):255–277, June 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/link.springer.com/article/10.1007/s00211-015-0748-z;
 http://link.springer.com/content/pdf/10.1007/s00211-
015-0748-z.pdf.

Chertock:2014:CUS

[CKL14] Alina Chertock, Alexander Kurganov, and Yu Liu. Centralupwind schemes for the system of shallow water equations with horizontal temperature gradients. Numerische Mathematik, 127 (4):595–639, August 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-013-0597-6.

Carstensen:2019:CEE

[CLA19] Carsten Carstensen, Dongjie Liu, and Jochen Alberty. Convergence of dG(1) in elastoplastic evolution. Numerische Mathematik, 141(3):715–742, March 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/s00211-018-0999-6.

Cai:2019:AFD

[CLLS19] Wentao Cai, Buyang Li, Yanping Lin, and Weiwei Sun. Analysis of fully discrete FEM for miscible displacement in porous media with Bear–Scheidegger diffusion tensor. Numerische Mathematik, 141(4):1009–1042, April 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/s00211-019-01030-0.

Costeseque:2015:CSH

[CLM15] Guillaume Costeseque, Jean-Patrick Lebacque, and Régis Monneau. A convergent scheme for Hamilton–Jacobi equations on a junction: application to traffic. Numerische Mathematik, 129 (3):405–447, March 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-014-0643-z.

Chartier:2017:HOE

[CLM17] Philippe Chartier, Mohammed Lemou, and Florian Méhats. Highly-oscillatory evolution equations with multiple frequencies: averaging and numerics. Numerische Mathematik, 136(4):907– 939, August 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chan:2018:CNC

[CLMS18] Raymond Chan, Alessandro Lanza, Serena Morigi, and Fiorella Sgallari. Convex non-convex image segmentation. Numerische Mathematik, 138(3):635–680, March 2018. CODEN NUMMA7.

ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-017-0916-4.

Costantini:2010:GAH

[CM10] Paolo Costantini and Carla Manni. A geometric approach for Hermite subdivision. Numerische Mathematik, 115(3): 333–369, May 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=115&issue=3&spage=333.

Carstensen:2013:EPE

[CM13] C. Carstensen and C. Merdon. Effective postprocessing for equilibration a posteriori error estimators. Numerische Mathematik, 123(3):425–459, March 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-012-0494-4.

Cockburn:2015:HDG

[CM15] Bernardo Cockburn and Kassem Mustapha. A hybridizable discontinuous Galerkin method for fractional diffusion problems. Numerische Mathematik, 130(2):293–314, June 2015. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-014-0661-x.

Christof:2018:NPE

[CM18] Constantin Christof and Christian Meyer. A note on a priori Lp-error estimates for the obstacle problem. Numerische Mathematik, 139(1):27–45, May 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/s00211-017-0931-5.

Chartier:2014:MRC

[CMMV14] Philippe Chartier, Joseba Makazaga, Ander Murua, and Gilles Vilmart. Multi-revolution composition methods for highly oscillatory differential equations. Numerische Mathematik, 128(1): 167–192, September 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-013-0602-0.

Cindea:2013:APS

[CMP13] Nicolae Cîndea, Sorin Micu, and Jardel Morais Pereira. Approximation of periodic solutions for a dissipative hyperbolic equation. Numerische Mathematik, 124(3):559–601, July 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-013-0523-y.

Chouly:2018:UNA

[CMR18] Franz Chouly, Rabii Mlika, and Yves Renard. An unbiased Nitsche’s approximation of the frictional contact between two elastic structures. Numerische Mathematik, 139(3):593–631, July 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL https://link.springer.com/article/10.
1007/s00211-018-0950-x.

Clason:2019:BLI

[CN19] Christian Clason and Vu Huu Nhu. Bouligand–Landweber iteration for a non-smooth ill-posed problem. Numerische Mathematik, 142(4):789–832, August 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/s00211-019-01038-6.

Canuto:2017:COA

[CNSV17] Claudio Canuto, Ricardo H. Nochetto, Rob Stevenson, and Marco Verani. Convergence and optimality of hp-AFEM. Numerische Mathematik, 135(4):1073–1119, April 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/link.springer.com/accesspage/article/10.1007/s00211-
016-0826-x;   http://link.springer.com/article/10.1007/
s00211-016-0826-x.

Chen:2012:OMM

[CNX12] Long Chen, Ricardo H. Nochetto, and Jinchao Xu. Optimal multilevel methods for graded bisection grids. Numerische Mathematik, 120(1):1–34, January 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=120&issue=1&spage=1.

Cherfils:2014:NAC

[CP14] Laurence Cherfils and Madalina Petcu. A numerical analysis of the Cahn–Hilliard equation with non-permeable walls. Numerische Mathematik, 128(3):517–549, November 2014. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-014-0618-0.

Carstensen:2018:LOD

[CP18] Carsten Carstensen and Sophie Puttkammer. A low-order discontinuous Petrov–Galerkin method for the Stokes equations. Numerische Mathematik, 140(1):1–34, September 2018. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
018-0965-3.

Chambolle:2019:TRT

[CP19] Antonin Chambolle and Thomas Pock. Total roto-translational variation. Numerische Mathematik, 142(3):611–666, July 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00211-019-01026-w.

Carstensen:2017:CNA

[CPB17] Carsten Carstensen, Eun-Jae Park, and Philipp Bringmann. Convergence of natural adaptive least squares finite element methods. Numerische Mathematik, 136(4):1097–1115, August 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Carstensen:2013:OAN

[CPR13] Carsten Carstensen, Daniel Peterseim, and Hella Rabus. Optimal adaptive nonconforming FEM for the Stokes problem. Numerische Mathematik, 123(2):291–308, February 2013. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-012-0490-8.

Carstensen:2016:NDG

[CPS16] Carsten Carstensen, Daniel Peterseim, and Andreas Schröder. The norm of a discretized gradient in H(div)^{∗} for a posteriori finite element error analysis. Numerische Mathematik, 132(3): 519–539, March 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-015-0728-3.

Chen:2016:MWR

[CQR16] Peng Chen, Alfio Quarteroni, and Gianluigi Rozza. Multilevel and weighted reduced basis method for stochastic optimal control problems constrained by Stokes equations. Numerische Mathematik, 133(1):67–102, May 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-015-0743-4.

Cohen:2012:CNS

[CR12] David Cohen and Xavier Raynaud. Convergent numerical schemes for the compressible hyperelastic rod wave equation. Numerische Mathematik, 122(1):1–59, September 2012. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=122&issue=1&spage=
1.

Carstensen:2016:NNF

[CRS16] C. Carstensen, B. D. Reddy, and M. Schedensack. A natural nonconforming FEM for the Bingham flow problem is quasioptimal. Numerische Mathematik, 133(1):37–66, May 2016. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-015-0738-1.

Cohen:2012:CAT

[CS12] David Cohen and Magdalena Sigg. Convergence analysis of trigonometric methods for stiff second-order stochastic differential equations. Numerische Mathematik, 121(1):1–29, May 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=121&issue=1&spage=
1.

Chen:2017:RBP

[CS17] Huangxin Chen and Shuyu Sun. A residual-based a posteriori error estimator for single-phase Darcy flow in fractured porous media. Numerische Mathematik, 136(3):805–839, July 2017. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Carstensen:2016:OAF

[CSW16] Carsten Carstensen, Andreas Schröder, and Sebastian Wiedemann. An optimal adaptive finite element method for elastoplasticity. Numerische Mathematik, 132(1):131–154, January 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-015-0714-9.

Cox:2013:PHC

[CvN13] Sonja Cox and Jan van Neerven. Pathwise Hölder convergence of the implicit-linear Euler scheme for semi-linear SPDEs with multiplicative noise. Numerische Mathematik, 125(2):259–345, October 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-013-0538-4.

Chen:2012:MFE

[CW12] Wenbin Chen and Yanqiu Wang. A mixed finite element method for thin film epitaxy. Numerische Mathematik, 122(4):771– 793, December 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-012-0473-9.

Chandler-Wilde:2015:HFB

[CWHLT15] S. N. Chandler-Wilde, D. P. Hewett, S. Langdon, and A. Twigger. A high frequency boundary element method for scattering by a class of nonconvex obstacles. Numerische Mathematik, 129 (4):647–689, April 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-014-0648-7.

Chen:2010:CQO

[CXH10] Huangxin Chen, Xuejun Xu, and Ronald H. W. Hoppe. Convergence and quasi-optimality of adaptive nonconforming finite element methods for some nonsymmetric and indefinite problems. Numerische Mathematik, 116(3):383–419, September 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=116&issue=3&spage=
383.

Dalik:2010:ADD

[Dal10] Josef Dalík. Averaging of directional derivatives in vertices of nonobtuse regular triangulations. Numerische Mathematik, 116(4):619–644, October 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=116&issue=4&spage=619.

Deckers:2011:GEP

[DBV11] K. Deckers, A. Bultheel, and J. Van Deun. A generalized eigenvalue problem for quasi-orthogonal rational functions. Numerische Mathematik, 117(3):463–506, March 2011. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=117&issue=3&spage=
463.

Dhia:2018:MRF

[DCC18] Anne-Sophie Bonnet-Ben Dhia, Camille Carvalho, and Patrick Ciarlet, Jr. Mesh requirements for the finite element approximation of problems with sign-changing coefficients. Numerische Mathematik, 138(4):801–838, April 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-017-0923-5.

deCamargo:2017:SEF

[dCM17] André Pierro de Camargo and Walter F. Mascarenhas. The stability of extended Floater–Hormann interpolants. Numerische Mathematik, 136(1):287–313, May 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

deCastro:2016:FFR

[dCMT16] Pedro Machado Manhães de Castro, Quentin Mérigot, and Boris Thibert. Far-field reflector problem and intersection of paraboloids. Numerische Mathematik, 134(2):389–411, October 2016. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s00211-015-0780-z;   http:
//link.springer.com/article/10.1007/s00211-015-0780-z.

DeLosReyes:2016:EEO

[DD16] Juan Carlos De Los Reyes and Vili Dhamo. Error estimates for optimal control problems of a class of quasilinear equations arising in variable viscosity fluid flow. Numerische Mathematik, 132 (4):691–720, April 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-015-0737-2.

Droniou:2016:UTC

[DE16] Jérôme Droniou and Robert Eymard. Uniform-in-time convergence of numerical methods for non-linear degenerate parabolic equations. Numerische Mathematik, 132(4):721–766, April 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-015-0733-6.

Demlow:2016:QOA

[Dem16] Alan Demlow. Quasi-optimality of adaptive finite element methods for controlling local energy errors. Numerische Mathematik, 134(1):27–60, September 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/link.springer.com/accesspage/article/10.1007/s00211-
015-0774-x;   http://link.springer.com/article/10.1007/
s00211-015-0774-x.

Demlow:2017:CQO

[Dem17] Alan Demlow. Convergence and quasi-optimality of adaptive finite element methods for harmonic forms. Numerische Mathematik, 136(4):941–971, August 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Deckelnick:2011:NAI

[DES11] Klaus Deckelnick, Charles M. Elliott, and Vanessa Styles. Numerical analysis of an inverse problem for the eikonal equation. Numerische Mathematik, 119(2):245–269, October 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=119&issue=2&spage=
245.

Dang:2016:CEM

[DFL16] Duy-Minh Dang, Peter A. Forsyth, and Yuying Li. Convergence of the embedded mean-variance optimal points with discrete sampling. Numerische Mathematik, 132(2):271–302, February 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-015-0723-8.

Demoures:2015:DVL

[DGBL^{+}15] F. Demoures, F. Gay-Balmaz, S. Leyendecker, S. Ober-Blöbaum, T. S. Ratiu, and Y. Weinand. Discrete variational Lie group formulation of geometrically exact beam dynamics. Numerische Mathematik, 130(1):73–123, May 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.com/article/10.1007/s00211-014-0659-4.

deGournay:2019:DRS

[dGKL19] Frédéric de Gournay, Jonas Kahn, and Léo Lebrat. Differentiation and regularity of semi-discrete optimal transport with respect to the parameters of the discrete measure. Numerische Mathematik, 141(2):429–453, February 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
018-1000-4.

Dryja:2015:AFD

[DGS15] Maksymilian Dryja, Juan Galvis, and Marcus Sarkis. The analysis of a FETI–DP preconditioner for a full DG discretization of elliptic problems in two dimensions. Numerische Mathematik, 131(4):737–770, December 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-015-0705-x.

Dick:2017:CIP

[DGSY17] Josef Dick, Takashi Goda, Kosuke Suzuki, and Takehito Yoshiki. Construction of interlaced polynomial lattice rules for infinitely differentiable functions. Numerische Mathematik, 137(2):257– 288, October 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-017-0882-x.

Drouet:2017:ALA

[DH17] Guillaume Drouet and Patrick Hild. An accurate local average contact method for nonmatching meshes. Numerische Mathematik, 136(2):467–502, June 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dutta:2016:CFD

[DHKR16] Rajib Dutta, Helge Holden, Ujjwal Koley, and Nils Henrik Risebro. Convergence of finite difference schemes for the Benjamin– Ono equation. Numerische Mathematik, 134(2):249–274, October 2016. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s00211-015-0778-6;   http:
//link.springer.com/article/10.1007/s00211-015-0778-6.

Droniou:2019:NAT

[DHM19] Jérôme Droniou, Julian Hennicker, and Roland Masson. Numerical analysis of a two-phase flow discrete fracture matrix model. Numerische Mathematik, 141(1):21–62, January 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00211-018-0994-y.

deHoop:2015:AML

[dHQS15] Maarten V. de Hoop, Lingyun Qiu, and Otmar Scherzer. An analysis of a multi-level projected steepest descent iteration for nonlinear inverse problems in Banach spaces subject to stability constraints. Numerische Mathematik, 129(1):127–148, January 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-014-0629-x.

Dolz:2017:CRM

[DHS17] J. Dölz, H. Harbrecht, and Ch. Schwab. Covariance regularity and H-matrix approximation for rough random fields. Numerische Mathematik, 135(4):1045–1071, April 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/link.springer.com/accesspage/article/10.1007/s00211-
016-0825-y;   http://link.springer.com/article/10.1007/
s00211-016-0825-y.

Dick:2012:RWR

[Dic12] Josef Dick. Random weights, robust lattice rules and the geometry of the cbcrc algorithm. Numerische Mathematik, 122(3):443–467, November 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=122&issue=3&spage=443.

Dopico:2011:PTL

[DK11] Froilán M. Dopico and Plamen Koev. Perturbation theory for the LDU factorization and accurate computations for diagonally dominant matrices. Numerische Mathematik, 119(2):337–371, October 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=119&issue=2&spage=337.

Demlow:2016:MNP

[DK16a] Alan Demlow and Natalia Kopteva. Maximum-norm a posteriori error estimates for singularly perturbed elliptic reactiondiffusion problems. Numerische Mathematik, 133(4):707–742, August 2016. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s00211-015-0763-0;   http:
//link.springer.com/article/10.1007/s00211-015-0763-0.

Dryja:2016:MPN

[DK16b] Maksymilian Dryja and Piotr Krzyzanowski. A massively parallel nonoverlapping additive Schwarz method for discontinuous Galerkin discretization of elliptic problems. Numerische Mathematik, 132(2):347–367, February 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/link.springer.com/article/10.1007/s00211-015-0718-5;
 http://link.springer.com/content/pdf/10.1007/s00211-
015-0718-5.pdf.

Diening:2013:CHP

[DKS13] Lars Diening, Christian Kreuzer, and Sebastian Schwarzacher. Convex hull property and maximum principle for finite element minimisers of general convex functionals. Numerische Mathematik, 124(4):685–700, August 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-013-0527-7.

Dieci:2011:SMD

[DL11] Luca Dieci and Luciano Lopez. Sliding motion on discontinuity surfaces of high co-dimension. A construction for selecting a Filippov vector field. Numerische Mathematik, 117 (4):779–811, April 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=117&issue=4&spage=779.

Droniou:2015:GSL

[DL15] Jérôme Droniou and Bishnu P. Lamichhane. Gradient schemes for linear and non-linear elasticity equations. Numerische Mathematik, 129(2):251–277, February 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-014-0636-y.

Dopico:2018:BKL

[DLPV18] Froilán M. Dopico, Piers W. Lawrence, Javier Pérez, and Paul Van Dooren. Block Kronecker linearizations of matrix polynomials and their backward errors. Numerische Mathematik, 140(2): 373–426, October 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-018-0969-z.

Duan:2012:CEG

[DLT12] Huoyuan Duan, Ping Lin, and Roger C. E. Tan. C^{0} elements for generalized indefinite Maxwell equations. Numerische Mathematik, 122(1):61–99, September 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=122&issue=1&spage=61.

Duan:2013:ACF

[DLT13] Huoyuan Duan, Ping Lin, and Roger C. E. Tan. Analysis of a continuous finite element method for H(curl, div)-elliptic interface problem. Numerische Mathematik, 123(4):671–707, April 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-012-0500-x.

Digne:2014:NAD

[DM14] Julie Digne and Jean-Michel Morel. Numerical analysis of differential operators on raw point clouds. Numerische Mathematik, 127(2):255–289, June 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-013-0584-y.

Dedner:2016:ADG

[DM16] Andreas Dedner and Pravin Madhavan. Adaptive discontinuous Galerkin methods on surfaces. Numerische Mathematik, 132(2): 369–398, February 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-015-0719-4.

Dereich:2019:GMA

[DMG19] Steffen Dereich and Thomas Müller-Gronbach. General multilevel adaptations for stochastic approximation algorithms of Robbins– Monro and Polyak–Ruppert type. Numerische Mathematik, 142 (2):279–328, June 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-019-01024-y.

Drineas:2011:FLS

[DMMS11] Petros Drineas, Michael W. Mahoney, S. Muthukrishnan, and Tamás Sarlós. Faster least squares approximation. Numerische Mathematik, 117(2):219–249, February 2011. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=117&issue=2&spage=
219.

DeBonis:2012:GQR

[DMN12] M. C. De Bonis, G. Mastroianni, and I. Notarangelo. Gaussian quadrature rules with exponential weights on (−1, 1). Numerische Mathematik, 120(3):433–464, March 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=120&issue=3&spage=433.

Dick:2014:LRN

[DNP14] Josef Dick, Dirk Nuyens, and Friedrich Pillichshammer. Lattice rules for nonperiodic smooth integrands. Numerische Mathematik, 126(2):259–291, February 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-013-0566-0.

Discacciati:2017:CMF

[DO17] Marco Discacciati and Ricardo Oyarzúa. A conforming mixed finite element method for the Navier–Stokes/Darcy coupled problem. Numerische Mathematik, 135(2):571–606, February 2017. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s00211-016-0811-4;   http:
//link.springer.com/article/10.1007/s00211-016-0811-4.

Dopico:2016:SEC

[DP16] Froilán M. Dopico and Kenet Pomés. Structured eigenvalue condition numbers for parameterized quasiseparable matrices. Numerische Mathematik, 134(3):473–512, November 2016. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s00211-015-0779-5;   http:
//link.springer.com/article/10.1007/s00211-015-0779-5.

Driver:2012:IZG

[Dri12] Kathy Driver. Interlacing of zeros of Gegenbauer polynomials of non-consecutive degree from different sequences. Numerische Mathematik, 120(1):35–44, January 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=120&issue=1&spage=35.

Dijkema:2010:SLT

[DS10] Tammo Jan Dijkema and Rob Stevenson. A sparse Laplacian in tensor product wavelet coordinates. Numerische Mathematik, 115(3):433–449, May 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=115&issue=3&spage=433.

Demlow:2011:CQO

[DS11] Alan Demlow and Rob Stevenson. Convergence and quasioptimality of an adaptive finite element method for controlling L_{2} errors. Numerische Mathematik, 117(2):185– 218, February 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=117&issue=2&spage=185.

Davydov:2016:EBK

[DS16] Oleg Davydov and Robert Schaback. Error bounds for kernelbased numerical differentiation. Numerische Mathematik, 132(2): 243–269, February 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-015-0722-9.

Davydov:2018:MND

[DS18a] Oleg Davydov and Robert Schaback. Minimal numerical differentiation formulas. Numerische Mathematik, 140(3):555– 592, November 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-018-0973-3.

delaCalleYsern:2018:MSP

[dS18b] B. de la Calle Ysern and M. M. Spalević. Modified Stieltjes polynomials and Gauss–Kronrod quadrature rules. Numerische Mathematik, 138(1):1–35, January 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/s00211-017-0901-y.

Deckelnick:2018:SEA

[DS18c] Klaus Deckelnick and Vanessa Styles. Stability and error analysis for a diffuse interface approach to an advection– diffusion equation on a moving surface. Numerische Mathematik, 139(3):709–741, July 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https:/

/link.springer.com/article/10.1007/s00211-018-0946-6;
 https://link.springer.com/content/pdf/10.1007/s00211-
018-0946-6.pdf.

Du:2016:RBP

[DSX16] Shaohong Du, Shuyu Sun, and Xiaoping Xie. Residual-based a posteriori error estimation for multipoint flux mixed finite element methods. Numerische Mathematik, 134(1):197–222, September 2016. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s00211-015-0770-1;   http:
//link.springer.com/article/10.1007/s00211-015-0770-1.

Dione:2015:PFE

[DU15] Ibrahima Dione and José M. Urquiza. Penalty: finite element approximation of Stokes equations with slip boundary conditions. Numerische Mathematik, 129(3):587–610, March 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-014-0646-9.

DeMicheli:2011:NRL

[DV11] Enrico De Micheli and Giovanni A. Viano. Numerical recovery of location and residue of poles of meromorphic functions. Numerische Mathematik, 117(1):147–183, January 2011. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=117&issue=1&spage=
147.

daVeiga:2016:CVE

[dVBMR16] L. Beirão da Veiga, F. Brezzi, L. D. Marini, and A. Russo. H(div) and H(curl)-conforming virtual element methods. Numerische Mathematik, 133(2):303–332, June 2016. CODEN NUMMA7.

ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.com/article/10.1007/s00211-015-0746-1.

daVeiga:2011:SEH

[dVBRS11] L. Beirão da Veiga, A. Buffa, J. Rivas, and G. Sangalli. Some estimates for h-p-k-refinement in Isogeometric Analysis. Numerische Mathematik, 118(2):271–305, June 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=118&issue=2&spage=271.

daVeiga:2018:ECH

[dVCMR18] L. Beirão da Veiga, A. Chernov, L. Mascotto, and A. Russo. Exponential convergence of the hp virtual element method in presence of corner singularities. Numerische Mathematik, 138(3):581–613, March 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https:/

/link.springer.com/article/10.1007/s00211-017-0921-7;
 https://link.springer.com/content/pdf/10.1007/s00211-
017-0921-7.pdf.

daVeiga:2011:MDR

[dVM11] L. Beirão da Veiga and D. Mora. A mimetic discretization of the Reissner–Mindlin plate bending problem. Numerische Mathematik, 117(3):425–462, March 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=117&issue=3&spage=425.

daVeiga:2019:PEE

[dVMM19] L. Beirão da Veiga, G. Manzini, and L. Mascotto. A posteriori error estimation and adaptivity in hp virtual elements. Numerische Mathematik, 143(1):139–175, September 2019. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
019-01054-6;   https://link.springer.com/content/pdf/
10.1007/s00211-019-01054-6.pdf.

daVeiga:2017:VEM

[dVMRR17] Lourenço Beirão da Veiga, David Mora, Gonzalo Rivera, and Rodolfo Rodríguez. A virtual element method for the acoustic vibration problem. Numerische Mathematik, 136(3):725–763, July 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Diegel:2017:CAE

[DWWW17] Amanda E. Diegel, Cheng Wang, Xiaoming Wang, and Steven M. Wise. Convergence analysis and error estimates for a second order accurate finite element method for the Cahn–Hilliard–Navier– Stokes system. Numerische Mathematik, 137(3):495–534, November 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL https://link.springer.com/article/
10.1007/s00211-017-0887-5.

Egger:2019:SPA

[Egg19] H. Egger. Structure preserving approximation of dissipative evolution problems. Numerische Mathematik, 143(1):85–106, September 2019. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL https://link.springer.com/
article/10.1007/s00211-019-01050-w.

Elfverson:2015:MMP

[EGH15] Daniel Elfverson, Victor Ginting, and Patrick Henning. On multiscale methods in Petrov–Galerkin formulation. Numerische Mathematik, 131(4):643–682, December 2015. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.com/article/10.1007/s00211-
015-0703-z.

Estatico:2017:CGL

[EGLTP17] Claudio Estatico, Serge Gratton, Flavia Lenti, and David Titley-Peloquin. A conjugate gradient like method for p-norm minimization in functional spaces. Numerische Mathematik, 137(4): 895–922, December 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-017-0893-7.

Espig:2012:RNM

[EH12] Mike Espig and Wolfgang Hackbusch. A regularized Newton method for the efficient approximation of tensors represented in the canonical tensor format. Numerische Mathematik, 122(3):489–525, November 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=122&issue=3&spage=489.

Evans:2013:ETI

[EH13] John A. Evans and Thomas J. R. Hughes. Explicit trace inequalities for isogeometric analysis and parametric hexahedral finite elements. Numerische Mathematik, 123(2):259–290, February 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-012-0484-6.

Eisenmann:2018:CAD

[EH18] Monika Eisenmann and Eskil Hansen. Convergence analysis of domain decomposition based time integrators for degenerate parabolic equations. Numerische Mathematik, 140 (4):913–938, December 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https:/

/link.springer.com/article/10.1007/s00211-018-0985-z;
 https://link.springer.com/content/pdf/10.1007/s00211-
018-0985-z.pdf.

Espig:2012:VCS

[EHRS12] Mike Espig, Wolfgang Hackbusch, Thorsten Rohwedder, and Reinhold Schneider. Variational calculus with sums of elementary tensors of fixed rank. Numerische Mathematik, 122 (3):469–488, November 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=122&issue=3&spage=469.

Egger:2018:DWS

[EK18] H. Egger and T. Kugler. Damped wave systems on networks: exponential stability and uniform approximations. Numerische Mathematik, 138(4):839–867, April 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-017-0924-4.

Erb:2016:BLI

[EKAB16] Wolfgang Erb, Christian Kaethner, Mandy Ahlborg, and Thorsten M. Buzug. Bivariate Lagrange interpolation at the node points of non-degenerate Lissajous curves. Numerische Mathematik, 133(4):685–705, August 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/link.springer.com/accesspage/article/10.1007/s00211-
015-0762-1;   http://link.springer.com/article/10.1007/
s00211-015-0762-1.

Ekeland:2010:ACS

[EMB10] Ivar Ekeland and Santiago Moreno-Bromberg. An algorithm for computing solutions of variational problems with global convexity constraints. Numerische Mathematik, 115 (1):45–69, March 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=115&issue=1&spage=45.

Eikeland:2019:OSM

[EMR19] Erik Eikeland, Leszek Marcinkowski, and Talal Rahman. Overlapping Schwarz methods with adaptive coarse spaces for multiscale problems in 3D. Numerische Mathematik, 142(1):103–128, May 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https:/

/link.springer.com/article/10.1007/s00211-018-1008-9;
 https://link.springer.com/content/pdf/10.1007/s00211-
018-1008-9.pdf.

Engstrom:2014:SAQ

[Eng14] Christian Engström. Spectral approximation of quadratic operator polynomials arising in photonic band structure calculations. Numerische Mathematik, 126(3):413–440, March 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-013-0568-y.

Ecevit:2017:FAG

[EÖ17] Fatih Ecevit and Hasan Çagan Özen. Frequency-adapted Galerkin boundary element methods for convex scattering problems. Numerische Mathematik, 135(1):27–71, January 2017. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s00211-016-0800-7;   http://link.springer.com/article/
10.1007/s00211-016-0800-7.

Erath:2017:NSC

[EOS17] Christoph Erath, Günther Of, and Francisco-Javier Sayas. A non-symmetric coupling of the finite volume method and the boundary element method. Numerische Mathematik, 135(3):895– 922, March 2017. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s00211-016-0820-3;   http:
//link.springer.com/article/10.1007/s00211-016-0820-3.

Eigel:2017:ASG

[EPS17] Martin Eigel, Max Pfeffer, and Reinhold Schneider. Adaptive stochastic Galerkin FEM with hierarchical tensor representations. Numerische Mathematik, 136(3):765–803, July 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Elliott:2015:ESF

[ER15] Charles M. Elliott and Thomas Ranner. Evolving surface finite element method for the Cahn–Hilliard equation. Numerische Mathematik, 129(3):483–534, March 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-014-0644-y.

Egger:2018:SCP

[ER18] Herbert Egger and Bogdan Radu. Super-convergence and postprocessing for mixed finite element approximations of the wave equation. Numerische Mathematik, 140(2):427–447, October 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL https://link.springer.com/article/10.
1007/s00211-018-0966-2.

Erath:2015:NPE

[Era15] Christoph Erath. A nonconforming a posteriori estimator for the coupling of cell-centered finite volume and boundary element methods. Numerische Mathematik, 131(3):425–451, November 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-014-0694-1.

Ervedoza:2015:TTO

[EZ15] Sylvain Ervedoza and Enrique Zuazua. Transmutation techniques and observability for time-discrete approximation schemes of conservative systems. Numerische Mathematik, 130(3):425–466, July 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-014-0668-3.

Fernandez:2013:IDC

[Fer13a] Miguel A. Fernández. Incremental displacement-correction schemes for incompressible fluid-structure interaction. Numerische Mathematik, 123(1):21–65, January 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.com/article/10.1007/s00211-
012-0481-9.

Ferretti:2013:RBS

[Fer13b] Roberto Ferretti. On the relationship between Semi–Lagrangian and Lagrange–Galerkin schemes. Numerische Mathematik, 124 (1):31–56, May 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-012-0505-5.

Feischl:2015:SSN

[FFKP15] M. Feischl, T. Führer, M. Karkulik, and D. Praetorius. Stability of symmetric and nonsymmetric FEM–BEM couplings for nonlinear elasticity problems. Numerische Mathematik, 130(2):199–223, June 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-014-0662-9.

Feischl:2016:ABE

[FGH^{+}16] Michael Feischl, Gregor Gantner, Alexander Haberl, Dirk Praetorius, and Thomas Führer. Adaptive boundary element methods for optimal convergence of point errors. Numerische Mathematik, 132(3):541–567, March 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-015-0727-4.

Feischl:2017:OCA

[FGHP17] Michael Feischl, Gregor Gantner, Alexander Haberl, and Dirk Praetorius. Optimal convergence for adaptive IGA boundary element methods for weakly-singular integral equations. Numerische Mathematik, 136(1):147–182, May 2017. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL https://link.springer.com/content/pdf/10.
1007/s00211-016-0836-8.pdf.

Fiebach:2014:UGB

[FGL14] André Fiebach, Annegret Glitzky, and Alexander Linke. Uniform global bounds for solutions of an implicit Voronoi finite volume method for reaction–diffusion problems. Numerische Mathematik, 128(1):31–72, September 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-014-0604-6.

Faou:2010:BNFa

[FGP10a] Erwan Faou, Benoît Grébert, and Eric Paturel. Birkhoff normal form for splitting methods applied to semilinear Hamiltonian PDEs. Part I. Finite-dimensional discretization. Numerische Mathematik, 114(3):429–458, January 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=114&issue=3&spage=429.

Faou:2010:BNFb

[FGP10b] Erwan Faou, Benoît Grébert, and Eric Paturel. Birkhoff normal form for splitting methods applied to semilinear Hamiltonian PDEs. Part II. Abstract splitting. Numerische Mathematik, 114(3):459–490, January 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=114&issue=3&spage=459.

Filbet:2017:FVS

[FH17] Francis Filbet and Maxime Herda. A finite volume scheme for boundary-driven convection–diffusion equations with relative entropy structure. Numerische Mathematik, 137(3):535– 577, November 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-017-0885-7.

Fuselier:2014:KBQ

[FHN^{+}14] E. Fuselier, T. Hangelbroek, F. J. Narcowich, J. D. Ward, and G. B. Wright. Kernel based quadrature on spheres and other homogeneous spaces. Numerische Mathematik, 127(1):57–92, May 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-013-0581-1.

Fuhrer:2019:ABI

[FHPS19] Thomas Führer, Alexander Haberl, Dirk Praetorius, and Stefan Schimanko. Adaptive BEM with inexact PCG solver yields almost optimal computational costs. Numerische Mathematik, 141(4):967–1008, April 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https:/

/link.springer.com/article/10.1007/s00211-018-1011-1;
 https://link.springer.com/content/pdf/10.1007/s00211-
018-1011-1.pdf.

Feireisl:2016:CNM

[FKM16] Eduard Feireisl, Trygve Karper, and Martin Michálek. Convergence of a numerical method for the compressible Navier– Stokes system on general domains. Numerische Mathematik, 134(4):667–704, December 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/link.springer.com/accesspage/article/10.1007/s00211-
015-0786-6;   http://link.springer.com/article/10.1007/
s00211-015-0786-6.

Feistauer:2011:AST

[FKNP11] Miloslav Feistauer, Václav Kucera, Karel Najzar, and Jaroslava Prokopová. Analysis of space–time discontinuous Galerkin method for nonlinear convection-diffusion problems. Numerische Mathematik, 117(2):251–288, February 2011. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=117&issue=2&spage=
251.

Feischl:2018:FRF

[FKS18] Michael Feischl, Frances Y. Kuo, and Ian H. Sloan. Fast random field generation with H-matrices. Numerische Mathematik, 140(3):639–676, November 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/s00211-018-0974-2. See cor-

rection [FKS19].

Feischl:2019:CFR

[FKS19] Michael Feischl, Frances Y. Kuo, and Ian H. Sloan. Correction to: Fast random field generation with H-matrices. Numerische Mathematik, 142(3):787, July 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
019-01033-x;   https://link.springer.com/content/pdf/
10.1007/s00211-019-01033-x.pdf. See [FKS18].

Fermo:2015:NMB

[FL15] Luisa Fermo and Concetta Laurita. A Nyström method for a boundary integral equation related to the Dirichlet problem on domains with corners. Numerische Mathematik, 130(1):35–71, May 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-014-0657-6.

Ferraz-Leite:2012:NQE

[FLMP12] Samuel Ferraz-Leite, Jens Markus Melenk, and Dirk Praetorius. Numerical quadratic energy minimization bound to convex constraints in thin-film micromagnetics. Numerische Mathematik, 122(1):101–131, September 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=122&issue=1&spage=101.

Ferraz-Leite:2010:CSA

[FLOP10] S. Ferraz-Leite, C. Ortner, and D. Praetorius. Convergence of simple adaptive Galerkin schemes based on h–h/2 error estimators. Numerische Mathematik, 116(2):291–316, August 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=116&issue=2&spage=
291.

Fornasier:2010:COD

[FLS10] Massimo Fornasier, Andreas Langer, and Carola-Bibiane Schönlieb. A convergent overlapping domain decomposition method for total variation minimization. Numerische Mathematik, 116(4): 645–685, October 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=116&issue=4&spage=645.

Friedland:2019:GCN

[FLZ19] Shmuel Friedland, Lek-Heng Lim, and Jinjie Zhang. Grothendieck constant is norm of Strassen matrix multiplication tensor. Numerische Mathematik, 143(4):905–922, December 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL

https://link.springer.com/article/10.1007/s00211-019-
01070-6.

Faustmann:2018:LCB

[FM18] Markus Faustmann and Jens Markus Melenk. Local convergence of the boundary element method on polyhedral domains. Numerische Mathematik, 140(3):593–637, November 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL

https://link.springer.com/article/10.1007/s00211-018-
0975-1;   https://link.springer.com/content/pdf/10.1007/
s00211-018-0975-1.pdf.

Faustmann:2015:MAI

[FMP15] Markus Faustmann, Jens Markus Melenk, and Dirk Praetorius. H-matrix approximability of the inverses of FEM matrices. Numerische Mathematik, 131(4):615–642, December 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-015-0706-9.

Falco:2012:PGD

[FN12] Antonio Falcó and Anthony Nouy. Proper generalized decomposition for nonlinear convex problems in tensor Banach spaces. Numerische Mathematik, 121(3):503–530, July 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=121&issue=3&spage=
503.

Ferreira:2012:SEU

[FPD12] Carla Ferreira, Beresford Parlett, and Froilán M. Dopico. Sensitivity of eigenvalues of an unsymmetric tridiagonal matrix. Numerische Mathematik, 122(3):527–555, November 2012. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=122&issue=3&spage=
527.

Fornberg:2019:IGL

[FR19] Bengt Fornberg and Jonah A. Reeger. An improved Gregorylike method for 1-D quadrature. Numerische Mathematik, 141 (1):1–19, January 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-018-0992-0.

Fritz:2015:NRD

[Fri15] Hans Fritz. Numerical Ricci–DeTurck flow. Numerische Mathematik, 131(2):241–271, October 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-014-0690-5.

Froese:2018:MFD

[Fro18] Brittany D. Froese. Meshfree finite difference approximations for functions of the eigenvalues of the Hessian. Numerische Mathematik, 138(1):75–99, January 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/s00211-017-0898-2.

Faou:2014:APS

[FS14] Erwan Faou and Katharina Schratz. Asymptotic preserving schemes for the Klein–Gordon equation in the non-relativistic limit regime. Numerische Mathematik, 126(3):441–469, March 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-013-0567-z.

Fathi:2017:IDP

[FS17] M. Fathi and G. Stoltz. Improving dynamical properties of metropolized discretizations of overdamped Langevin dynamics. Numerische Mathematik, 136(2):545–602, June 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ferre:2019:EEE

[FS19] Grégoire Ferré and Gabriel Stoltz. Error estimates on ergodic properties of discretized Feynman–Kac semigroups. Numerische Mathematik, 143(2):261–313, October 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-019-01059-1.

Fukushima:2010:FCI

[Fuk10] Toshio Fukushima. Fast computation of incomplete elliptic integral of first kind by half argument transformation. Numerische Mathematik, 116(4):687–719, October 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=116&issue=4&spage=687.

Fukushima:2013:PFC

[Fuk13] Toshio Fukushima. Precise and fast computation of Jacobian elliptic functions by conditional duplication. Numerische Mathematik, 123(4):585–605, April 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-012-0498-0.

Fasshauer:2011:RKG

[FY11] Gregory E. Fasshauer and Qi Ye. Reproducing kernels of generalized Sobolev spaces via a Green function approach with distributional operators. Numerische Mathematik, 119 (3):585–611, November 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=119&issue=3&spage=585.

Greenhalgh:2013:PDP

[GAB13] Scott Greenhalgh, Vincent Acary, and Bernard Brogliato. On preserving dissipativity properties of linear complementarity dynamical systems with the θ-method. Numerische Mathematik, 125(4): 601–637, December 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-013-0553-5.

Gallistl:2015:OAF

[Gal15] Dietmar Gallistl. An optimal adaptive FEM for eigenvalue clusters. Numerische Mathematik, 130(3):467–496, July 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-014-0671-8.

Gantumur:2013:ABE

[Gan13] Tsogtgerel Gantumur. Adaptive boundary element methods with convergence rates. Numerische Mathematik, 124(3):471–516, July 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-013-0524-x.

Gauckler:2018:SMZ

[Gau18] Ludwig Gauckler. On a splitting method for the Zakharov system. Numerische Mathematik, 139(2):349–379, June 2018. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
017-0942-2.

Gray:2017:DTA

[GEEF17] W. Steven Gray, Luis A. Duffaut Espinosa, and Kurusch Ebrahimi-Fard. Discrete-time approximations of Fliess operators. Numerische Mathematik, 137(1):35–62, September 2017.

CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Giani:2012:AFE

[GG12] S. Giani and I. G. Graham. Adaptive finite element methods for computing band gaps in photonic crystals. Numerische Mathematik, 121(1):31–64, May 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=121&issue=1&spage=31.

Guillen-Gonzalez:2011:EEL

[GGGS11] Francisco Guillén-González and Juan Vicente Gutiérrez-Santacreu. Error estimates of a linear decoupled Euler–FEM scheme for a mass diffusion model. Numerische Mathematik, 117(2): 333–371, February 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=117&issue=2&spage=333.

Guillen-Gonzalez:2019:CMA

[GGGS19] Francisco Guillén-González and Juan Vicente Gutiérrez-Santacreu. From a cell model with active motion to a Hele–Shaw-like system: a numerical approach. Numerische Mathematik, 143(1): 107–137, September 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-019-01053-7.

Gilbert:2019:AQM

[GGK^{+}19] A. D. Gilbert, I. G. Graham, F. Y. Kuo, R. Scheichl, and I. H. Sloan. Analysis of quasi-Monte Carlo methods for elliptic eigenvalue problems with stochastic coefficients. Numerische Mathematik, 142(4):863–915, August 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/s00211-019-01046-6.

Gatica:2014:APB

[GGM14] Gabriel N. Gatica, Luis F. Gatica, and Antonio Márquez. Analysis of a pseudostress-based mixed finite element method for the Brinkman model of porous media flow. Numerische Mathematik, 126(4):635–677, April 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-013-0577-x.

Giani:2016:REE

[GGMO16] Stefano Giani, Luka Grubisić, Agnieszka Miedlar, and Jeffrey S. Ovall. Robust error estimates for approximations of non-self-adjoint eigenvalue problems. Numerische Mathematik, 133(3): 471–495, July 2016. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s00211-015-0752-3;   http:
//link.springer.com/article/10.1007/s00211-015-0752-3.

Guillen-Gonzalez:2015:AHS

[GGRG15] F. Guillén-González and J. R. Rodríguez-Galván. Analysis of the hydrostatic Stokes problem and finite-element approximation in unstructured meshes. Numerische Mathematik, 130(2):225–256, June 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-014-0663-8.

Gander:2015:AGH

[GGS15] M. J. Gander, I. G. Graham, and E. A. Spence. Applying GMRES to the Helmholtz equation with shifted Laplacian preconditioning: what is the largest shift for which wavenumber-independent convergence is guaranteed? Numerische Mathematik, 131(3): 567–614, November 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-015-0700-2.

Georgieva:2014:IHF

[GH14a] Irina Georgieva and Clemens Hofreither. Interpolation of harmonic functions based on Radon projections. Numerische Mathematik, 127(3):423–445, July 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-013-0592-y.

Gradinaru:2014:CSW

[GH14b] Vasile Gradinaru and George A. Hagedorn. Convergence of a semiclassical wavepacket based time-splitting for the Schrödinger equation. Numerische Mathematik, 126(1):53–73, January 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-013-0560-6.

Gallistl:2017:SRR

[GHP17] D. Gallistl, P. Huber, and D. Peterseim. On the stability of the Rayleigh–Ritz method for eigenvalues. Numerische Mathematik, 137(2):339–351, October 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/s00211-017-0876-8.

Gatica:2012:RHB

[GHS12] Gabriel N. Gatica, George C. Hsiao, and Francisco-Javier Sayas. Relaxing the hypotheses of Bielak–MacCamy’s BEM-FEM coupling. Numerische Mathematik, 120(3):465–487, March 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=120&issue=3&spage=
465.

Gittelson:2014:AWM

[Git14] Claude Jeffrey Gittelson. Adaptive wavelet methods for elliptic partial differential equations with random operators. Numerische Mathematik, 126(3):471–513, March 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.com/article/10.1007/s00211-013-0572-2.

Grasedyck:2019:SAT

[GK19a] Lars Grasedyck and Sebastian Krämer. Stable ALS approximation in the TT-format for rank-adaptive tensor completion. Numerische Mathematik, 143(4):855–904, December 2019. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
019-01072-4.

Guerand:2019:EEF

[GK19b] Jessica Guerand and Marwa Koumaiha. Error estimates for a finite difference scheme associated with Hamilton–Jacobi equations on a junction. Numerische Mathematik, 142(3):525–575, July 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL https://link.springer.com/article/10.
1007/s00211-019-01043-9;   https://link.springer.com/
content/pdf/10.1007/s00211-019-01043-9.pdf.

Guglielmi:2015:LRD

[GKL15] Nicola Guglielmi, Daniel Kressner, and Christian Lubich. Low rank differential equations for Hamiltonian matrix nearness problems. Numerische Mathematik, 129(2):279–319, February 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-014-0637-x.

Goudon:2019:DMN

[GKL19] Thierry Goudon, Stella Krell, and Giulia Lissoni. DDFV method for Navier–Stokes problem with outflow boundary conditions. Numerische Mathematik, 142(1):55–102, May 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
018-1014-y.

Graham:2015:QMC

[GKN^{+}15] I. G. Graham, F. Y. Kuo, J. A. Nichols, R. Scheichl, and Ch. Schwab. Quasi-Monte Carlo finite element methods for elliptic PDEs with lognormal random coefficients. Numerische Mathematik, 131(2):329–368, October 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-014-0689-y.

Graham:2018:CEQ

[GKN^{+}18] Ivan G. Graham, Frances Y. Kuo, Dirk Nuyens, Rob Scheichl, and Ian H. Sloan. Circulant embedding with QMC: analysis for elliptic PDE with lognormal coefficients. Numerische Mathematik, 140(2):479–511, October 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https:/

/link.springer.com/article/10.1007/s00211-018-0968-0;
 https://link.springer.com/content/pdf/10.1007/s00211-
018-0968-0.pdf.

Gittelson:2013:MLM

[GKSS13] Claude J. Gittelson, Juho Könnö, Christoph Schwab, and Rolf Stenberg. The multi-level Monte Carlo finite element method for a stochastic Brinkman Problem. Numerische Mathematik, 125(2): 347–386, October 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-013-0537-5.

Gameiro:2011:RCS

[GL11] Marcio Gameiro and Jean-Philippe Lessard. Rigorous computation of smooth branches of equilibria for the three dimensional Cahn–Hilliard equation. Numerische Mathematik, 117(4):753–778, April 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=117&issue=4&spage=753.

Gao:2017:SCF

[GLS17] Huadong Gao, Buyang Li, and Weiwei Sun. Stability and convergence of fully discrete Galerkin FEMs for the nonlinear thermistor equations in a nonconvex polygon. Numerische Mathematik, 136 (2):383–409, June 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gunzburger:2019:CFE

[GLW19] Max Gunzburger, Buyang Li, and Jilu Wang. Convergence of finite element solutions of stochastic partial integro-differential equations driven by white noise. Numerische Mathematik, 141 (4):1043–1077, April 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-019-01028-8.

Giesselmann:2014:GEF

[GM14] Jan Giesselmann and Thomas Müller. Geometric error of finite volume schemes for conservation laws on evolving surfaces. Numerische Mathematik, 128(3):489–516, November 2014. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-014-0621-5.

Gallouet:2019:EEI

[GMN19] Thierry Gallouët, David Maltese, and Antonin Novotny. Error estimates for the implicit MAC scheme for the compressible Navier–Stokes equations. Numerische Mathematik, 141(2): 495–567, February 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-018-1007-x.

Gimperlein:2018:BEM

[GMÖ^{+}18] Heiko Gimperlein, Fabian Meyer, Ceyhun Özdemir, David Stark, and Ernst P. Stephan. Boundary elements with mesh refinements for the wave equation. Numerische Mathematik, 139(4):867–912, August 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL https://link.springer.com/article/
10.1007/s00211-018-0954-6.

Garoni:2014:SSM

[GMP^{+}14] Carlo Garoni, Carla Manni, Francesca Pelosi, Stefano Serra-Capizzano, and Hendrik Speleers. On the spectrum of stiffness matrices arising from isogeometric analysis. Numerische Mathematik, 127(4):751–799, August 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-013-0600-2.

Galkowski:2019:WEA

[GMS19] Jeffrey Galkowski, Eike H. Müller, and Euan A. Spence. Wavenumber-explicit analysis for the Helmholtz h-BEM: error estimates and iteration counts for the Dirichlet problem. Numerische Mathematik, 142(2):329–357, June 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
019-01032-y;   https://link.springer.com/content/pdf/
10.1007/s00211-019-01032-y.pdf.

Gimperlein:2011:AFC

[GMSS11] Heiko Gimperlein, Matthias Maischak, Elmar Schrohe, and Ernst P. Stephan. Adaptive FE–BE coupling for strongly nonlinear transmission problems with Coulomb friction. Numerische Mathematik, 117(2):307–332, February 2011. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=117&issue=2&spage=
307.

Gratton:2011:SRB

[GMT11] Serge Gratton, Mélodie Mouffe, and Philippe L. Toint. Stopping rules and backward error analysis for bound-constrained optimization. Numerische Mathematik, 119(1):163–187, September 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0029-599X&volume=119&issue=1&spage=
163.

Guzman:2014:SCM

[GN14] Johnny Guzmán and Michael Neilan. Symmetric and conforming mixed finite elements for plane elasticity using rational bubble functions. Numerische Mathematik, 126(1):153–171, January 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-013-0557-1.

Girault:2015:MNE

[GNS15] V. Girault, R. H. Nochetto, and L. R. Scott. Max-norm estimates for Stokes and Navier–Stokes approximations in convex polyhedra. Numerische Mathematik, 131(4):771–822, December 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-015-0707-8.

Graf:2011:CSD

[GP11] Manuel Gräf and Daniel Potts. On the computation of spherical designs by a new optimization approach based on fast spherical Fourier transforms. Numerische Mathematik, 119 (4):699–724, December 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=119&issue=4&spage=699.

Gajardo:2013:SIC

[GS13] Pedro Gajardo and Alberto Seeger. Solving inverse coneconstrained eigenvalue problems. Numerische Mathematik, 123 (2):309–331, February 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-012-0487-3.

Gunther:2016:MGA

[GS16] Michael Günther and Adrian Sandu. Multirate generalized additive Runge–Kutta methods. Numerische Mathematik, 133(3): 497–524, July 2016. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s00211-015-0756-z;   http:
//link.springer.com/article/10.1007/s00211-015-0756-z.

Garde:2017:CRM

[GS17] Henrik Garde and Stratos Staboulis. Convergence and regularization for monotonicity-based shape reconstruction in electrical impedance tomography. Numerische Mathematik, 135(4):1221– 1251, April 2017. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s00211-016-0830-1;   http:
//link.springer.com/article/10.1007/s00211-016-0830-1.

Gil:2019:FRU

[GST19] Amparo Gil, Javier Segura, and Nico M. Temme. Fast, reliable and unrestricted iterative computation of Gauss–Hermite and Gauss–Laguerre quadratures. Numerische Mathematik, 143(3): 649–682, November 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-019-01066-2.

Gustafsson:2019:EAN

[GSV19] Tom Gustafsson, Rolf Stenberg, and Juha Videman. Error analysis of Nitsche’s mortar method. Numerische Mathematik, 142(4):973–994, August 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-019-01039-5;
 https://link.springer.com/content/pdf/10.1007/s00211-
019-01039-5.pdf.

Grohs:2017:SMV

[GSY17] Philipp Grohs, Markus Sprecher, and Thomas Yu. Scattered manifold-valued data approximation. Numerische Mathematik, 135(4):987–1010, April 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/link.springer.com/article/10.1007/s00211-016-0823-0;
 http://link.springer.com/content/pdf/10.1007/s00211-
016-0823-0.pdf.

Goda:2019:LRN

[GSY19] Takashi Goda, Kosuke Suzuki, and Takehito Yoshiki. Lattice rules in non-periodic subspaces of Sobolev spaces. Numerische Mathematik, 141(2):399–427, February 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
018-1003-1.

Garvie:2014:TLF

[GT14] Marcus R. Garvie and Catalin Trenchea. A three level finite element approximation of a pattern formation model in developmental biology. Numerische Mathematik, 127(3):397–422, July 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-013-0591-z.

Gopal:2019:RCM

[GT19] Abinand Gopal and Lloyd N. Trefethen. Representation of conformal maps by rational functions. Numerische Mathematik, 142 (2):359–382, June 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-019-01023-z.

Gigante:2015:AOG

[GV15] Giacomo Gigante and Christian Vergara. Analysis and optimization of the generalized Schwarz method for elliptic problems with application to fluid-structure interaction. Numerische Mathematik, 131(2):369–404, October 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-014-0693-2.

Gosse:2019:SEK

[GV19] Laurent Gosse and Nicolas Vauchelet. Some examples of kinetic schemes whose diffusion limit is Il’in’s exponential-fitting. Numerische Mathematik, 141(3):627–680, March 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
018-01020-8.

Girault:2014:MMF

[GVY14] Vivette Girault, Danail Vassilev, and Ivan Yotov. Mortar multiscale finite element methods for Stokes–Darcy flows. Numerische Mathematik, 127(1):93–165, May 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.com/article/10.1007/s00211-013-0583-z.

Gander:2019:CAP

[GW19] Martin J. Gander and Shu-Lin Wu. Convergence analysis of a periodic-like waveform relaxation method for initial-value problems via the diagonalization technique. Numerische Mathematik, 143(2):489–527, October 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/s00211-019-01060-8.

Gong:2010:NAS

[GWCH10] Rongfang Gong, Ge Wang, Xiaoliang Cheng, and Weimin Han. A novel approach for studies of multispectral bioluminescence tomography. Numerische Mathematik, 115(4):

553–583, June 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=115&issue=4&spage=553.

Guan:2014:CCS

[GWW14] Zhen Guan, Cheng Wang, and Steven M. Wise. A convergent convex splitting scheme for the periodic nonlocal Cahn– Hilliard equation. Numerische Mathematik, 128(2):377–406, October 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-014-0608-2.

Grasedyck:2016:NOM

[GWX16] Lars Grasedyck, Lu Wang, and Jinchao Xu. A nearly optimal multigrid method for general unstructured grids. Numerische Mathematik, 134(3):637–666, November 2016. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s00211-015-0785-7;   http://link.springer.com/article/
10.1007/s00211-015-0785-7.

Gong:2019:NHM

[GWX19] Shihua Gong, Shuonan Wu, and Jinchao Xu. New hybridized mixed methods for linear elasticity and optimal multilevel solvers. Numerische Mathematik, 141(2):569–604, February 2019. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
018-1001-3.

Gagelman:2012:SMS

[GY12] Jerry Gagelman and Harry Yserentant. A spectral method for Schrödinger equations with smooth confinement potentials. Numerische Mathematik, 122(2):383–398, October 2012. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=122&issue=2&spage=
383.

Gong:2017:AFE

[GY17] Wei Gong and Ningning Yan. Adaptive finite element method for elliptic optimal control problems: convergence and optimality. Numerische Mathematik, 135(4):1121–1170, April 2017. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s00211-016-0827-9;   http:
//link.springer.com/article/10.1007/s00211-016-0827-9.

Hackbusch:2011:TVT

[Hac11] Wolfgang Hackbusch. Tensorisation of vectors and their efficient convolution. Numerische Mathematik, 119(3):465– 488, November 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=119&issue=3&spage=465.

Hackbusch:2013:ETT

[Hac13] Wolfgang Hackbusch. L^{∞} estimation of tensor truncations. Numerische Mathematik, 125(3):419–440, November 2013. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-013-0544-6.

Hackbusch:2016:NER

[Hac16] Wolfgang Hackbusch. New estimates for the recursive low-rank truncation of block-structured matrices. Numerische Mathematik, 132(2):303–328, February 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-015-0716-7.

Huang:2017:IRT

[HAG17] Wen Huang, P.-A. Absil, and K. A. Gallivan. Intrinsic representation of tangent vectors and vector transports on matrix manifolds. Numerische Mathematik, 136(2):523–543, June 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Haji-Ali:2016:MIM

[HANT16] Abdul-Lateef Haji-Ali, Fabio Nobile, and Raúl Tempone. Multiindex Monte Carlo: when sparsity meets sampling. Numerische Mathematik, 132(4):767–806, April 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.com/article/10.1007/s00211-015-0734-5.

Hari:2015:CDF

[Har15] Vjeran Hari. Convergence to diagonal form of block Jacobi-type methods. Numerische Mathematik, 129(3):449–481, March 2015.

CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-014-0647-8.

Hardering:2018:DEB

[Har18] Hanne Hardering. L^{2}-discretization error bounds for maps into Riemannian manifolds. Numerische Mathematik, 139(2):381–410, June 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL https://link.springer.com/article/
10.1007/s00211-017-0941-3.

He:2013:EIE

[He13] Yinnian He. Euler implicit/explicit iterative scheme for the stationary Navier–Stokes equations. Numerische Mathematik, 123 (1):67–96, January 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-012-0482-8.

Horvath:2019:QPD

[HFK19] Róbert Horváth, István Faragó, and János Karátson. Qualitative properties of discrete nonlinear parabolic operators. Numerische Mathematik, 143(3):529–554, November 2019. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
019-01062-6;   https://link.springer.com/content/pdf/
10.1007/s00211-019-01062-6.pdf.

Huang:2012:IMS

[HFL12] Y. Huang, P. A. Forsyth, and G. Labahn. Iterative methods for the solution of a singular control formulation of a GMWB pricing problem. Numerische Mathematik, 122(1): 133–167, September 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=122&issue=1&spage=133.

Hochbruck:2010:CRL

[HH10] Marlis Hochbruck and Michael Hönig. On the convergence of a regularizing Levenberg–Marquardt scheme for nonlinear ill-posed problems. Numerische Mathematik, 115(1): 71–79, March 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=115&issue=1&spage=71.

Huang:2011:LPA

[HH11] Jianguo Huang and Xuehai Huang. Local and parallel algorithms for fourth order problems discretized by the Morley– Wang–Xu element method. Numerische Mathematik, 119 (4):667–697, December 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=119&issue=4&spage=667.

Harwood:2016:ESO

[HHB16] Stuart M. Harwood, Kai Höffner, and Paul I. Barton. Efficient solution of ordinary differential equations with a parametric lexicographic linear program embedded. Numerische Mathematik, 133(4):623–653, August 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/link.springer.com/accesspage/article/10.1007/s00211-
015-0760-3;   http://link.springer.com/article/10.1007/
s00211-015-0760-3.

Halla:2016:HSI

[HHNS16] Martin Halla, Thorsten Hohage, Lothar Nannen, and Joachim Schöberl. Hardy space infinite elements for time harmonic wave equations with phase and group velocities of different signs. Numerische Mathematik, 133(1):103–139, May 2016. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-015-0739-0.

Hanke:2011:CBS

[HHR11] Martin Hanke, Nuutti Hyvönen, and Stefanie Reusswig. Convex backscattering support in electric impedance tomography. Numerische Mathematik, 117(2):373–396, February 2011. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=117&issue=2&spage=
373.

Hu:2015:CBL

[HHS15] Jun Hu, Yunqing Huang, and Quan Shen. Constructing both lower and upper bounds for the eigenvalues of elliptic operators by nonconforming finite element methods. Numerische Mathematik, 131(2):273–302, October 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-014-0688-z.

He:2019:GGI

[HHX19] Juncai He, Kaibo Hu, and Jinchao Xu. Generalized Gaffney inequality and discrete compactness for discrete differential forms. Numerische Mathematik, 143(4):781–795, December 2019. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
019-01076-0.

Han:2014:CDR

[HHYY14] Deren Han, Hongjin He, Hai Yang, and Xiaoming Yuan. A customized Douglas–Rachford splitting algorithm for separable convex minimization with linear constraints. Numerische Mathematik, 127(1):167–200, May 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-013-0580-2.

Henriquez:2017:BIF

[HJHA17] Fernando Henríquez, Carlos Jerez-Hanckes, and Fernando Altermatt. Boundary integral formulation and semi-implicit scheme coupling for modeling cells under electrical stimulation. Numerische Mathematik, 136(1):101–145, May 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hsia:2017:TPS

[HJNS17] Chun-Hsiung Hsia, Chang-Yeol Jung, Thien Binh Nguyen, and Ming-Cheng Shiue. On time periodic solutions, asymptotic stability and bifurcations of Navier–Stokes equations. Numerische Mathematik, 135(2):607–638, February 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s00211-016-0812-3;   http://link.springer.com/article/
10.1007/s00211-016-0812-3.

Hochbruck:2015:CAS

[HJS15] Marlis Hochbruck, Tobias Jahnke, and Roland Schnaubelt. Convergence of an ADI splitting for Maxwell’s equations. Numerische Mathematik, 129(3):535–561, March 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.com/article/10.1007/s00211-014-0642-0.

Hong:2014:NAS

[HJT14] Youngjoon Hong, Chang-Yeol Jung, and Roger Temam. On the numerical approximations of stiff convection-diffusion equations in a circle. Numerische Mathematik, 127(2):291–313, June 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-013-0585-x.

Hakopian:2014:GMC

[HJZ14] Hakop Hakopian, Kurt Jetter, and Georg Zimmermann. The Gasca–Maeztu conjecture for n = 5. Numerische Mathematik, 127(4):685–713, August 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-013-0599-4.

Heuer:2017:DPG

[HK17] Norbert Heuer and Michael Karkulik. Discontinuous Petrov– Galerkin boundary elements. Numerische Mathematik, 135 (4):1011–1043, April 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/link.springer.com/accesspage/article/10.1007/s00211-
016-0824-z;   http://link.springer.com/article/10.1007/
s00211-016-0824-z.

Hagstrom:2019:CRB

[HK19a] Thomas Hagstrom and Seungil Kim. Complete radiation boundary conditions for the Helmholtz equation I: waveguides. Numerische Mathematik, 141(4):917–966, April 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
018-1012-0.

Hofstatter:2019:NSP

[HK19b] Harald Hofstätter and Othmar Koch. Non-satisfiability of a positivity condition for commutator-free exponential integrators of order higher than four. Numerische Mathematik, 141(3):681–691, March 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL https://link.springer.com/article/
10.1007/s00211-018-1015-x.

Hannukainen:2012:MAC

[HKK12] Antti Hannukainen, Sergey Korotov, and Michal Krízek. The maximum angle condition is not necessary for convergence of the finite element method. Numerische Mathematik, 120 (1):79–88, January 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=120&issue=1&spage=79.

Hinze:2018:ICE

[HKQ18] Michael Hinze, Barbara Kaltenbacher, and Tran Nhan Tam Quyen. Identifying conductivity in electrical impedance tomography with total variation regularization. Numerische Mathematik, 138(3):723–765, March 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/s00211-017-0920-8.

Hofstatter:2014:CAH

[HKT14] Harald Hofstätter, Othmar Koch, and Mechthild Thalhammer. Convergence analysis of high-order time-splitting pseudo-spectral methods for rotational Gross–Pitaevskii equations. Numerische Mathematik, 127(2):315–364, June 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.com/article/10.1007/s00211-013-0586-9.

Hong:2016:RMM

[HKXZ16] Qingguo Hong, Johannes Kraus, Jinchao Xu, and Ludmil Zikatanov. A robust multigrid method for discontinuous Galerkin discretizations of Stokes and linear elasticity equations. Numerische Mathematik, 132(1):23–49, January 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-015-0712-y.

Hall:2015:SVI

[HL15] James Hall and Melvin Leok. Spectral variational integrators. Numerische Mathematik, 130(4):681–740, August 2015. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-014-0679-0.

Hairer:2016:LTA

[HL16] Ernst Hairer and Christian Lubich. Long-term analysis of the Störmer–Verlet method for Hamiltonian systems with a solution-dependent high frequency. Numerische Mathematik, 134(1):119–138, September 2016. CODEN NUMMA7.

ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/link.springer.com/accesspage/article/10.1007/s00211-
015-0766-x;   http://link.springer.com/article/10.1007/
s00211-015-0766-x.

Hu:2018:CRA

[HL18] Shenglong Hu and Guoyin Li. Convergence rate analysis for the higher order power method in best rank one approximations of tensors. Numerische Mathematik, 140(4):993–1031, December 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL https://link.springer.com/article/10.
1007/s00211-018-0981-3.

Hakula:2019:ACS

[HL19] Harri Hakula and Mikael Laaksonen. Asymptotic convergence of spectral inverse iterations for stochastic eigenvalue problems. Numerische Mathematik, 142(3):577–609, July 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
019-01034-w;   https://link.springer.com/content/pdf/
10.1007/s00211-019-01034-w.pdf.

Huang:2011:PQS

[HLS11] Tsung-Ming Huang, Wen-Wei Lin, and Wei-Shuo Su. Palindromic quadratization and structure-preserving algorithm for palindromic matrix polynomials of even degree. Numerische Mathematik, 118(4):713–735, August 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=118&issue=4&spage=713.

Hiptmair:2012:CAF

[HLZ12] Ralf Hiptmair, Jingzhi Li, and Jun Zou. Convergence analysis of finite element methods for H(curl; Ω)-elliptic interface problems. Numerische Mathematik, 122(3):557–578, November 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=122&issue=3&spage=
557.

Heuer:2013:DGB

[HM13] Norbert Heuer and Salim Meddahi. Discontinuous Galerkin hp-bem with quasi-uniform meshes. Numerische Mathematik, 125(4):679–703, December 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-013-0547-3.

Hiltebrand:2014:ESS

[HM14] Andreas Hiltebrand and Siddhartha Mishra. Entropy stable shock capturing space–time discontinuous Galerkin schemes for systems of conservation laws. Numerische Mathematik, 126(1): 103–151, January 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-013-0558-0.

Hu:2016:SBC

[HM16] Jun Hu and Rui Ma. Superconvergence of both the Crouzeix– Raviart and Morley elements. Numerische Mathematik, 132(3): 491–509, March 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-015-0729-2.

Hyvonen:2018:GLT

[HM18] Nuutti Hyvönen and Lauri Mustonen. Generalized linearization techniques in electrical impedance tomography. Numerische Mathematik, 140(1):95–120, September 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
018-0959-1.

Hinrichs:2016:OQM

[HMOU16] Aicke Hinrichs, Lev Markhasin, Jens Oettershagen, and Tino Ullrich. Optimal quasi-Monte Carlo rules on order 2 digital nets for the numerical integration of multivariate periodic functions. Numerische Mathematik, 134(1):163–196, September 2016. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s00211-015-0765-y;   http:
//link.springer.com/article/10.1007/s00211-015-0765-y.

Hu:2017:SFE

[HMX17] Kaibo Hu, Yicong Ma, and Jinchao Xu. Stable finite element methods preserving ∇ · B = 0 exactly for MHD models. Numerische Mathematik, 135(2):371–396, February 2017. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s00211-016-0803-4;   http:
//link.springer.com/article/10.1007/s00211-016-0803-4.

Huhtanen:2016:PLI

[HN16] Marko Huhtanen and Olavi Nevanlinna. Polynomials and lemniscates of indefiniteness. Numerische Mathematik, 133(2):233–253, June 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-015-0745-2.

Heister:2017:ULT

[HOR17a] Timo Heister, Maxim A. Olshanskii, and Leo G. Rebholz. Unconditional long-time stability of a velocity-vorticity method for the 2D Navier–Stokes equations. Numerische Mathematik, 135(1):143–167, January 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/link.springer.com/accesspage/article/10.1007/s00211-
016-0794-1;   http://link.springer.com/article/10.1007/
s00211-016-0794-1.

Horsley:2017:BPF

[Hor17b] David E. Horsley. Bessel phase functions: calculation and application. Numerische Mathematik, 136(3):679–702, July 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hochbruck:2017:EAI

[HP17] Marlis Hochbruck and Tomislav Pazur. Error analysis of implicit Euler methods for quasilinear hyperbolic evolution equations. Numerische Mathematik, 135(2):547–569, February 2017. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s00211-016-0810-5;   http:
//link.springer.com/article/10.1007/s00211-016-0810-5.

Harbrecht:2016:ADM

[HPS16] H. Harbrecht, M. Peters, and M. Siebenmorgen. Analysis of the domain mapping method for elliptic diffusion problems on random domains. Numerische Mathematik, 134(4):823–856, December 2016. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s00211-016-0791-4;   http:
//link.springer.com/article/10.1007/s00211-016-0791-4.

Hochbruck:2018:EAI

[HPS18] Marlis Hochbruck, Tomislav Pazur, and Roland Schnaubelt. Error analysis of implicit Runge–Kutta methods for quasilinear hyperbolic evolution equations. Numerische Mathematik, 138(3): 557–579, March 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-017-0914-6.

Hao:2012:CRT

[HQ12] Dinh Nho Hào and Tran Nhan Tam Quyen. Convergence rates for Tikhonov regularization of a two-coefficient identification problem in an elliptic boundary value problem. Numerische Mathematik, 120(1):45–77, January 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=120&issue=1&spage=45.

Hild:2010:SLM

[HR10] Patrick Hild and Yves Renard. A stabilized Lagrange multiplier method for the finite element approximation of contact problems in elastostatics. Numerische Mathematik, 115 (1):101–129, March 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=115&issue=1&spage=101.

Haine:2012:RID

[HR12] Ghislain Haine and Karim Ramdani. Reconstructing initial data using observers: error analysis of the semi-discrete and fully discrete approximations. Numerische Mathematik, 120(2):307–343, February 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=120&issue=2&spage=307.

Haggblad:2014:ASA

[HR14] Jon Häggblad and Olof Runborg. Accuracy of staircase approximations in finite-difference methods for wave propagation. Numerische Mathematik, 128(4):741–771, December 2014. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-014-0625-1.

Hausenblas:2019:TDS

[HR19] Erika Hausenblas and Tsiry A. Randrianasolo. Time-discretization of stochastic 2-D Navier–Stokes equations with a penaltyprojection method. Numerische Mathematik, 143(2):339–378, October 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL https://link.springer.com/article/
10.1007/s00211-019-01057-3.

Holtz:2012:MTF

[HRS12] Sebastian Holtz, Thorsten Rohwedder, and Reinhold Schneider. On manifolds of tensors of fixed TT-rank. Numerische Mathematik, 120(4):701–731, April 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=120&issue=4&spage=701.

Halvorsen:2013:SGT

[HS13] Tore Gunnar Halvorsen and Torquil Macdonald Sørensen. Simplicial gauge theory on spacetime. Numerische Mathematik, 125(4): 733–760, December 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-013-0552-6.

Hasegawa:2015:EEC

[HS15] Takemitsu Hasegawa and Hiroshi Sugiura. Error estimate for a corrected Clenshaw–Curtis quadrature rule. Numerische Mathematik, 130(1):135–149, May 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-014-0660-y.

Hofmanova:2017:ETI

[HS17] Martina Hofmanová and Katharina Schratz. An exponential-type integrator for the KdV equation. Numerische Mathematik, 136 (4):1117–1137, August 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hsia:2018:LTS

[HS18] Chun-Hsiung Hsia and Ming-Cheng Shiue. On the long-time stability of a temporal discretization scheme for the three dimensional viscous primitive equations. Numerische Mathematik, 139 (1):187–245, May 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-017-0934-2.

Han:2019:CAP

[HS19a] Weimin Han and Mircea Sofonea. Convergence analysis of penalty based numerical methods for constrained inequality problems. Numerische Mathematik, 142(4):917–940, August 2019. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
019-01036-8.

Herrmann:2019:QIL

[HS19b] Lukas Herrmann and Christoph Schwab. QMC integration for lognormal-parametric, elliptic PDEs: local supports and product weights. Numerische Mathematik, 141(1):63–102, January 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00211-018-0991-1.

Han:2018:NAS

[HSD18] Weimin Han, Mircea Sofonea, and David Danan. Numerical analysis of stationary variational–hemivariational inequalities. Numerische Mathematik, 139(3):563–592, July 2018. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
018-0951-9.

Hernandez-Santamaria:2019:GOC

[HSLZ19] Víctor Hernández-Santamaría, Martin Lazar, and Enrique Zuazua. Greedy optimal control for elliptic problems and its application to turnpike problems. Numerische Mathematik, 141(2): 455–493, February 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-018-1005-z.

Hannukainen:2012:UFP

[HSV12] Antti Hannukainen, Rolf Stenberg, and Martin Vohralík. A unified framework for a posteriori error estimation for the Stokes problem. Numerische Mathematik, 122(4):725–769, December 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-012-0472-x.

Hesse:2017:RBF

[HSW17] Kerstin Hesse, Ian H. Sloan, and Robert S. Womersley. Radial basis function approximation of noisy scattered data on the sphere.

Numerische Mathematik, 137(3):579–605, November 2017. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
017-0886-6.

Hu:2012:COA

[HSX12] Jun Hu, Zhongci Shi, and Jinchao Xu. Convergence and optimality of the adaptive Morley element method. Numerische Mathematik, 121(4):731–752, August 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=121&issue=4&spage=731.

Hu:2013:DWH

[HSZ13] Qiya Hu, Shi Shu, and Jun Zou. A discrete weighted Helmholtz decomposition and its application. Numerische Mathematik, 125(1):153–189, September 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-013-0536-6.

Harbrecht:2018:FSG

[HT18] Helmut Harbrecht and Johannes Tausch. A fast sparse grid based space-time boundary element method for the nonstationary heat equation. Numerische Mathematik, 140(1):239–264, September 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL https://link.springer.com/article/10.
1007/s00211-018-0963-5.

Hackbusch:2017:IBH

[HU17] Wolfgang Hackbusch and André Uschmajew. On the interconnection between the higher-order singular values of real tensors. Numerische Mathematik, 135(3):875–894, March 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.com/article/10.1007/s00211-
016-0819-9;   http://link.springer.com/content/pdf/10.
1007/s00211-016-0819-9.pdf.

Huang:2012:USD

[HV12] Chengming Huang and Stefan Vandewalle. Unconditionally stable difference methods for delay partial differential equations. Numerische Mathematik, 122(3):579–601, November 2012. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?

genre=article&issn=0029-599X&volume=122&issue=3&spage=
579.

Howell:2011:ISC

[HW11] Jason S. Howell and Noel J. Walkington. Inf-sup conditions for twofold saddle point problems. Numerische Mathematik, 118(4):663–693, August 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=118&issue=4&spage=663.

Hohage:2013:IRN

[HW13] Thorsten Hohage and Frank Werner. Iteratively regularized Newton-type methods for general data misfit functionals and applications to Poisson data. Numerische Mathematik, 123(4):745–779, April 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/link.springer.com/article/10.1007/s00211-012-0499-z;
 http://link.springer.com/content/pdf/10.1007/s00211-
012-0499-z.pdf.

Holm:2018:CDG

[HW18] Bärbel Holm and Thomas P. Wihler. Continuous and discontinuous Galerkin time stepping methods for nonlinear initial value problems with application to finite time blow-up. Numerische Mathematik, 138(3):767–799, March 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-017-0918-2.

He:2015:NEC

[HY15] Bingsheng He and Xiaoming Yuan. On non-ergodic convergence rate of Douglas–Rachford alternating direction method of multipliers. Numerische Mathematik, 130(3):567–577, July 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-014-0673-6.

He:2018:MNE

[HZZ18] Wenming He, Zhimin Zhang, and Qingsong Zou. Maximumnorms error estimates for high-order finite volume schemes over quadrilateral meshes. Numerische Mathematik, 138(2):473–500, February 2018. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL https://link.springer.com/
article/10.1007/s00211-017-0912-8.

Imbert-Gerard:2015:IPG

[IG15] Lise-Marie Imbert-Gérard. Interpolation properties of generalized plane waves. Numerische Mathematik, 131(4):683–711, December 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-015-0704-y.

Ito:2018:SPR

[IN18] Shinji Ito and Yuji Nakatsukasa. Stable polefinding and rational least-squares fitting via eigenvalues. Numerische Mathematik, 139(3):633–682, July 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https:/

/link.springer.com/article/10.1007/s00211-018-0948-4;
 https://link.springer.com/content/pdf/10.1007/s00211-
018-0948-4.pdf.

Iserles:2011:FSA

[Ise11] Arieh Iserles. A fast and simple algorithm for the computation of Legendre coefficients. Numerische Mathematik, 117(3):529–553, March 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=117&issue=3&spage=529.

Tanaka:2010:FCD

[iTSMM10] Ken ichiro Tanaka, Masaaki Sugihara, Kazuo Murota, and Masatake Mori. Function classes for double exponential integration formulas. Numerische Mathematik, 114(4):631– 655, February 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=114&issue=4&spage=631.

Jin:2010:RLM

[Jin10] Qinian Jin. On a regularized Levenberg–Marquardt method for solving nonlinear inverse problems. Numerische Mathematik, 115(2):229–259, April 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=115&issue=2&spage=229.

Jin:2012:OOR

[Jin12] Qinian Jin. On the order optimality of the regularization via inexact Newton iterations. Numerische Mathematik, 121 (2):237–260, June 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=121&issue=2&spage=237.

Jin:2017:HSR

[Jin17] Qinian Jin. On a heuristic stopping rule for the regularization of inverse problems by the augmented Lagrangian method. Numerische Mathematik, 136(4):973–992, August 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Jooste:2013:ZMP

[JJT13] A. Jooste, K. Jordaan, and F. Toókos. On the zeros of Meixner polynomials. Numerische Mathematik, 124(1):57–71, May 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-012-0504-6.

Jia:2013:CNT

[JL13a] Zhongxiao Jia and Bingyu Li. On the condition number of the total least squares problem. Numerische Mathematik, 125(1): 61–87, September 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-013-0533-9.

Johansson:2013:HOD

[JL13b] August Johansson and Mats G. Larson. A high order discontinuous Galerkin Nitsche method for elliptic problems with fictitious boundary. Numerische Mathematik, 123(4):607–628, April 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-012-0497-1.

Jia:2015:PPI

[JLL15] Zhongxiao Jia, Wen-Wei Lin, and Ching-Sung Liu. A positivity preserving inexact Noda iteration for computing the smallest eigenpair of a large irreducible M-matrix. Numerische Mathematik, 130(4):645–679, August 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-014-0677-2.

Jin:2018:DMR

[JLZ18] Bangti Jin, Buyang Li, and Zhi Zhou. Discrete maximal regularity of time-stepping schemes for fractional evolution equations. Numerische Mathematik, 138(1):101–131, January 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL

https://link.springer.com/article/10.1007/s00211-017-
0904-8;   https://link.springer.com/content/pdf/10.1007/
s00211-017-0904-8.pdf.

Jahnke:2018:AMR

[JM18] Tobias Jahnke and Marcel Mikl. Adiabatic midpoint rule for the dispersion-managed nonlinear Schrödinger equation. Numerische Mathematik, 138(4):975–1009, April 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-017-0926-2.

Jarlebring:2012:LEA

[JMM12] Elias Jarlebring, Wim Michiels, and Karl Meerbergen. A linear eigenvalue algorithm for the nonlinear eigenvalue problem. Numerische Mathematik, 122(1):169–195, September 2012. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=122&issue=1&spage=
169.

Janovska:2010:CCZ

[JO10]  Drahoslava Janovská and Gerhard Opfer.  The classifi-

cation and the computation of the zeros of quaternionic, two-sided polynomials. Numerische Mathematik, 115(1): 81–100, March 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=115&issue=1&spage=81.

Jeon:2013:NLC

[JP13] Youngmok Jeon and Eun-Jae Park. New locally conservative finite element methods on a rectangular mesh. Numerische Mathematik, 123(1):97–119, January 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-012-0477-5.

John:2018:ECF

[JSW18] Lorenz John, Piotr Swierczynski, and Barbara Wohlmuth. Energy corrected FEM for optimal Dirichlet boundary control problems. Numerische Mathematik, 139(4):913–938, August 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00211-018-0952-8.

Jin:2011:INR

[JT11] Qinian Jin and Ulrich Tautenhahn. Inexact Newton regularization methods in Hilbert scales. Numerische Mathematik, 117(3):555–579, March 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=117&issue=3&spage=555.

Javed:2016:EMG

[JT16] Mohsin Javed and Lloyd N. Trefethen. Euler–Maclaurin and Gregory interpolants. Numerische Mathematik, 132(1):201–216, January 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-015-0713-x.

Juntunen:2015:CBS

[Jun15] Mika Juntunen. On the connection between the stabilized Lagrange multiplier and Nitsche’s methods. Numerische Mathematik, 131(3):453–471, November 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.com/article/10.1007/s00211-015-0701-1.

Jin:2016:LMM

[JY16] Qinian Jin and Hongqi Yang. Levenberg–Marquardt method in Banach spaces with general convex regularization terms. Numerische Mathematik, 133(4):655–684, August 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s00211-015-0764-z;   http://link.springer.com/article/
10.1007/s00211-015-0764-z.

Jin:2013:IRG

[JZ13] Qinian Jin and Min Zhong. On the iteratively regularized Gauss– Newton method in Banach spaces with applications to parameter identification problems. Numerische Mathematik, 124(4):

647–683, August 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-013-0529-5.

Jin:2014:NIT

[JZ14] Qinian Jin and Min Zhong. Nonstationary iterated Tikhonov regularization in Banach spaces with uniformly convex penalty terms. Numerische Mathematik, 127(3):485–513, July 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-013-0594-9.

Kaltenbacher:2015:IRG

[Kal15] Barbara Kaltenbacher. An iteratively regularized Gauss–Newton-Halley method for solving nonlinear ill-posed problems. Numerische Mathematik, 131(1):33–57, September 2015. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-014-0682-5.

Karper:2013:CFD

[Kar13] Trygve K. Karper. A convergent FEM–DG method for the compressible Navier–Stokes equations. Numerische Mathematik, 125(3):441–510, November 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-013-0543-7.

Kaltenbacher:2018:CAD

[KdS18] Barbara Kaltenbacher and Mario Luiz Previatti de Souza. Convergence and adaptive discretization of the IRGNM Tikhonov and the IRGNM Ivanov method under a tangential cone condition in Banach space. Numerische Mathematik, 140 (2):449–478, October 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https:/

/link.springer.com/article/10.1007/s00211-018-0971-5;
 https://link.springer.com/content/pdf/10.1007/s00211-
018-0971-5.pdf.

Kim:2019:NMC

[Kim19] Yunho Kim. A Newton’s method characterization for real eigenvalue problems. Numerische Mathematik, 142(4):941–971, August 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL https://link.springer.com/article/
10.1007/s00211-019-01037-7.

Kirby:2011:FSF

[Kir11] Robert C. Kirby. Fast simplicial finite element algorithms using Bernstein polynomials. Numerische Mathematik, 117 (4):631–652, April 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=117&issue=4&spage=631.

Kirby:2017:FIS

[Kir17] Robert C. Kirby. Fast inversion of the simplicial Bernstein mass matrix. Numerische Mathematik, 135(1):73–95, January 2017. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s00211-016-0795-0;   http:
//link.springer.com/article/10.1007/s00211-016-0795-0.

Kammerer:2011:SHC

[KK11] Lutz Kämmerer and Stefan Kunis. On the stability of the hyperbolic cross discrete Fourier transform. Numerische Mathematik, 117(3):581–600, March 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=117&issue=3&spage=581.

Katsaounis:2015:PEC

[KK15a] Theodoros Katsaounis and Irene Kyza. A posteriori error control and adaptivity for Crank–Nicolson finite element approximations for the linear Schrödinger equation. Numerische Mathematik, 129 (1):55–90, January 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-014-0634-0.

Kim:2015:NCC

[KK15b] J. H. Kim and Do Y. Kwak. New curl conforming finite elements on parallelepiped. Numerische Mathematik, 131(3):473– 488, November 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-015-0696-7.

Kirby:2015:SMF

[KK15c] Robert C. Kirby and Thinh Tri Kieu. Symplectic-mixed finite element approximation of linear acoustic wave equations.

Numerische Mathematik, 130(2):257–291, June 2015. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-014-0667-4.

Kinoshita:2014:PEI

[KKN14] Takehiko Kinoshita, Takuma Kimura, and Mitsuhiro T. Nakao. On the a posteriori estimates for inverse operators of linear parabolic equations with applications to the numerical enclosure of solutions for nonlinear problems. Numerische Mathematik, 126(4):679–701, April 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/link.springer.com/article/10.1007/s00211-013-0575-z;
 http://link.springer.com/content/pdf/10.1007/s00211-
013-0575-z.pdf.

Karlsen:2012:EEF

[KKR12] K. H. Karlsen, U. Koley, and N. H. Risebro. An error estimate for the finite difference approximation to degenerate convection-diffusion equations. Numerische Mathematik, 121 (2):367–395, June 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=121&issue=2&spage=367.

Kovacs:2017:SCF

[KL17] Balázs Kovács and Christian Lubich. Stable and convergent fully discrete interior–exterior coupling of Maxwell’s equations. Numerische Mathematik, 137(1):91–117, September 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See correction [NKL21].

Kovacs:2018:LIF

[KL18a] Balázs Kovács and Christian Lubich. Linearly implicit full discretization of surface evolution. Numerische Mathematik, 140(1): 121–152, September 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-018-0962-6.

Kovacs:2018:SCT

[KL18b] Balázs Kovács and Christian Lubich. Stability and convergence of time discretizations of quasi-linear evolution equations of Kato type. Numerische Mathematik, 138(2):365–388, February 2018.

CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00211-017-0909-3.

Klapproth:2015:CSN

[Kla15] Corinna Klapproth. The contact-stabilized Newmark method: consistency error of a spatiotemporal discretization. Numerische Mathematik, 131(1):59–82, September 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.com/article/10.1007/s00211-014-0686-1.

Kovacs:2019:CEF

[KLL19] Balázs Kovács, Buyang Li, and Christian Lubich. A convergent evolving finite element algorithm for mean curvature flow of closed surfaces. Numerische Mathematik, 143(4):797–853, December 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL https://link.springer.com/article/
10.1007/s00211-019-01074-2.

Kovacs:2017:CFE

[KLLG17] Balázs Kovács, Buyang Li, Christian Lubich, and Christian A. Power Guerra. Convergence of finite elements on an evolving surface driven by diffusion on the surface. Numerische Mathematik, 137(3):643–689, November 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-017-0888-4.

Kuo:2019:SPF

[KLS19] Yueh-Cheng Kuo, Wen-Wei Lin, and Shih-Feng Shieh. A structure preserving flow for computing Hamiltonian matrix exponential. Numerische Mathematik, 143(3):555–582, November 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00211-019-01065-3.

Kunemund:2019:HOM

[KNWW19] Jens Künemund, Francis J. Narcowich, Joseph D. Ward, and Holger Wendland. A high-order meshless Galerkin method for semilinear parabolic equations on spheres. Numerische Mathematik, 142(2):383–419, June 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-018-01021-7.

Koev:2019:AEE

[Koe19] Plamen Koev. Accurate eigenvalues and exact zero Jordan blocks of totally nonnegative matrices. Numerische Mathematik, 141 (3):693–713, March 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-019-01022-0.

Kopteva:2017:ENP

[Kop17] Natalia Kopteva. Energy-norm a posteriori error estimates for singularly perturbed reaction-diffusion problems on anisotropic meshes. Numerische Mathematik, 137(3):607–642, November 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL https://link.springer.com/article/10.
1007/s00211-017-0889-3.

Kashiwabara:2016:PMP

[KOZ16] Takahito Kashiwabara, Issei Oikawa, and Guanyu Zhou. Penalty method with P1/P1 finite element approximation for the Stokes equations under the slip boundary condition. Numerische Mathematik, 134(4):705–740, December 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/link.springer.com/accesspage/article/10.1007/s00211-
016-0790-5;   http://link.springer.com/article/10.1007/
s00211-016-0790-5.

Karpinski:2017:AIP

[KP17] Stefan Karpinski and Iuliu Sorin Pop. Analysis of an interior penalty discontinuous Galerkin scheme for two phase flow in porous media with dynamic capillary effects. Numerische Mathematik, 136(1):249–286, May 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kumar:2014:CAC

[KPR14] K. Kumar, I. S. Pop, and F. A. Radu. Convergence analysis for a conformal discretization of a model for precipitation and dissolution in porous media. Numerische Mathematik, 127(4): 715–749, August 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-013-0601-1.

Kristensen:2016:PAA

[KR16] Jan Kristensen and Filip Rindler. Piecewise affine approximations for functions of bounded variation. Numerische Mathematik, 132(2):329–346, February 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-015-0721-x.

Kress:2017:TIE

[Kre17] Rainer Kress. On Trefftz’ integral equation for the Bernoulli free boundary value problem. Numerische Mathematik, 136(2):503– 522, June 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Knizhnerman:2011:CAE

[KS11a] L. Knizhnerman and V. Simoncini. Convergence analysis of the extended Krylov subspace method for the Lyapunov equation. Numerische Mathematik, 118(3):567–586, July 2011. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=118&issue=3&spage=
567.

Kopteva:2011:SAI

[KS11b] Natalia Kopteva and Martin Stynes. Stabilised approximation of interior-layer solutions of a singularly perturbed semilinear reaction-diffusion problem. Numerische Mathematik, 119(4):787–810, December 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=119&issue=4&spage=787.

Kreuzer:2011:DRA

[KS11c] Christian Kreuzer and Kunibert G. Siebert. Decay rates of adaptive finite elements with Dörfler marking. Numerische Mathematik, 117(4):679–716, April 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=117&issue=4&spage=679.

Kazeev:2018:QTS

[KS18a] Vladimir Kazeev and Christoph Schwab. Quantized tensorstructured finite elements for second-order elliptic PDEs in two dimensions. Numerische Mathematik, 138(1):133–190, January 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL https://link.springer.com/article/10.
1007/s00211-017-0899-1.

Kemmochi:2018:DMR

[KS18b] Tomoya Kemmochi and Norikazu Saito. Discrete maximal regularity and the finite element method for parabolic equations. Numerische Mathematik, 138(4):905–937, April 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
017-0929-z.

Klapproth:2010:CRN

[KSD10] Corinna Klapproth, Anton Schiela, and Peter Deuflhard. Consistency results on Newmark methods for dynamical contact problems. Numerische Mathematik, 116(1):65–94, July 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=116&issue=1&spage=
65.

Klapproth:2011:ATC

[KSD11] Corinna Klapproth, Anton Schiela, and Peter Deuflhard. Adaptive timestep control for the contact-stabilized Newmark method. Numerische Mathematik, 119(1):49–81, September 2011. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=119&issue=1&spage=
49.

Krendl:2013:SES

[KSZ13] Wolfgang Krendl, Valeria Simoncini, and Walter Zulehner. Stability estimates and structural spectral properties of saddle point problems. Numerische Mathematik, 124(1):183–213, May 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-012-0507-3.

Kirby:2012:FSQ

[KT12] Robert C. Kirby and Kieu Tri Thinh. Fast simplicial quadraturebased finite element operators using Bernstein polynomials. Numerische Mathematik, 121(2):261–279, June 2012. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=121&issue=2&spage=

Krause:2015:ERR

[KVW15] Rolf Krause, Andreas Veeser, and Mirjam Walloth. An efficient and reliable residual-type a posteriori error estimator for the Signorini problem. Numerische Mathematik, 130(1):151–197, May 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-014-0655-8.

Kreusler:2012:MRE

[KY12] Hans-Christian Kreusler and Harry Yserentant. The mixed regularity of electronic wave functions in fractional order and weighted Sobolev spaces. Numerische Mathematik, 121(4): 781–802, August 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=121&issue=4&spage=781.

Li:2011:SDS

[LA11] Yuan Li and Rong An. Semi-discrete stabilized finite element methods for Navier–Stokes equations with nonlinear slip boundary conditions based on regularization procedure. Numerische Mathematik, 117(1):1–36, January 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=117&issue=1&spage=1.

Li:2014:OTL

[LC14] Jian Li and Zhangxin Chen. Optimal L^{2}, H^{1} and L^{∞} analysis of finite volume methods for the stationary Navier–Stokes equations with large data. Numerische Mathematik, 126(1):75–101, January 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-013-0556-2.

Li:2012:SML

[LCH12] Jian Li, Zhangxin Chen, and Yinnian He. A stabilized multilevel method for non-singular finite volume solutions of the stationary 3D Navier–Stokes equations. Numerische Mathematik, 122(2):279–304, October 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=122&issue=2&spage=279.

Liu:2017:EAM

[LCWW17] Yuan Liu, Wenbin Chen, Cheng Wang, and Steven M. Wise. Error analysis of a mixed finite element method for a Cahn–Hilliard– Hele–Shaw system. Numerische Mathematik, 135(3):679–709, March 2017. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s00211-016-0813-2;   http:
//link.springer.com/article/10.1007/s00211-016-0813-2.

Lesinigo:2011:MDB

[LDQ11] Matteo Lesinigo, Carlo D’Angelo, and Alfio Quarteroni. A multiscale Darcy–Brinkman model for fluid flow in fractured porous media. Numerische Mathematik, 117(4):717–752, April 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=117&issue=4&spage=
717.

Lopez-Fernandez:2016:GCQ

[LFS16] M. Lopez-Fernandez and S. Sauter. Generalized convolution quadrature based on Runge–Kutta methods. Numerische Mathematik, 133(4):743–779, August 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/link.springer.com/accesspage/article/10.1007/s00211-
015-0761-2;   http://link.springer.com/article/10.1007/
s00211-015-0761-2.

Li:2014:FSL

[LGC14] Shengguo Li, Ming Gu, and Lizhi Cheng. Fast structured LU factorization for nonsymmetric matrices. Numerische Mathematik, 127(1):35–55, May 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-013-0582-0.

Liu:2017:NNI

[LGL17] Ching-Sung Liu, Chun-Hua Guo, and Wen-Wei Lin. Newton– Noda iteration for finding the Perron pair of a weakly irreducible nonnegative tensor. Numerische Mathematik, 137(1): 63–90, September 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lu:2014:MPL

[LHQ14] Changna Lu, Weizhang Huang, and Jianxian Qiu. Maximum principle in linear finite element approximations of anisotropic diffusion-convection-reaction problems. Numerische Mathematik, 127(3):515–537, July 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-013-0595-8.

Li:2015:MNS

[Li15] Buyang Li. Maximum-norm stability and maximal L^{p} regularity of FEMs for parabolic equations with Lipschitz continuous coefficients. Numerische Mathematik, 131(3):489–516, November 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-015-0698-5.

Lin:2017:RES

[Lin17] Lin Lin. Randomized estimation of spectral densities of large matrices made accurate. Numerische Mathematik, 136(1):183– 213, May 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Luthen:2018:IPE

[LJS18] Nora Lüthen, Mika Juntunen, and Rolf Stenberg. An improved a priori error analysis of Nitsche’s method for Robin boundary conditions. Numerische Mathematik, 138(4):1011–1026, April 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00211-017-0927-1.

Lu:2010:EEN

[LL10] Xiliang Lu and Ping Lin. Error estimate of the P_{1} nonconforming finite element method for the penalized unsteady Navier–Stokes equations. Numerische Mathematik, 115(2): 261–287, April 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=115&issue=2&spage=261.

Larsson:2012:CPL

[LL12] Karl Larsson and Mats G. Larson. Continuous piecewise linear finite elements for the Kirchhoff–Love plate equation. Numerische Mathematik, 121(1):65–97, May 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=121&issue=1&spage=65.

Li:2010:SBE

[LLW10] Ren-Cang Li, Wen-Wei Lin, and Chern-Shuh Wang. Structured backward error for palindromic polynomial eigenvalue problems. Numerische Mathematik, 116(1):95–122, July 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=116&issue=1&spage=
95.

Liang:2015:FEA

[LLY15] Kewei Liang, Xiliang Lu, and Jerry Zhijian Yang. Finite element approximation to the extremal eigenvalue problem for inhomogeneous materials. Numerische Mathematik, 130(4):741–762, August 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-014-0678-1.

Larson:2011:PEE

[LM11] Mats G. Larson and Axel Målqvist. A posteriori error estimates for mixed finite element approximations of parabolic problems. Numerische Mathematik, 118(1):33–48, May 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=118&issue=1&spage=
33.

Lubich:2013:IPE

[LM13] Christian Lubich and Charalambos Makridakis. Interior a posteriori error estimates for time discrete approximations of parabolic problems. Numerische Mathematik, 124(3):541–557, July 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-013-0520-1.

Louka:2015:CES

[LM15] M. A. Louka and N. M. Missirlis. A comparison of the Extrapolated Successive Overrelaxation and the Preconditioned Simultaneous Displacement methods for augmented linear systems. Numerische Mathematik, 131(3):517–540, November 2015.

CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-015-0697-6.

Li:2017:TRN

[LMcS17] Hongliang Li, Pingbing Ming, and Zhong ci Shi. Two robust nonconforming H^{2}-elements for linear strain gradient elasticity. Numerische Mathematik, 137(3):691–711, November 2017. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
017-0890-x.

Lepe:2019:MDG

[LMMR19] Felipe Lepe, Salim Meddahi, David Mora, and Rodolfo Rodríguez. Mixed discontinuous Galerkin approximation of the elasticity eigenproblem. Numerische Mathematik, 142(3):749– 786, July 2019. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL https://link.springer.com/
article/10.1007/s00211-019-01035-9.

Loneland:2016:AAS

[LMR16] A. Loneland, L. Marcinkowski, and T. Rahman. Additive average Schwarz method for a Crouzeix–Raviart finite volume element discretization of elliptic problems with heterogeneous coefficients. Numerische Mathematik, 134(1):91–118, September 2016. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s00211-015-0771-0;   http:
//link.springer.com/article/10.1007/s00211-015-0771-0.

Lederer:2019:RPE

[LMS19] Philip Lukas Lederer, Christian Merdon, and Joachim Schöberl. Refined a posteriori error estimation for classical and pressurerobust Stokes finite element methods. Numerische Mathematik, 142(3):713–748, July 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-019-01049-3.

Lanza:2017:NNO

[LMSS17] A. Lanza, S. Morigi, I. Selesnick, and F. Sgallari. Nonconvex nonsmooth optimization via convex-nonconvex majorizationminimization. Numerische Mathematik, 136(2):343–381, June 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Li:2018:IEE

[LMY18] Hengguang Li, Lin Mu, and Xiu Ye. Interior energy error estimates for the weak Galerkin finite element method. Numerische Mathematik, 139(2):447–478, June 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-017-0940-4.

Li:2015:SBS

[LN15] Wen Li and Michael K. Ng. Some bounds for the spectral radius of nonnegative tensors. Numerische Mathematik, 130(2):315–335, June 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-014-0666-5.

Li:2018:RPE

[LN18] Hengguang Li and Serge Nicaise. Regularity and a priori error analysis on anisotropic meshes of a Dirichlet problem in polyhedral domains. Numerische Mathematik, 139(1):47–92, May 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00211-017-0936-0.

Li:2014:PEE

[LO14] Hengguang Li and Jeffrey S. Ovall. A posteriori error estimation of hierarchical type for the Schrödinger operator with inverse square potential. Numerische Mathematik, 128(4):707–740, December 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-014-0628-y.

Li:2016:ABA

[LOSV16] Xingjie Helen Li, Christoph Ortner, Alexander V. Shapeev, and Brian Van Koten. Analysis of blended atomistic/continuum hybrid methods. Numerische Mathematik, 134(2):275–326, October 2016. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s00211-015-0772-z;   http:
//link.springer.com/article/10.1007/s00211-015-0772-z.

Lozinski:2019:PDG

[Loz19] Alexei Lozinski. A primal discontinuous Galerkin method with static condensation on very general meshes. Numerische Mathematik, 143(3):583–604, November 2019. CODEN NUMMA7.

ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-019-01067-1.

Lu:2011:MPR

[LP11] Shuai Lu and Sergei V. Pereverzev. Multi-parameter regularization and its numerical realization. Numerische Mathematik, 118(1):1–31, May 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=118&issue=1&spage=1.

Liu:2015:LDG

[LP15] Hailiang Liu and Nattapol Ploymaklam. A local discontinuous Galerkin method for the Burgers–Poisson equation. Numerische Mathematik, 129(2):321–351, February 2015. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-014-0641-1.

Lehrenfeld:2017:OPN

[LR17] Christoph Lehrenfeld and Arnold Reusken. Optimal preconditioners for Nitsche–XFEM discretizations of interface problems. Numerische Mathematik, 135(2):313–332, February 2017. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s00211-016-0801-6;   http:
//link.springer.com/article/10.1007/s00211-016-0801-6.

Lelievre:2019:HMC

[LRS19] Tony Lelièvre, Mathias Rousset, and Gabriel Stoltz. Hybrid Monte Carlo methods for sampling probability measures on submanifolds. Numerische Mathematik, 143(2):379–421, October 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL https://link.springer.com/article/10.
1007/s00211-019-01056-4. See correction [LRS20].

Lelievre:2020:CHM

[LRS20] Tony Lelièvre, Mathias Rousset, and Gabriel Stoltz. Correction to: Hybrid Monte Carlo methods for sampling probability measures on submanifolds. Numerische Mathematik, 144(2):447–449, February 2020. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-019-01086-y;

 https://link.springer.com/content/pdf/10.1007/s00211-
019-01086-y.pdf. See [LRS19].

Li:2016:SAY

[LS16] Jichun Li and Sidney Shields. Superconvergence analysis of Yee scheme for metamaterial Maxwell’s equations on non-uniform rectangular meshes. Numerische Mathematik, 134(4):741–781, December 2016. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s00211-015-0788-4;   http:
//link.springer.com/article/10.1007/s00211-015-0788-4.

Lasser:2017:DHK

[LS17] Caroline Lasser and David Sattlegger. Discretising the Herman– Kluk propagator. Numerische Mathematik, 137(1):119–157, September 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Leobacher:2018:CEM

[LS18] Gunther Leobacher and Michaela Szölgyenyi. Convergence of the Euler–Maruyama method for multidimensional SDEs with discontinuous drift and degenerate diffusion coefficient. Numerische Mathematik, 138(1):219–239, January 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https:/

/link.springer.com/article/10.1007/s00211-017-0903-9;
 https://link.springer.com/content/pdf/10.1007/s00211-
017-0903-9.pdf.

LeGia:2011:SPH

[LSW11] Q. T. Le Gia, I. H. Sloan, and Andrew J. Wathen. Stability and preconditioning for a hybrid approximation on the sphere. Numerische Mathematik, 118(4):695–711, August 2011. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=118&issue=4&spage=
695.

LeGia:2012:MRC

[LSW12] Q. T. Le Gia, I. H. Sloan, and H. Wendland. Multiscale RBF collocation for solving PDEs on spheres. Numerische Mathematik, 121(1):99–125, May 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=121&issue=1&spage=99.

Lopez:2010:ANP

[LT10] José Luis López and Nico M. Temme. Asymptotics and numerics of polynomials used in Tricomi and Buchholz expansions of Kummer functions. Numerische Mathematik, 116 (2):269–289, August 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=116&issue=2&spage=269.

Lui:2017:LSC

[Lui17] S. H. Lui. Legendre spectral collocation in space and time for PDEs. Numerische Mathematik, 136(1):75–99, May 2017. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lang:2013:MOC

[LV13] J. Lang and J. G. Verwer. W -methods in optimal control. Numerische Mathematik, 124(2):337–360, June 2013. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-013-0516-x.

Leykekhman:2017:DMP

[LV17] Dmitriy Leykekhman and Boris Vexler. Discrete maximal parabolic regularity for Galerkin finite element methods. Numerische Mathematik, 135(3):923–952, March 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s00211-016-0821-2;   http://link.springer.com/article/
10.1007/s00211-016-0821-2.

Lipnikov:2014:DGM

[LVY14] Konstantin Lipnikov, Danail Vassilev, and Ivan Yotov. Discontinuous Galerkin and mimetic finite difference methods for coupled Stokes–Darcy flows on polygonal and polyhedral grids. Numerische Mathematik, 126(2):321–360, February 2014. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-013-0563-3.

Li:2015:CBL

[LZ15] Ren-Cang Li and Lei-Hong Zhang. Convergence of the block Lanczos method for eigenvalue clusters. Numerische Mathematik, 131(1):83–113, September 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-014-0681-6.

Makridakis:2018:BOA

[Mak18] Charalambos G. Makridakis. On the Babuska–Osborn approach to finite element analysis: L^{2} estimates for unstructured meshes. Numerische Mathematik, 139(4):831–844, August 2018. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
018-0955-5.

Mantica:2013:DIC

[Man13] Giorgio Mantica. Direct and inverse computation of Jacobi matrices of infinite iterated function systems. Numerische Mathematik, 125(4):705–731, December 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.com/article/10.1007/s00211-013-0551-7.

Mansour:2015:GRK

[Man15a] Dhia Mansour. Gauss–Runge–Kutta time discretization of wave equations on evolving surfaces. Numerische Mathematik, 129 (1):21–53, January 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-014-0632-2.

Manton:2015:FGN

[Man15b] Jonathan H. Manton. A framework for generalising the Newton method and other iterative methods from Euclidean space to manifolds. Numerische Mathematik, 129(1):91–125, January 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-014-0630-4.

Mascarenhas:2014:SBI

[Mas14] Walter F. Mascarenhas. The stability of barycentric interpolation at the Chebyshev points of the second kind. Numerische Mathematik, 128(2):265–300, October 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.com/article/10.1007/s00211-014-0612-6.

Maset:2016:AFN

[Mas16] Stefano Maset. An abstract framework in the numerical solution of boundary value problems for neutral functional differential equations. Numerische Mathematik, 133(3):525–555, July 2016. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s00211-015-0754-1;   http:
//link.springer.com/article/10.1007/s00211-015-0754-1.

Matt:2013:TME

[Mat13] Michael A. Matt. A C^{r} trivariate macro-element based on the Worsey–Farin split of a tetrahedron. Numerische Mathematik, 123(1):121–144, January 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-012-0478-4.

Mazure:2011:HBA

[Maz11a] Marie-Laurence Mazure. How to build all Chebyshevian spline spaces good for geometric design? Numerische Mathematik, 119(3):517–556, November 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=119&issue=3&spage=517.

Mazure:2011:QEC

[Maz11b] Marie-Laurence Mazure. Quasi Extended Chebyshev spaces and weight functions. Numerische Mathematik, 118(1):79–108, May 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0029-599X&volume=118&issue=1&spage=
79.

Mascarenhas:2017:ERE

[MdC17] Walter F. Mascarenhas and André Pierro de Camargo. The effects of rounding errors in the nodes on barycentric interpolation. Numerische Mathematik, 135(1):113–141, January 2017. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s00211-016-0798-x;   http:
//link.springer.com/article/10.1007/s00211-016-0798-x.

Mengi:2011:LNM

[Men11] Emre Mengi. Locating a nearest matrix with an eigenvalue of prespecified algebraic multiplicity. Numerische Mathematik, 118(1):109–135, May 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=118&issue=1&spage=109.

He:2017:UHO

[mHZZ17] Wen ming He, Zhimin Zhang, and Qingsong Zou. Ultraconvergence of high order FEMs for elliptic problems with variable coefficients. Numerische Mathematik, 136(1):215–248, May 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Mirebeau:2012:OAM

[Mir12] Jean-Marie Mirebeau. Optimally adapted meshes for finite elements of arbitrary order and W^{1,p} norms. Numerische Mathematik, 120(2):271–305, February 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=120&issue=2&spage=271.

Mirebeau:2014:EFM

[Mir14] Jean-Marie Mirebeau. Efficient fast marching with Finsler metrics. Numerische Mathematik, 126(3):515–557, March 2014. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-013-0571-3.

Mirebeau:2016:AAH

[Mir16] Jean-Marie Mirebeau. Adaptive, anisotropic and hierarchical cones of discrete convex functions. Numerische Mathematik, 132 (4):807–853, April 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-015-0732-7.

Massing:2014:SNO

[MLLR14] André Massing, Mats G. Larson, Anders Logg, and Marie E. Rognes. A stabilized Nitsche overlapping mesh method for the Stokes problem. Numerische Mathematik, 128(1):73–101, September 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-013-0603-z.

McLachlan:2016:SME

[MMMKV16] Robert I. McLachlan, Klas Modin, Hans Munthe-Kaas, and Olivier Verdier. B-series methods are exactly the affine equivariant methods. Numerische Mathematik, 133(3):599–622, July 2016. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s00211-015-0753-2;   http:
//link.springer.com/article/10.1007/s00211-015-0753-2.

Manna:2010:FNA

[MMMS10] Dante V. Manna, Luis A. Medina, Victor H. Moll, and Armin Straub. A fast numerical algorithm for the integration of rational functions. Numerische Mathematik, 115 (2):289–307, April 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=115&issue=2&spage=289.

Motamed:2013:SCM

[MNT13] Mohammad Motamed, Fabio Nobile, and Raúl Tempone. A stochastic collocation method for the second order wave equation with a discontinuous random speed. Numerische Mathematik, 123 (3):493–536, March 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-012-0493-5.

Maree:2017:PEE

[MOGO17] S. C. Maree, L. Ortiz-Gracia, and C. W. Oosterlee. Pricing earlyexercise and discrete barrier options by Shannon wavelet expansions. Numerische Mathematik, 136(4):1035–1070, August 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Monzon:2019:AIE

[Mon19] Gabriel Monzón. Anisotropic interpolation error estimate for arbitrary quadrilateral isoparametric elements. Numerische Mathematik, 143(2):315–338, October 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/s00211-019-01061-7.

Massoudi:2017:AMB

[MOR17] Arash Massoudi, Mark R. Opmeer, and Timo Reis. The ADI method for bounded real and positive real Lur’e equations. Numerische Mathematik, 135(2):431–458, February 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s00211-016-0805-2;   http://link.springer.com/article/
10.1007/s00211-016-0805-2.

Makridakis:2011:PEA

[MOS11] Charalambos Makridakis, Christoph Ortner, and Endre Süli. A priori error analysis of two force-based atomistic/continuum models of a periodic chain. Numerische Mathematik, 119 (1):83–121, September 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=119&issue=1&spage=83.

Maalqvist:2015:CEN

[MP15] Axel Målqvist and Daniel Peterseim. Computation of eigenvalues by numerical upscaling. Numerische Mathematik, 130 (2):337–361, June 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-014-0665-6.

Maalqvist:2018:MTP

[MP18a] Axel Målqvist and Anna Persson. Multiscale techniques for parabolic equations. Numerische Mathematik, 138(1):191–217, January 2018. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL https://link.springer.com/
article/10.1007/s00211-017-0905-7;   https://link.springer.
com/content/pdf/10.1007/s00211-017-0905-7.pdf.

Moiola:2018:STT

[MP18b] Andrea Moiola and Ilaria Perugia. A space-time Trefftz discontinuous Galerkin method for the acoustic wave equation in first-order formulation. Numerische Mathematik, 138(2):389– 435, February 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-017-0910-x.

Meng:2017:DGM

[MR17] Xiong Meng and Jennifer K. Ryan. Discontinuous Galerkin methods for nonlinear scalar hyperbolic conservation laws: divided difference estimates and accuracy enhancement. Numerische Mathematik, 136(1):27–73, May 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL

https://link.springer.com/content/pdf/10.1007/s00211-
016-0833-y.pdf.

Merrien:2019:GTO

[MS19] Jean-Louis Merrien and Tomas Sauer. Generalized Taylor operators and polynomial chains for Hermite subdivision schemes. Numerische Mathematik, 142(1):167–203, May 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
018-0996-9.

Magoules:2017:AOS

[MSV17] Frédéric Magoulès, Daniel B. Szyld, and Cédric Venet. Asynchronous optimized Schwarz methods with and without overlap. Numerische Mathematik, 137(1):199–227, September 2017. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Mardal:2013:USF

[MSW13] Kent-Andre Mardal, Joachim Schöberl, and Ragnar Winther. A uniformly stable Fortin operator for the Taylor–Hood element. Numerische Mathematik, 123(3):537–551, March 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-012-0492-6.

Murakawa:2017:LFV

[Mur17] Hideki Murakawa. A linear finite volume method for nonlinear cross-diffusion systems. Numerische Mathematik, 136(1):1–26, May 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Mustapha:2015:TSD

[Mus15] Kassem Mustapha. Time-stepping discontinuous Galerkin methods for fractional diffusion problems. Numerische Mathematik, 130(3):497–516, July 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-014-0669-2.

Mertens:2015:SRC

[MV15] Clara Mertens and Raf Vandebril. Short recurrences for computing extended Krylov bases for Hermitian and unitary matrices. Numerische Mathematik, 131(2):303–328, October 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-014-0692-3.

Nakatsukasa:2012:CNM

[Nak12] Yuji Nakatsukasa. On the condition numbers of a multiple eigenvalue of a generalized eigenvalue problem. Numerische Mathematik, 121(3):531–544, July 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=121&issue=3&spage=531.

Nakano:2017:CMC

[Nak17] Yumiharu Nakano. Convergence of meshfree collocation methods for fully nonlinear parabolic equations. Numerische Mathematik, 136(3):703–723, July 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Noumir:2014:NEM

[NDL14] Youness Noumir, François Dubois, and Olivier Lafitte. Numerical Eulerian method for linearized gas dynamics in the high frequency regime. Numerische Mathematik, 127(4):641–683, August 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-013-0598-5.

Neilan:2010:NMF

[Nei10] Michael Neilan. A nonconforming Morley finite element method for the fully nonlinear Monge–Ampère equation. Numerische Mathematik, 115(3):371–394, May 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=115&issue=3&spage=371.

Niu:2010:MTF

[NGKN10] Qiang Niu, Laura Grigori, Pawan Kumar, and Frédéric Nataf. Modified tangential frequency filtering decomposition and its Fourier analysis. Numerische Mathematik, 116 (1):123–148, July 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=116&issue=1&spage=123.

Nievergelt:2010:MST

[Nie10] Yves Nievergelt. Median spheres: theory, algorithms, applications. Numerische Mathematik, 114(4):573–606, February 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=114&issue=4&spage=
573.

Nick:2021:CSC

[NKL21] Jörg Nick, Balázs Kovács, and Christian Lubich. Correction to: Stable and convergent fully discrete interior–exterior coupling of Maxwell’s equations. Numerische Mathematik, 147(4):997–1000, April 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.com/article/10.1007/s00211-021-01196-6;
 https://www.math.utah.edu/pub/tex/bib/nummath2010.bib;
 https://link.springer.com/content/pdf/10.1007/s00211-
021-01196-6.pdf. See [KL17].

Napov:2011:SFO

[NN11] Artem Napov and Yvan Notay. Smoothing factor, order of prolongation and actual multigrid convergence. Numerische Mathematik, 118(3):457–483, July 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=118&issue=3&spage=457.

Nakatsukasa:2015:CCZ

[NNT15] Yuji Nakatsukasa, Vanni Noferini, and Alex Townsend. Computing the common zeros of two bivariate functions via Bézout resultants. Numerische Mathematik, 129(1):181–209, January 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-014-0635-z.

Nochetto:2016:PPI

[NOS16] Ricardo H. Nochetto, Enrique Otárola, and Abner J. Salgado. Piecewise polynomial interpolation in Muckenhoupt weighted Sobolev spaces and applications. Numerische Mathematik, 132 (1):85–130, January 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-015-0709-6.

Notaris:2016:CFQ

[Not16a] Sotirios E. Notaris. On a corrected Fejér quadrature formula of the second kind. Numerische Mathematik, 133(2):279–302, June 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-015-0750-5.

Notay:2016:NAM

[Not16b] Yvan Notay. A new algebraic multigrid approach for Stokes problems. Numerische Mathematik, 132(1):51–84, January 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-015-0710-0.

Notaris:2019:SPR

[Not19] Sotirios E. Notaris. Stieltjes polynomials and related quadrature formulae for a class of weight functions, II. Numerische Mathematik, 142(1):129–147, May 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/s00211-018-1009-8.

Nouy:2019:HOP

[Nou19] Anthony Nouy. Higher-order principal component analysis for the approximation of tensors in tree-based low-rank formats. Numerische Mathematik, 141(3):743–789, March 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
018-1017-8.

Novati:2014:NAF

[Nov14] P. Novati. Numerical approximation to the fractional derivative operator. Numerische Mathematik, 127(3):539–566, July 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-013-0596-7.

Nguyen:2015:CAF

[NP15] Giang T. Nguyen and Federico Poloni. Componentwise accurate fluid queue computations using doubling algorithms. Numerische Mathematik, 130(4):763–792, August 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.com/article/10.1007/s00211-014-0675-4.

Neitzel:2019:SCA

[NPVW19] Ira Neitzel, Konstantin Pieper, Boris Vexler, and Daniel Walter. A sparse control approach to optimal sensor placement in PDEconstrained parameter estimation problems. Numerische Mathematik, 143(4):943–984, December 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-019-01073-3.

Neuenkirch:2014:FOS

[NS14] Andreas Neuenkirch and Lukasz Szpruch. First order strong approximations of scalar SDEs defined in a domain. Numerische Mathematik, 128(1):103–136, September 2014. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-014-0606-4.

Nobile:2016:CQO

[NTT16] F. Nobile, L. Tamellini, and R. Tempone. Convergence of quasioptimal sparse-grid approximation of Hilbert-space-valued functions: application to random elliptic PDEs. Numerische Mathematik, 134(2):343–388, October 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/link.springer.com/accesspage/article/10.1007/s00211-
015-0773-y;   http://link.springer.com/article/10.1007/
s00211-015-0773-y.

Nobile:2018:SAM

[NTW18] Fabio Nobile, Raúl Tempone, and Sören Wolfers. Sparse approximation of multilinear problems with applications to kernel-based methods in UQ. Numerische Mathematik, 139(1):247–280, May 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL https://link.springer.com/article/10.
1007/s00211-017-0932-4.

Neitzel:2012:PEE

[NV12] Ira Neitzel and Boris Vexler. A priori error estimates for space–time finite element discretization of semilinear parabolic optimal control problems. Numerische Mathematik, 120(2):

345–386, February 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=120&issue=2&spage=345.

Nochetto:2010:PEA

[NvPZ10] Ricardo H. Nochetto, Tobias von Petersdorff, and Chen-Song Zhang. A posteriori error analysis for a class of integral equations and variational inequalities. Numerische Mathematik, 116(3):519–552, September 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=116&issue=3&spage=519.

Neitzel:2018:PLL

[NW18] I. Neitzel and W. Wollner. A priori L^{2}-discretization error estimates for the state in elliptic optimization problems with pointwise inequality state constraints. Numerische Mathematik, 138(2):273–299, February 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/s00211-017-0906-6.

Nochetto:2019:PRC

[NZ19] Ricardo H. Nochetto and Wujun Zhang. Pointwise rates of convergence for the Oliker–Prussner method for the Monge–Ampère equation. Numerische Mathematik, 141(1):253–288, January 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL https://link.springer.com/article/10.
1007/s00211-018-0988-9.

Ovcharova:2017:CRA

[OB17] Nina Ovcharova and Lothar Banz. Coupling regularization and adaptive hp-BEM for the solution of a delamination problem. Numerische Mathematik, 137(2):303–337, October 2017. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
017-0879-5.

Olson:2018:FBA

[OLOV18] Derek Olson, Xingjie Li, Christoph Ortner, and Brian Van Koten. Force-based atomistic/continuum blending for multilattices. Numerische Mathematik, 140(3):703–754, November 2018. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
018-0979-x.

Olver:2010:SGO

[Olv10] Sheehan Olver. Shifted GMRES for oscillatory integrals. Numerische Mathematik, 114(4):607–628, February 2010. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=114&issue=4&spage=
607.

Olver:2012:GFS

[Olv12] Sheehan Olver. A general framework for solving Riemann– Hilbert problems numerically. Numerische Mathematik, 122 (2):305–340, October 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=122&issue=2&spage=305.

Okayama:2013:EEE

[OMS13] Tomoaki Okayama, Takayasu Matsuo, and Masaaki Sugihara. Error estimates with explicit constants for Sinc approximation, Sinc quadrature and Sinc indefinite integration. Numerische Mathematik, 124(2):361–394, June 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-013-0515-y.

Of:2015:ESF

[OPS15] G. Of, T. X. Phan, and O. Steinbach. An energy space finite element approach for elliptic Dirichlet boundary control problems. Numerische Mathematik, 129(4):723–748, April 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-014-0653-x.

Olshanskii:2010:FEM

[OR10] Maxim A. Olshanskii and Arnold Reusken. A finite element method for surface PDEs: matrix properties. Numerische Mathematik, 114(3):491–520, January 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=114&issue=3&spage=491.

Offner:2013:SCO

[ÖS13] Philipp Öffner and Thomas Sonar. Spectral convergence for orthogonal polynomials on triangles. Numerische Mathematik, 124 (4):701–721, August 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-013-0530-z.

Of:2014:ECF

[OS14] G. Of and O. Steinbach. On the ellipticity of coupled finite element and one-equation boundary element methods for boundary value problems. Numerische Mathematik, 127(3):567–593, July 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-013-0593-x.

Ostermann:2019:TET

[OS19] Alexander Ostermann and Chunmei Su. Two exponential-type integrators for the ”good” Boussinesq equation. Numerische Mathematik, 143(3):683–712, November 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-019-01064-4.

Okayama:2013:SMA

[OTMS13] Tomoaki Okayama, Ken’ichiro Tanaka, Takayasu Matsuo, and Masaaki Sugihara. DE-Sinc methods have almost the same convergence property as SE–Sinc methods even for a family of functions fitting the SE–Sinc methods. Numerische Mathematik, 125(3):511–543, November 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-013-0540-x.

Ortner:2011:PEE

[OW11] C. Ortner and W. Wollner. A priori error estimates for optimal control problems with pointwise constraints on the gradient of the state. Numerische Mathematik, 118(3): 587–600, July 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=118&issue=3&spage=587.

Owen:2016:CEQ

[Owe16] Art B. Owen. A constraint on extensible quadrature rules. Numerische Mathematik, 132(3):511–518, March 2016. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-015-0724-7.

Oksa:2017:AQC

[OYV17] Gabriel Oksa, Yusaku Yamamoto, and Marián Vajtersic. Asymptotic quadratic convergence of the serial block-Jacobi EVD algorithm for Hermitian matrices. Numerische Mathematik, 136(4): 1071–1095, August 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Oswald:2017:RRS

[OZ17] Peter Oswald and Weiqi Zhou. Random reordering in SORtype methods. Numerische Mathematik, 135(4):1207–1220, April 2017. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s00211-016-0829-7;   http:
//link.springer.com/article/10.1007/s00211-016-0829-7.

See correction [OZ23].

Oswald:2023:CRR

[OZ23] Peter Oswald and Weiqi Zhou. Correction: Random reordering in SOR-type methods. Numerische Mathematik, 154(3-4):521–525, August 2023. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL https://link.springer.com/article/
10.1007/s00211-023-01365-9.

Peterseim:2013:FEN

[PC13] Daniel Peterseim and Carsten Carstensen. Finite element network approximation of conductivity in particle composites. Numerische Mathematik, 124(1):73–97, May 2013. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-012-0509-1.

Pinto:2018:CLT

[PCCC18] Martin Campos Pinto, José A. Carrillo, Frédérique Charles, and Young-Pil Choi. Convergence of a linearly transformed particle method for aggregation equations. Numerische Mathematik, 139(4):743–793, August 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https:/

/link.springer.com/article/10.1007/s00211-018-0958-2;
 https://link.springer.com/content/pdf/10.1007/s00211-
018-0958-2.pdf.

Petcu:2012:EDP

[Pet12] Madalina Petcu. Exponential decay of the power spectrum and finite dimensionality for solutions of the three dimensional primitive equations. Numerische Mathematik, 120(1): 89–116, January 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=120&issue=1&spage=89.

Pearson:2017:FIP

[PG17] John W. Pearson and Jacek Gondzio. Fast interior point solution of quadratic programming problems arising from PDE-constrained optimization. Numerische Mathematik, 137 (4):959–999, December 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https:/

/link.springer.com/article/10.1007/s00211-017-0892-8;
 https://link.springer.com/content/pdf/10.1007/s00211-
017-0892-8.pdf.

Peherstorfer:2018:CAM

[PGW18] Benjamin Peherstorfer, Max Gunzburger, and Karen Willcox. Convergence analysis of multifidelity Monte Carlo estimation. Numerische Mathematik, 139(3):683–707, July 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
018-0945-7.

Pierre:2018:CEA

[Pie18] Morgan Pierre. Convergence of exponential attractors for a time semi-discrete reaction-diffusion equation. Numerische Mathematik, 139(1):121–153, May 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/s00211-017-0937-z.

Pinelis:2018:AEM

[Pin18] Iosif Pinelis. An alternative to the Euler–Maclaurin summation formula: approximating sums by integrals only. Numerische Mathematik, 140(3):755–790, November 2018. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
018-0978-y.

Plaskota:2015:AIU

[Pla15] Leszek Plaskota. Automatic integration using asymptotically optimal adaptive Simpson quadrature. Numerische Mathematik, 131(1):173–198, September 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/link.springer.com/article/10.1007/s00211-014-0684-3;
 http://link.springer.com/content/pdf/10.1007/s00211-
014-0684-3.pdf.

Pearson:2018:RSP

[PPS18] John W. Pearson, Jennifer Pestana, and David J. Silvester. Refined saddle-point preconditioners for discretized Stokes problems. Numerische Mathematik, 138(2):331–363, February 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00211-017-0908-4;   https://link.springer.com/content/
pdf/10.1007/s00211-017-0908-4.pdf.

Pranic:2013:RRO

[PR13] Miroslav S. Pranić and Lothar Reichel. Recurrence relations for orthogonal rational functions. Numerische Mathematik, 123 (4):629–642, April 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-012-0502-8.

Pruitt:2014:LTS

[Pru14] Michael Pruitt. Large time step maximum norm regularity of L-stable difference methods for parabolic equations. Numerische Mathematik, 128(3):551–587, November 2014. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-014-0617-1.

Pechstein:2011:AFM

[PS11] Clemens Pechstein and Robert Scheichl. Analysis of FETI methods for multiscale PDEs. Part II: interface variation. Numerische Mathematik, 118(3):485–529, July 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=118&issue=3&spage=485.

Park:2013:QME

[PS13] Chunjae Park and Dongwoo Sheen. A quadrilateral Morley element for biharmonic equations. Numerische Mathematik, 124 (2):395–413, June 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-013-0517-9.

Pejcev:2016:EBG

[PS16] Aleksandar V. Pejcev and Miodrag M. Spalević. The error bounds of Gauss–Radau quadrature formulae with Bernstein– Szegő weight functions. Numerische Mathematik, 133(1):177–201, May 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-015-0740-7.

Pechstein:2017:TMR

[PS17a] Astrid S. Pechstein and Joachim Schöberl. The TDNNS method for Reissner–Mindlin plates. Numerische Mathematik, 137(3):713–740, November 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https:/

/link.springer.com/article/10.1007/s00211-017-0883-9;
 https://link.springer.com/content/pdf/10.1007/s00211-
017-0883-9.pdf.

Pozzi:2017:CSF

[PS17b] Paola Pozzi and Björn Stinner. Curve shortening flow coupled to lateral diffusion. Numerische Mathematik, 135(4):1171– 1205, April 2017. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s00211-016-0828-8;   http:
//link.springer.com/article/10.1007/s00211-016-0828-8.

Pechstein:2018:ATM

[PS18] Astrid S. Pechstein and Joachim Schöberl. An analysis of the TDNNS method using natural norms. Numerische Mathematik, 139(1):93–120, May 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https:/

/link.springer.com/article/10.1007/s00211-017-0933-3;
 https://link.springer.com/content/pdf/10.1007/s00211-
017-0933-3.pdf.

Pazoto:2010:USN

[PSV10] A. F. Pazoto, M. Sepúlveda, and O. Vera Villagrán. Uniform stabilization of numerical schemes for the critical generalized Korteweg–de Vries equation with damping. Numerische Mathematik, 116(2):317–356, August 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=116&issue=2&spage=317.

Papez:2018:ELA

[PSV18] J. Papez, Z. Strakos, and M. Vohralík. Estimating and localizing the algebraic and total numerical errors using flux reconstructions. Numerische Mathematik, 138(3):681–721, March 2018. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
017-0915-5.

Pytlak:2010:PCG

[PT10] R. Pytlak and T. Tarnawski. Preconditioned conjugate gradient algorithms for nonconvex problems with box constraints. Numerische Mathematik, 116(1):149–175, July 2010. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=116&issue=1&spage=
149.

Pham:2012:DDM

[PT12] Duong Pham and Thanh Tran. A domain decomposition method for solving the hypersingular integral equation on the sphere with spherical splines. Numerische Mathematik, 120(1):117–151, January 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=120&issue=1&spage=117.

Pham:2014:SEP

[PT14] T. D. Pham and T. Tran. Strongly elliptic pseudodifferential equations on the sphere with radial basis functions. Numerische Mathematik, 128(3):589–614, November 2014. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-014-0614-4.

Pollock:2018:UDS

[PZ18] Sara Pollock and Yunrong Zhu. Uniqueness of discrete solutions of nonmonotone PDEs without a globally fine mesh condition. Numerische Mathematik, 139(4):845–865, August 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00211-018-0956-4.

Quarteroni:2011:AFV

[QRB11] Alfio Quarteroni and Ricardo Ruiz-Baier. Analysis of a finite volume element method for the Stokes problem. Numerische Mathematik, 118(4):737–764, August 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=118&issue=4&spage=737.

Qiu:2019:TDB

[QRSZ19] Tianyu Qiu, Alexander Rieder, Francisco-Javier Sayas, and Shougui Zhang. Time-domain boundary integral equation modeling of heat transmission problems. Numerische Mathematik, 143(1):223–259, September 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-019-01040-y;
 https://link.springer.com/content/pdf/10.1007/s00211-
019-01040-y.pdf.

Rademacher:2019:MMA

[Rad19] Andreas Rademacher. Mesh and model adaptivity for frictional contact problems. Numerische Mathematik, 142(3):465–523, July 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL https://link.springer.com/article/10.
1007/s00211-019-01044-8.

Rafiei:2011:RIF

[RB11] A. Rafiei and M. Bollhöfer. Robust incomplete factorization for nonsymmetric matrices. Numerische Mathematik, 118 (2):247–269, June 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=118&issue=2&spage=247.

Reusken:2014:ATD

[RE14] Arnold Reusken and Patrick Esser. Analysis of time discretization methods for Stokes equations with a nonsmooth forcing term. Numerische Mathematik, 126(2):293–319, February 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-013-0564-2.

Reichmann:2012:OST

[Rei12] O. Reichmann. Optimal space–time adaptive wavelet methods for degenerate parabolic PDEs. Numerische Mathematik, 121(2):337–365, June 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=121&issue=2&spage=337.

Renac:2017:RHO

[Ren17] Florent Renac. A robust high-order Lagrange-projection like scheme with large time steps for the isentropic Euler equations. Numerische Mathematik, 135(2):493–519, February 2017. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s00211-016-0807-0;   http:
//link.springer.com/article/10.1007/s00211-016-0807-0.

Rozza:2013:RBA

[RHM13] Gianluigi Rozza, D. B. Phuong Huynh, and Andrea Manzoni. Reduced basis approximation and a posteriori error estimation for Stokes flows in parametrized geometries: roles of the infsup stability constants. Numerische Mathematik, 125(1):115– 152, September 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-013-0534-8.

Ramos:2015:NSS

[RI15] Alberto Gil C. P. Ramos and Arieh Iserles. Numerical solution of Sturm–Liouville problems via Fer streamers. Numerische Mathematik, 131(3):541–565, November 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.com/article/10.1007/s00211-014-0695-0.

Rieger:2019:PCI

[Rie19] Janosch Rieger. Provably convergent implementations of the subdivision algorithm for the computation of invariant objects. Numerische Mathematik, 142(1):149–165, May 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
018-0997-8.

Roman:2017:SAM

[RMS17] Fabio Roman, Carla Manni, and Hendrik Speleers. Spectral analysis of matrices in Galerkin methods based on generalized B-splines with high smoothness. Numerische Mathematik, 135(1):169–216, January 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/link.springer.com/accesspage/article/10.1007/s00211-
016-0796-z;   http://link.springer.com/article/10.1007/
s00211-016-0796-z.

Respaud:2011:ANC

[RS11] Thomas Respaud and Eric Sonnendrücker. Analysis of a new class of forward semi-Lagrangian schemes for the 1D Vlasov Poisson equations. Numerische Mathematik, 118(2): 329–366, June 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=118&issue=2&spage=329.

Rekatsinas:2018:OAW

[RS18] Nikolaos Rekatsinas and Rob Stevenson. An optimal adaptive wavelet method for first order system least squares. Numerische Mathematik, 140(1):191–237, September 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL

https://link.springer.com/article/10.1007/s00211-018-
0961-7;   https://link.springer.com/content/pdf/10.1007/
s00211-018-0961-7.pdf.

Richter:2013:ENR

[RSV13] Thomas Richter, Andreas Springer, and Boris Vexler. Efficient numerical realization of discontinuous Galerkin methods for temporal discretization of parabolic problems. Numerische Mathematik, 124(1):151–182, May 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-012-0511-7.

Runborg:2014:AHO

[Run14] Olof Runborg. Analysis of high order fast interface tracking methods. Numerische Mathematik, 128(2):339–375, October 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-014-0613-5.

Rushai:2010:NSS

[Rus10] V. D. Rushai. Numerical solution of some partial differential equations by means of a deterministic method of approximate functional integration. Numerische Mathematik, 115(1):153–167, March 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=115&issue=1&spage=153.

Rosch:2012:PEE

[RW12] Arnd Rösch and Daniel Wachsmuth. A-posteriori error estimates for optimal control problems with state and control constraints. Numerische Mathematik, 120(4):733–762, April 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=120&issue=4&spage=
733.

Rieger:2017:SIS

[RW17] Christian Rieger and Holger Wendland. Sampling inequalities for sparse grids. Numerische Mathematik, 136(2):439–466, June 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stewart:2010:OCS

[SA10] David E. Stewart and Mihai Anitescu. Optimal control of systems with discontinuous differential equations. Numerische Mathematik, 114(4):653–695, February 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=114&issue=4&spage=653.

Saibaba:2017:RMF

[SAI17] Arvind K. Saibaba, Alen Alexanderian, and Ilse C. F. Ipsen. Randomized matrix-free trace and log-determinant estimators.

Numerische Mathematik, 137(2):353–395, October 2017. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
017-0880-z.

Sauer:2017:PMS

[Sau17] Tomas Sauer. Prony’s method in several variables. Numerische Mathematik, 136(2):411–438, June 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sayas:2013:EEG

[Say13] Francisco-Javier Sayas. Energy estimates for Galerkin semidiscretizations of time domain boundary integral equations. Numerische Mathematik, 124(1):121–149, May 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.com/article/10.1007/s00211-
012-0506-4.

Scott:2013:IBSa

[SB13a] Joseph K. Scott and Paul I. Barton. Interval bounds on the solutions of semi-explicit index-one DAEs. Part 1: analysis. Numerische Mathematik, 125(1):1–25, September 2013. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-013-0531-y.

Scott:2013:IBSb

[SB13b] Joseph K. Scott and Paul I. Barton. Interval bounds on the solutions of semi-explicit index-one DAEs. Part 2: computation. Numerische Mathematik, 125(1):27–60, September 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-013-0532-x.

Schaback:2010:UMM

[Sch10] Robert Schaback. Unsymmetric meshless methods for operator equations. Numerische Mathematik, 114(4):629–651, February 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0029-599X&volume=114&issue=4&spage=

Schweitzer:2011:MPP

[Sch11a] Marc Alexander Schweitzer. Multilevel particle-partition of unity method. Numerische Mathematik, 118(2):307–328, June 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=118&issue=2&spage=
307.

Schweitzer:2011:SEL

[Sch11b] Marc Alexander Schweitzer. Stable enrichment and local preconditioning in the particle-partition of unity method. Numerische Mathematik, 118(1):137–170, May 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=118&issue=1&spage=137.

Schaback:2016:AWP

[Sch16] Robert Schaback. All well-posed problems have uniformly stable and convergent discretizations. Numerische Mathematik, 132 (3):597–630, March 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-015-0731-8.

Sun:2013:EMS

[SCHH13] Jian-Qing Sun, Xiang-Ke Chang, Yi He, and Xing-Biao Hu. An extended multistep Shanks transformation and convergence acceleration algorithm with their convergence and stability analysis. Numerische Mathematik, 125(4):785–809, December 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-013-0549-1.

Spillane:2014:ARC

[SDH^{+}14] N. Spillane, V. Dolean, P. Hauret, F. Nataf, C. Pechstein, and R. Scheichl. Abstract robust coarse spaces for systems of PDEs via generalized eigenproblems in the overlaps. Numerische Mathematik, 126(4):741–770, April 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-013-0576-y.

Segura:2013:CCZ

[Seg13] Javier Segura. Computing the complex zeros of special functions. Numerische Mathematik, 124(4):723–752, August 2013.

CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-013-0528-6.

Schiela:2011:IPA

[SG11] Anton Schiela and Andreas Günther. An interior point algorithm with inexact step computation in function space for state constrained optimal control. Numerische Mathematik, 119(2):373–407, October 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=119&issue=2&spage=373.

Seidman:2012:FEA

[SG… 12] Thomas I. Seidman, Matthias K. Gobbert, and David W. Trott . . .. Finite element approximation for time-dependent diffusion with measure-valued source. Numerische Mathematik, 122(4): 709–723, December 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-012-0474-8.

Schmidt:2017:AET

[SH17] Kersten Schmidt and Ralf Hiptmair. Asymptotic expansion techniques for singularly perturbed boundary integral equations. Numerische Mathematik, 137(2):397–415, October 2017. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
017-0881-y.

Sun:2018:QOC

[SH18] Pengtao Sun and Xuehai Huang. Quasi-optimal convergence rate for an adaptive hybridizable C^{0} discontinuous Galerkin method for Kirchhoff plates. Numerische Mathematik, 139(4):795–829, August 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL https://link.springer.com/article/
10.1007/s00211-018-0953-7.

Sprung:2019:HOC

[SH19] Benjamin Sprung and Thorsten Hohage. Higher order convergence rates for Bregman iterated variational regularization of inverse problems. Numerische Mathematik, 141(1):215–252, January 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL https://link.springer.com/article/
10.1007/s00211-018-0987-x.

Sidi:2011:ACG

[Sid11a] Avram Sidi. Acceleration of convergence of general linear sequences by the Shanks transformation. Numerische Mathematik, 119(4):725–764, December 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=119&issue=4&spage=725.

Sidje:2011:STT

[Sid11b] Roger B. Sidje. On the simultaneous tridiagonalization of two symmetric matrices. Numerische Mathematik, 118(3): 549–566, July 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=118&issue=3&spage=549.

Singler:2012:BPM

[Sin12] John R. Singler. Balanced POD for model reduction of linear PDE systems: convergence theory. Numerische Mathematik, 121(1):127–164, May 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=121&issue=1&spage=127.

Storath:2019:SSD

[SKW19] Martin Storath, Lukas Kiefer, and Andreas Weinmann. Smoothing for signals with discontinuities using higher order Mumford– Shah models. Numerische Mathematik, 143(2):423–460, October 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245

(electronic). URL https://link.springer.com/article/10.
1007/s00211-019-01052-8.

Speleers:2016:EQI

[SM16] Hendrik Speleers and Carla Manni. Effortless quasi-interpolation in hierarchical spaces. Numerische Mathematik, 132(1):155– 184, January 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-015-0711-z.

Smyrlis:2010:MFM

[Smy10] Yiorgos-Sokratis Smyrlis. Mathematical foundation of the MFS for certain elliptic systems in linear elasticity. Numerische Mathematik, 115(2):319–340, April 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=115&issue=2&spage=319.

Sorokina:2010:ISM

[Sor10] T. Sorokina. Intrinsic supersmoothness of multivariate splines. Numerische Mathematik, 116(3):421–434, September 2010. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=116&issue=3&spage=
421.

Sousedik:2013:NBS

[Sou13] Bedrich Sousedík. Nested BDDC for a saddle-point problem. Numerische Mathematik, 125(4):761–783, December 2013. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-013-0548-2.

Schwab:2011:FEN

[SS11] Christoph Schwab and Rob Stevenson. Fast evaluation of nonlinear functionals of tensor product wavelet expansions. Numerische Mathematik, 119(4):765–786, December 2011. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=119&issue=4&spage=
765.

Saad:2015:CFV

[SS15] Bilal Saad and Mazen Saad. A combined finite volumenonconforming finite element scheme for compressible two phase flow in porous media. Numerische Mathematik, 129(4):691–722, April 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-014-0651-z.

Smears:2016:DGF

[SS16] Iain Smears and Endre Süli. Discontinuous Galerkin finite element methods for time-dependent Hamilton–Jacobi-Bellman equations with Cordes coefficients. Numerische Mathematik, 133 (1):141–176, May 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-015-0741-6.

Schlichting:2018:AIU

[SS18] André Schlichting and Christian Seis. Analysis of the implicit upwind finite volume scheme with rough coefficients. Numerische Mathematik, 139(1):155–186, May 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https:/

/link.springer.com/article/10.1007/s00211-017-0935-1.

SanMartin:2012:MLG

[SSS12] Jorge San Martín, Jean-François Scheid, and Loredana Smaranda. A modified Lagrange–Galerkin method for a fluid-rigid system with discontinuous density. Numerische Mathematik, 122(2):341–382, October 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=122&issue=2&spage=341.

Shank:2016:ELR

[SSS16] Stephen D. Shank, Valeria Simoncini, and Daniel B. Szyld. Efficient low-rank solution of generalized Lyapunov equations. Numerische Mathematik, 134(2):327–342, October 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s00211-015-0777-7;   http://link.springer.com/article/
10.1007/s00211-015-0777-7.

Schroder:2016:BEA

[ST16] Christian Schröder and Leo Taslaman. Backward error analysis of the shift-and-invert Arnoldi algorithm. Numerische Mathematik, 133(4):819–843, August 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/link.springer.com/article/10.1007/s00211-015-0759-9;
 http://link.springer.com/content/pdf/10.1007/s00211-
015-0759-9.pdf.

Stephan:2018:CTP

[ST18] Ernst P. Stephan and Matthias T. Teltscher. Collocation with trigonometric polynomials for integral equations to the mixed boundary value problem. Numerische Mathematik, 140(1):153– 190, September 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-018-0960-8.

Stenberg:2010:NMF

[Ste10] Rolf Stenberg. A nonstandard mixed finite element family. Numerische Mathematik, 115(1):131–139, March 2010. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=115&issue=1&spage=
131.

Steinbach:2014:BEM

[Ste14] O. Steinbach. Boundary element methods for variational inequalities. Numerische Mathematik, 126(1):173–197, January 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-013-0554-4.

Stern:2015:BSP

[Ste15] Ari Stern. Banach space projections and Petrov–Galerkin estimates. Numerische Mathematik, 130(1):125–133, May 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-014-0658-5.

Stiemer:2010:GMM

[Sti10] Marcus Stiemer. A Galerkin method for mixed parabolicelliptic partial differential equations. Numerische Mathematik, 116(3):435–462, September 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=116&issue=3&spage=435.

Strobach:2015:COO

[Str15] Peter Strobach. Constrained optimization of the overdetermined Zimmermann compressor for nonorthogonal joint matrix diagonalization. Numerische Mathematik, 129(3):563–586, March 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-014-0645-x.

Sun:2016:MFQ

[Sun16] Jiguang Sun. A mixed FEM for the quad-curl eigenvalue problem. Numerische Mathematik, 132(1):185–200, January 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-015-0708-7.

Spigler:2012:SFM

[SV12]  Renato Spigler and Marco Vianello.  The ‘phase func-

tion’ method to solve second-order asymptotically polynomial differential equations. Numerische Mathematik, 121 (3):565–586, July 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=121&issue=3&spage=565.

Sauter:2013:GMR

[SV13] S. Sauter and A. Veit. A Galerkin method for retarded boundary integral equations with smooth and compactly supported temporal basis functions. Numerische Mathematik, 123(1):145– 176, January 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-012-0483-7.

Sauter:2016:ATD

[SV16] S. Sauter and A. Veit. Adaptive time discretization for retarded potentials. Numerische Mathematik, 132(3):569–595, March 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-015-0726-5.

Schumaker:2011:SST

[SW11a] Larry L. Schumaker and Lujun Wang. Spline spaces on TRmeshes with hanging vertices. Numerische Mathematik, 118 (3):531–548, July 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=118&issue=3&spage=531.

Steinbach:2011:SBE

[SW11b] O. Steinbach and M. Windisch. Stable boundary element domain decomposition methods for the Helmholtz equation. Numerische Mathematik, 118(1):171–195, May 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=118&issue=1&spage=171.

Schumacher:2019:VCD

[SW19] Henrik Schumacher and Max Wardetzky. Variational convergence of discrete minimal surfaces. Numerische Mathematik, 141(1): 173–213, January 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-018-0993-z.

Si:2016:USE

[SWS16] Zhiyong Si, Jilu Wang, and Weiwei Sun. Unconditional stability and error estimates of modified characteristics FEMs for the Navier–Stokes equations. Numerische Mathematik, 134 (1):139–161, September 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/link.springer.com/accesspage/article/10.1007/s00211-
015-0767-9;   http://link.springer.com/article/10.1007/
s00211-015-0767-9.

Szyld:2013:LCA

[SX13] Daniel B. Szyld and Fei Xue. Local convergence analysis of several inexact Newton-type algorithms for general nonlinear eigenvalue problems. Numerische Mathematik, 123(2):333–362, February 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-012-0489-1.

Szyld:2015:LCNa

[SX15a] Daniel B. Szyld and Fei Xue. Local convergence of Newtonlike methods for degenerate eigenvalues of nonlinear eigenproblems. I. Classical algorithms. Numerische Mathematik, 129(2): 353–381, February 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-014-0639-8.

Szyld:2015:LCNb

[SX15b] Daniel B. Szyld and Fei Xue. Local convergence of Newton-like methods for degenerate eigenvalues of nonlinear eigenproblems: II. Accelerated algorithms. Numerische Mathematik, 129(2): 383–403, February 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-014-0640-2.

Scholz:2017:AEW

[SY17] Stephan Scholz and Harry Yserentant. On the approximation of electronic wavefunctions by anisotropic Gauss and Gauss– Hermite functions. Numerische Mathematik, 136(3):841–874, July 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Shi:2013:ERA

[SZ13] Zhongci Shi and Xuying Zhao. Error reduction of the adaptive conforming and nonconforming finite element methods with red-green refinement. Numerische Mathematik, 123(3):553–584, March 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-012-0495-3.

Sun:2012:FDM

[SzS12] Weiwei Sun and Zhi zhong Sun. Finite difference methods for a nonlinear and strongly coupled heat and moisture transport system in textile materials. Numerische Mathematik, 120(1):153–187, January 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=120&issue=1&spage=153.

Takacs:2015:RAO

[Tak15] Stefan Takacs. A robust all-at-once multigrid method for the Stokes control problem. Numerische Mathematik, 130(3): 517–540, July 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-014-0674-5.

Tanaka:2013:SMA

[TOMS13] Ken’ichiro Tanaka, Tomoaki Okayama, Takayasu Matsuo, and Masaaki Sugihara. DE-Sinc methods have almost the same convergence property as SE–Sinc methods even for a family of functions fitting the SE–Sinc methods. Numerische Mathematik, 125(3):545–568, November 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-013-0541-9.

Towers:2018:COG

[Tow18] John D. Towers. Convergence via OSLC of the Godunov scheme for a scalar conservation law with time and space flux discontinuities. Numerische Mathematik, 139(4):939–969, August 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00211-018-0957-3.

Teckentrup:2013:FAM

[TSGU13] A. L. Teckentrup, R. Scheichl, M. B. Giles, and E. Ullmann. Further analysis of multilevel Monte Carlo methods for elliptic PDEs with random coefficients. Numerische Mathematik, 125(3): 569–600, November 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-013-0546-4.

Tregubov:2015:GMS

[TT15] I. Tregubov and T. Tran. A Galerkin method with spherical splines for the shallow water equations on a sphere: error analysis. Numerische Mathematik, 129(4):783–814, April 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-014-0654-9.

Tu:2011:TLB

[Tu11] Xuemin Tu. A three-level BDDC algorithm for a saddle point problem. Numerische Mathematik, 119(1):189– 217, September 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=119&issue=1&spage=189.

Tone:2015:LTD

[TWW15] Florentina Tone, Xiaoming Wang, and Djoko Wirosoetisno. Long-time dynamics of 2 d double-diffusive convection: analysis and/of numerics. Numerische Mathematik, 130(3):541–566, July 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-014-0670-9.

Tran:2017:AQO

[TWZ17] Hoang Tran, Clayton G. Webster, and Guannan Zhang. Analysis of quasi-optimal polynomial approximations for parameterized PDEs with deterministic and stochastic coefficients. Numerische Mathematik, 137(2):451–493, October 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/s00211-017-0878-6.

Vandebril:2011:UUS

[VD11] Raf Vandebril and Gianna M. Del Corso. A unification of unitary similarity transforms to compressed representations. Numerische Mathematik, 119(4):641–665, December 2011. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=119&issue=4&spage=
641.

Venel:2011:NSC

[Ven11] Juliette Venel. A numerical scheme for a class of sweeping processes. Numerische Mathematik, 118(2):367–400, June 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=118&issue=2&spage=
367.

Verwer:2010:RKM

[Ver10] J. G. Verwer. Runge–Kutta methods and viscous wave equations. Numerische Mathematik, 115(3):485–507, May 2010. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=115&issue=3&spage=
485.

Vermeeren:2017:MEV

[Ver17] Mats Vermeeren. Modified equations for variational integrators. Numerische Mathematik, 137(4):1001–1037, December 2017. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
017-0896-4.

Vico:2014:BIE

[VGG14] Felipe Vico, Leslie Greengard, and Zydrunas Gimbutas. Boundary integral equation analysis on the sphere. Numerische Mathematik, 128(3):463–487, November 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.com/article/10.1007/s00211-014-0619-z.

Barel:2010:IDS

[VVVF10] Marc Van Barel, Raf Vandebril, Paul Van Dooren, and Katrijn Frederix. Implicit double shift QR-algorithm for companion matrices. Numerische Mathematik, 116(2):177–212, August 2010.

CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=116&issue=2&spage=
177.

vanyen:2014:ALS

[vyKS14] Romain Nguyen van yen, Dmitry Kolomenskiy, and Kai Schneider. Approximation of the Laplace and Stokes operators with Dirichlet boundary conditions through volume penalization: a spectral viewpoint. Numerische Mathematik, 128(2):301–338, October 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-014-0610-8.

Wachsmuth:2017:CAC

[Wac17] Gerd Wachsmuth. Conforming approximation of convex functions with the finite element method. Numerische Mathematik, 137(3): 741–772, November 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-017-0884-8.

Wang:2012:ESO

[Wan12] Xiaoming Wang. An efficient second order in time scheme for approximating long time statistical properties of the two dimensional Navier–Stokes equations. Numerische Mathematik, 121(4):753–779, August 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=121&issue=4&spage=753.

Wang:2018:USC

[Wan18] Jilu Wang. Unconditional stability and convergence of Crank– Nicolson Galerkin FEMs for a nonlinear Schrödinger–Helmholtz system. Numerische Mathematik, 139(2):479–503, June 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00211-017-0944-0.

Wulff:2016:RKT

[WE16] Claudia Wulff and Chris Evans. Runge–Kutta time semidiscretizations of semilinear PDEs with non-smooth data. Numerische Mathematik, 134(2):413–440, October 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.com/article/10.1007/s00211-
015-0776-8;   http://link.springer.com/content/pdf/10.
1007/s00211-015-0776-8.pdf.

Weisser:2011:REE

[Wei11] Steffen Weißer. Residual error estimate for BEM-based FEM on polygonal meshes. Numerische Mathematik, 118(4): 765–788, August 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=118&issue=4&spage=765.

Wendland:2010:MAS

[Wen10] Holger Wendland. Multiscale analysis in Sobolev spaces on bounded domains. Numerische Mathematik, 116(3):493– 517, September 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=116&issue=3&spage=493.

Wilkins:2013:MBM

[WG13] Gautam Wilkins and Ming Gu. A modified Brent’s method for finding zeros of functions. Numerische Mathematik, 123(1): 177–188, January 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-012-0480-x.

Wang:2014:DGM

[WHC14] Fei Wang, Weimin Han, and Xiaoliang Cheng. Discontinuous Galerkin methods for solving a quasistatic contact problem. Numerische Mathematik, 126(4):771–800, April 2014. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-013-0574-0.

Williams:2019:ADG

[Wil19] David M. Williams. An analysis of discontinuous Galerkin methods for the compressible Euler equations: entropy and L_{2} stability. Numerische Mathematik, 141(4):1079–1120, April 2019. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
019-01027-9.

Wu:2010:ASE

[WMS10] Fuke Wu, Xuerong Mao, and Lukas Szpruch. Almost sure exponential stability of numerical solutions for stochastic delay differential equations. Numerische Mathematik, 115(4): 681–697, June 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=115&issue=4&spage=681.

Wormell:2019:SGM

[Wor19] Caroline Wormell. Spectral Galerkin methods for transfer operators in uniformly expanding dynamics. Numerische Mathematik, 142(2):421–463, June 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/s00211-019-01031-z.

Wu:2019:PPH

[WS19] Kailiang Wu and Chi-Wang Shu. Provably positive high-order schemes for ideal magnetohydrodynamics: analysis on general meshes. Numerische Mathematik, 142(4):995–1047, August 2019. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00211-019-01042-w.

Wang:2012:SNR

[WX12] Feng Wang and Xuejun Xu. Some new residual-based a posteriori error estimators for the mortar finite element methods. Numerische Mathematik, 120(3):543–571, March 2012. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=120&issue=3&spage=
543.

Wheeler:2012:MFM

[WXY12] Mary Wheeler, Guangri Xue, and Ivan Yotov. A multipoint flux mixed finite element method on distorted quadrilaterals and hexahedra. Numerische Mathematik, 121(1): 165–204, May 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=121&issue=1&spage=165.

Wang:2010:PSI

[WZ10] Wansheng Wang and Chengjian Zhang. Preserving stability implicit Euler method for nonlinear Volterra and neutral functional differential equations in Banach space. Numerische Mathematik, 115(3):451–474, May 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=115&issue=3&spage=451.

Wang:2013:AEF

[WZH13] Haiyong Wang, Lun Zhang, and Daan Huybrechs. Asymptotic expansions and fast computation of oscillatory Hilbert transforms. Numerische Mathematik, 123(4):709–743, April 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-012-0501-9.

Xiang:2010:EBA

[XCW10] Shuhuang Xiang, Xiaojun Chen, and Haiyong Wang. Error bounds for approximation in Chebyshev points. Numerische Mathematik, 116(3):463–491, September 2010. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=116&issue=3&spage=
463.

Xue:2017:HAD

[XL17] Jungong Xue and Ren-Cang Li. Highly accurate doubling algorithms for M-matrix algebraic Riccati equations. Numerische Mathematik, 135(3):733–767, March 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/link.springer.com/accesspage/article/10.1007/s00211-
016-0815-0;   http://link.springer.com/article/10.1007/
s00211-016-0815-0.

Xue:2012:ASMb

[XXL12a] Jungong Xue, Shufang Xu, and Ren-Cang Li. Accurate solutions of M-matrix algebraic Riccati equations. Numerische Mathematik, 120(4):671–700, April 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=120&issue=4&spage=671.

Xue:2012:ASMa

[XXL12b] Jungong Xue, Shufang Xu, and Ren-Cang Li. Accurate solutions of M-matrix Sylvester equations. Numerische Mathematik, 120(4):639–670, April 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=120&issue=4&spage=639.

Xu:2010:ALQ

[XZ10] Jinchao Xu and Qingsong Zou. Analysis of linear and quadratic simplicial finite volume methods for elliptic equations. Numerische Mathematik, 114(3):469–492, January 2010. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=114&issue=3&spage=
469.

Yuan:2014:SCM

[YBTB14] Jing Yuan, Egil Bae, Xue-Cheng Tai, and Yuri Boykov. A spatially continuous max-flow and min-cut framework for binary labeling problems. Numerische Mathematik, 126(3):559–587, March 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-013-0569-x.

Yang:2013:DGM

[YS13] Yang Yang and Chi-Wang Shu. Discontinuous Galerkin method for hyperbolic equations involving δ-singularities: negative-order norm error estimates and applications. Numerische Mathematik, 124(4):753–781, August 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-013-0526-8.

Yin:2014:CFM

[YWY14] Li Yin, Jiming Wu, and Yanzhong Yao. A cell functional minimization scheme for domain decomposition method on nonorthogonal and non-matching meshes. Numerische Mathematik, 128(4):773–804, December 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-014-0623-3.

Zampini:2013:BNN

[Zam13] Stefano Zampini. Balancing Neumann–Neumann methods for the cardiac bidomain model. Numerische Mathematik, 123(2): 363–393, February 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-012-0488-2.

Zhang:2014:HOS

[ZBB14] Qinghui Zhang, Uday Banerjee, and Ivo Babuska. Higher order stable generalized finite element method. Numerische Mathematik, 128(1):1–29, September 2014. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-014-0609-1.

Zhao:2018:RIN

[ZBJ18] Zhi Zhao, Zheng-Jian Bai, and Xiao-Qing Jin. A Riemannian inexact Newton–CG method for constructing a nonnegative matrix with prescribed realizable spectrum. Numerische Mathematik, 140(4):827–855, December 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/s00211-018-0982-2.

Zhu:2017:IAP

[ZDQ17] Shengfeng Zhu, Luca Dedè, and Alfio Quarteroni. Isogeometric analysis and proper orthogonal decomposition for parabolic problems. Numerische Mathematik, 135(2):333–370, February 2017. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s00211-016-0802-5;   http:
//link.springer.com/article/10.1007/s00211-016-0802-5.

Zhang:2016:ADG

[Zha16a] Sheng Zhang. Analysis of a discontinuous Galerkin method for the bending problem of Koiter shell. Numerische Mathematik, 133 (2):333–370, June 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/s00211-015-0747-0.

Zhang:2016:SFE

[Zha16b] Shuo Zhang. Stable finite element pair for Stokes problem and discrete Stokes complex on quadrilateral grids. Numerische Mathematik, 133(2):371–408, June 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-015-0749-y.

Zhang:2018:RDF

[Zha18] Shuo Zhang. Regular decomposition and a framework of order reduced methods for fourth order problems. Numerische Mathematik, 138(1):241–271, January 2018. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/s00211-017-0902-x.

Zeng:2012:CQM

[ZLL^{+}12] Wei Zeng, Lok Ming Lui, Feng Luo, Tony Fan-Cheong Chan, Shing-Tung Yau, et al. Computing quasiconformal maps using an auxiliary metric and discrete curvature flow. Numerische Mathematik, 121(4):671–703, August 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=121&issue=4&spage=671.

Zhang:2016:SWC

[ZRK16] Zhongqiang Zhang, Boris Rozovskii, and George Em Karniadakis. Strong and weak convergence order of finite element methods for stochastic PDEs with spatial white noise. Numerische Mathematik, 134(1):61–89, September 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/link.springer.com/accesspage/article/10.1007/s00211-
015-0768-8;   http://link.springer.com/article/10.1007/
s00211-015-0768-8.

Zhang:2012:MEP

[ZS12] Xiangxiong Zhang and Chi-Wang Shu. A minimum entropy principle of high order schemes for gas dynamics equations. Numerische Mathematik, 121(3):545–563, July 2012. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=121&issue=3&spage=
545.

Zhang:2014:EET

[ZS14] Qiang Zhang and Chi-Wang Shu. Error estimates for the third order explicit Runge–Kutta discontinuous Galerkin method for a linear hyperbolic equation in one-dimension with discontinuous initial data. Numerische Mathematik, 126(4):703–740, April 2014.

CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00211-013-0573-1.

Zhou:2017:FVM

[ZS17] Guanyu Zhou and Norikazu Saito. Finite volume methods for a Keller–Segel system: discrete energy, error estimates and numerical blow-up analysis. Numerische Mathematik, 135(1):265–311, January 2017. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s00211-016-0793-2;   http:
//link.springer.com/article/10.1007/s00211-016-0793-2.

Zelati:2012:MAT

[ZT12] Michele Coti Zelati and Florentina Tone. Multivalued attractors and their approximation: applications to the Navier–Stokes equations. Numerische Mathematik, 122(3):421–441, November 2012. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=122&issue=3&spage=
421.

Zou:2011:HEE

[ZVKG11] Qingsong Zou, Andreas Veeser, Ralf Kornhuber, and Carsten Gräser. Hierarchical error estimates for the energy functional in obstacle problems. Numerische Mathematik, 117 (4):653–677, April 2011. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0029-
599X&volume=117&issue=4&spage=653.

Zhong:2019:RIP

[ZWJ19] Min Zhong, Wei Wang, and Qinian Jin. Regularization of inverse problems by two-point gradient methods in Banach spaces. Numerische Mathematik, 143(3):713–747, November 2019. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/s00211-
019-01068-0.

Zhang:2015:VCF

[ZZ15] Zhimin Zhang and Qingsong Zou. Vertex-centered finite volume schemes of any order over quadrilateral meshes for elliptic boundary value problems. Numerische Mathematik, 130(2):363–393, June 2015. CODEN NUMMA7. ISSN 0029-599X (print), 0945-

3245 (electronic). URL http://link.springer.com/article/
10.1007/s00211-014-0664-7.